### v1

# Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim1

Abstract
In this work, we bridge the gap between wearable sensor technology and personalized AI assistants by enabling Large Language Models (LLMs) to understand time-series tasks like human activity recognition (HAR). Despite the strong reasoning and generalization capabilities of LLMs, leveraging them for sensor data tasks remains largely unexplored. This gap stems from challenges like the lack of semantic context in time-series data, computational limitations, and LLMs’ difficulty processing numerical inputs. To address these issues, we introduce SensorLLM, a two-stage framework to unlock LLMs’ potential for sensor data tasks. In the Sensor-Language Alignment Stage, we introduce special tokens for each sensor channel and automatically generate trend-descriptive text to align sensor data with textual inputs, enabling SensorLLM to capture numerical changes, channel-specific information, and sensor data of varying lengths—capabilities that existing LLMs typically struggle with, all without the need for human annotations. Next, in Task-Aware Tuning Stage, we refine the model for HAR classification using the frozen LLM and alignment module, achieving performance on par with or surpassing state-of-the-art models. We further demonstrate that SensorLLM evolves into an effective sensor learner, reasoner, and classifier through Sensor-Language Alignment, enabling it to generalize across diverse datasets for HAR tasks. We strongly believe our work lays the stepstone for future time-series and text alignment research, offering a path toward foundation models for sensor data. Our codes are available at https://github.com/zechenli03/SensorLLM.
1 Introduction
Human Activity Recognition (HAR) is a time-series (TS) classification task that involves mapping sensor data, such as accelerometer and gyroscope signals, to human activities. Models such as LSTM (Guan & Plötz, 2017; Hammerla et al., 2016) and DeepConvLSTM (Ordóñez & Roggen, 2016) have gained popularity in HAR due to their ability to learn high-level features. However, these models are typically task-specific and struggle to scale across varying sensor configurations and activity sets. In contrast, Large Language Models (LLMs) (Han et al., 2021) have shown remarkable success in integrating diverse data types (Wu et al., 2023b; Yin et al., 2023), including text and images. In this work, we explore how LLMs can unlock new possibilities for HAR by improving the interpretation of sensor data and offering a more scalable, flexible solution to this challenging task.
To enable LLMs to understand effectively and process sensor data (Jin et al., 2023), there are two primary approaches. The first approach involves pretraining or fine-tuning LLMs on sensor data (Zhou et al., 2023b), optimizing them to handle time-series inputs directly. However, this method presents significant challenges, particularly related to computational resource requirements and data limitations. Pretraining LLMs on such data demands immense computational resources, and the limited availability of labeled sensor data, especially for rare activities, leads to class imbalance, making the training process difficult and inefficient. The second approach to enabling LLMs to understand sensor data focuses on using zero-shot and few-shot prompting, where sensor data is transformed into textual formats that LLMs can process without retraining (Ji et al., 2024). However, this introduces several challenges. One major issue is the numerical and temporal nature of data, which poses challenges due to how LLMs tokenize numerical inputs (Nate Gruver & Wilson, 2023). LLM tokenizers, designed for text, are not optimized for numerical values. They struggle with recognizing patterns and contextual nuances, treating consecutive values as independent tokens and ignoring temporal correlations. Sensor data, characterized by its continuous and extensive nature, often exceeds the maximum sequence length, resulting in truncation and potential loss of vital information. Another challenge lies in converting sensor data into text-like embeddings while preserving high-level time-series features, as this transformation risks losing the temporal relationships crucial for complex activity recognition (Spathis & Kawsar, 2024). Handling multi-channel sensor data further complicates this process, as encoding multi-dimensional information in a way that LLMs can interpret effectively is difficult cause the input for LLMs is univariate. Moreover, designing effective prompts that enable LLMs to interpret sensor readings, detect trends, and classify activities is complex, especially when dealing with intricate time-series patterns (Liu et al., 2023b).
To address these challenges, we introduce a novel approach for aligning sensor data with natural language, allowing LLMs to interactively analyze sensor data through text-based commands (see Figure 1) without requiring any modifications to the LLM itself. A significant challenge in this context is managing the annotation bottleneck, which is further exacerbated by the inherent heterogeneity and complexity of sensor data. Unlike image-text pairs that can be more straightforwardly aligned, sensor data comprises multiple modalities with diverse characteristics, making the interpretation and alignment process considerably more difficult. Previous approaches (Jin et al., 2024a; Sun et al., 2024b) have often used condensed text prototypes for alignment. While these methods provide a solution, they face limitations in interpretability, as the mapping between numerical time-series data and abstract text prototypes is not always intuitive. Moreover, such methods often require multiple experiments to select appropriate text prototypes, adding complexity to the process.
Figure 1: SensorLLM can analyze and summarize trends in captured sensor data, facilitating human activity recognition tasks.
In this work, we propose a novel approach that automatically generates descriptive text from time-series data, removing the need for manual annotations. These descriptions are generated using statistical analyses or simple descriptions of numerical trends. We employ pre-defined templates to automate the generation of descriptive texts, offering a scalable, precise, and interpretable solution for aligning time-series data with text in our SensorLLM framework. SensorLLM introduces a two-stage framework process that has been proven effective across various domains  (Xu et al., 2024; Huang et al., 2023; Jiang et al., 2024). In Sensor-Language Alignment Stage, we generate automatic question-answer pairs to align sensor data with text, providing a scalable and interpretable solution. To preserve the temporal features of the sensor data, we use a pretrained encoder to generate sensor embeddings, which are then mapped into a space that the LLM can interpret. This avoids the issues caused by using a text-specific embedding model used by LLM. Furthermore, we introduce special tokens for each sensor channel, enabling the LLM to capture and interpret multi-channel data effectively. This ensures that both time-series characteristics and sensor-specific information are fully aligned with the LLM’s interpretative capabilities. In Task-Aware Tuning Stage, the aligned data is leveraged to address downstream tasks HAR, utilizing the LLM’s reasoning abilities while keeping its parameters frozen. This alignment framework enhances the capabilities of LLMs beyond their original pretraining. Our work demonstrates the effectiveness of SensorLLM, addressing the questions raised by Tan et al. (2024) regarding the broader applicability of LLMs for time-series data. To our knowledge, this is the first approach to incorporate sensor data into LLMs for sensor data analysis and HAR tasks. The key contributions of this work are as follows:
  1. 1.
    TS-Language Alignment: We introduce a human-intuitive approach for aligning time-series data with automatically generated text, removing the need for manual annotations. By leveraging text similarity metrics and human evaluations, we demonstrate that SensorLLM effectively captures both time-series patterns and channel-specific information, enabling robust multimodal understanding.
  2. 2.
    Task-aware Sensor-Language Model: SensorLLM achieves competitive performance across four HAR tasks, either surpassing or matching SOTA models. This success is driven by our modality alignment and the inclusion of task-specific prompts, both of which significantly enhance SensorLLM’s ability to interpret and classify sensor data effectively.
  3. 3.
    Cross-Dataset Generalization: We demonstrate that SensorLLM maintains superior performance in Task-Aware Tuning Stage, even when applied to datasets distinct from those used in alignment stage, underscoring the model’s robustness and generalizability in HAR tasks.
2 Related Work
In this section, we explore recent works that utilize LLMs for human-centric tasks, such as human activity recognition and mobile health sensing. Current approaches can be categorized based on two main aspects: (1) how they integrate numerical time-series data with LLMs, and (2) the specific end task, whether it is time-series forecasting, classification, etc. Zhang et al. (2024) provides a comprehensive review of the existing methodologies that leverage LLM capabilities in processing time-series data. Several works investigated using LLM for numerical data by treating time-series signals as raw text. They used the same tokenization and embedding models as text such as PromptCast  (Xue & Salim, 2023) and LLMTime (Gruver et al., 2024). However, despite LLM’s versatility, they are inherently designed for processing sequential text data and not for heterogeneous time-series numerical data. The challenge lies in enabling LLMs to interpret and learn patterns within multimodal time-series data effectively. As highlighted by Spathis & Kawsar (2024), one significant barrier is the absence of suitable tokenizers for numerical data. LLMs typically employ tokenizers that segment text into smaller units, but these tokenizers are not optimized for representing numerical values. Consequently, they may struggle to recognise repetitive patterns and context, often treating consecutive values as independent tokens and overlooking the crucial temporal relationships inherent in time-series data. To extract more meaningful insights from numerical data before feeding them to LLM, several works proposed the use of time-series encoders and tried to align the embedding space with text embedding models. Liu et al. (2024a) aligns ECG data with corresponding textual descriptions using contrastive learning techniques. Similar approaches have been investigated for aligning motion/IoT sensors with text (Zhou et al., 2023c; Haresamudram et al., 2024; Moon et al., 2023; Xia et al., 2024). Please refer to Appendix A.1 for a more in-depth review of existing works.
3 Methods
In this work, we propose a novel approach to align sensor data with descriptive text using automatically generated question-answer pairs. The aim of this work is to develop a multimodal foundational model with reasoning capabilities to address challenges in analyzing wearable sensor data. As shown in Figure 2, our SensorLLM comprises three core components: 1) a pretrained LLM, 2) a pretrained TS embedder, and 3) an alignment module MLP. The three core modules cooperate in a two-stage framework. In the Sensor-Language Alignment Stage, a generative model aligns sensor readings with text based on user instructions (Liu et al., 2023a) and questions. In Task-Aware Tuning Stage, a classification model is used to perform HAR and maintain an end-to-end alignment. Notably, only the parameters of the MLP in both stages, as well as the classifier in the second stage, are updated during training, while the backbone LLM and TS embedder remain frozen. Consequently, only 5.67% of the parameters in stage 1 and 6% in stage 2 are trainable, ensuring a highly efficient and lightweight training process.
Figure 2: Our proposed SensorLLM framework: (a) Sensor-Language Alignment Stage, where a generative model aligns sensor readings with automatically generated text; (b) Task-Aware Tuning Stage, where a classification model leverages the aligned modalities to perform HAR.
3.1 Sensor-Text Data Generation
A key challenge in aligning time-series data with text for LLM-based tasks is effectively bridging these two modalities. Unlike other multimodal tasks, time-series data typically lacks rich semantic textual information beyond class labels, making manual annotation impractical  (Deldari et al., 2024; Haresamudram et al., 2024). While prior works have used predefined text prototypes for alignment (Sun et al., 2024a; Jin et al., 2024a), we aim to guide the model toward a more human-intended understanding of the data. We argue that time-series data inherently contains semantic information that can be expressed through descriptive text, from simple numerical trends to more complex insights derived from statistical analysis. Our approach focuses on automatically generating descriptive text based on observed trends and fluctuations in the input data. To achieve this, we use a variety of predefined templates to generate diverse question-answer (QA) pairs that describe trend changes in the sensor data, ensuring both accuracy and scalability, while avoiding the potential errors of manual annotation. These templates are randomly combined to produce diverse QA pairs. For example:
  • (1)
    The time-series data represents readings taken from a <S> sensor between $<t_{s}>$ and $<t_{e}>$ seconds.
  • (2)
    To sum up, the data exhibited a <T> trend for a cumulative period of $<t_{t}>$ seconds.
where T and S represent placeholders for a specific trend and sensor type, respectively, and t correspond to numerical values. These automatically generated QA pairs offer a scalable and interpretable method for aligning sensor data and text in the latent space, facilitating more effective understanding and interaction with LLMs. See Appendix A.2 for detailed templates.
3.2 Sensor-Language Alignment
As shown in Figure 2 (a), Sensor-Language Alignment stage employs a generative model designed to create multi-modal sentences that combine single-channel sensor readings with textual descriptions. The sensor data is represented as a matrix $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ indicates the number of sensor channels, and $T$ represents the total number of time steps in the data sequence. Each individual channel’s data, denoted as $\mathbf{X}^{c}$ for channel $c$, is processed independently to maintain the uniqueness of information captured by each channel. The data is then divided into non-overlapping segments, $\mathbf{X}_{S}^{c}$, where $S$ represents the total number of segments. Each segment $x_{s}$ is assigned a random length $l$ within a predefined range. The variation in the length of segments allows the model to effectively learn from a variety of temporal patterns and trend variations across different segments. By learning from segments with different lengths, the generative model becomes more adept at handling the diverse characteristics of sensor data, ensuring that both long-term trends and short-term fluctuations are adequately represented in the generated multi-modal sentences. We employ the Chronos (Ansari et al., 2024) as TS embedder, which generates the segment embedding $\hat{x}_{s}\in\mathbb{R}^{(l+1)\times d_{ts}}$, where $d_{ts}$ is the feature dimension, and the $(l+1)$ accounts for the [EOS] token appended during Chronos tokenization. Unlike the original Chronos implementation, which uses mean normalization A.4, our approach utilizes instance normalization $\tilde{x}_{s}=\frac{x_{s}-\text{mean}(x_{s})}{\text{std}(x_{s})}$ for improved performance. We use Llama3-8b (Touvron et al., 2023) as our LLM backbone.
Alignment Module. To transform the time-series embeddings $\hat{x}{s}$ into text embeddings $\hat{a}_{s}\in\mathbb{R}^{(l+1)\times D}$ for further downstream tasks, we introduce an alignment projection module. This module is implemented as a multi-layer perceptron (MLP), which first maps the sensor embeddings to an intermediate latent space $d_{m}$ and then projects them to the final target dimension $D$. Formally, the projection can be written as: $\hat{a}_{s}=\mathbf{W}_{2}\cdot\sigma(\mathbf{W}_{1}\hat{x}_{s}+\mathbf{b}_{1})+\mathbf{b}_{2}$ where $\mathbf{W}1\in\mathbb{R}^{d_{m}\times d_{ts}}$ and $\mathbf{W}2\in\mathbb{R}^{D\times d_{m}}$ are learnable weights, $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are biases, and $\sigma$ is GELU activation function (Hendrycks & Gimpel, 2016). This projection ensures that the transformed embeddings $\hat{a}_{s}$ are semantically aligned with the text embedding space, making them suitable for subsequent tasks such as generation or classification.
Input Embedding. To enable the LLM to process sensor channel information and align sensor features with text tokens, we introduce two special tokens for each channel (e.g.,<x_acc_start> and <x_acc_end> for the x-axis accelerometer), extending the LLM’s embedding matrix from $\mathbf{E}\in\mathbb{R}^{V\times D}$ to $\mathbf{E}\in\mathbb{R}^{V^{\prime}\times D}$, where $V^{\prime}=V+2c$ and $V$ is the vocabulary size, and $c$ is the number of channels. Then the special token embeddings are concatenated with the aligned sensor embeddings. Finally, the combined sensor embeddings $\hat{o}_{s}\in\mathbb{R}^{(l+3)\times D}$ are concatenated with instruction and question embeddings to form the final input $\hat{z}\in\mathbb{R}^{k\times D}$, where $k$ is the total number of tokens.
Loss Function. Our SensorLLM processes a mixed sequence of tokens, $\mathbf{Z}_{s}=\{z_{s}^{i}\}_{i=1}^{K}$, consisting of both sensor and text embeddings, and generates an output sequence, $\mathbf{Z}_{t}=\{z_{t}^{i}\}_{i=1}^{N}$, where $z_{s}^{i}$ and $z_{t}^{i}\in V^{\prime}$, and $K$ and $N$ represent the number of input and output tokens, respectively. The generative model is trained using a causal language modeling objective, where the next token in the sequence is predicted based on all preceding tokens. To optimize this, we minimize the negative log-likelihood of the generated text tokens ($\mathcal{L}_{gen}=-\sum_{i=0}^{N-1}\log P(z_{t}^{i}|Z_{t}^{<i},z_{s})$). The loss is computed solely for output tokens, excluding those corresponding to instructions, questions, channel special tokens, and sensor data. This ensures that SensorLLM effectively integrates sensor and text embeddings, enabling the generation of coherent and contextually appropriate responses.
3.3 Task-Aware Tuning
As shown in Figure 2 (b), the Task-Aware Tuning stage employs a classification model to align the multimodal sensor-text embeddings with human activity labels. This stage further integrates multi-channel sensor readings with activity labels for HAR. The input sensor data $\mathbf{X}$ is divided into segments of window size $L$ with 50% overlap (Li et al., 2018), forming segments $\mathbf{X}_{S}\in\mathbb{R}^{S\times C\times L}$, where $S$ is the number of segments and $C$ is the number of channels. The pretrained alignment module from first stage is used to map sensor data to activity labels, learning the relationship between temporal patterns and human activities while preserving inter-channel dependencies.
Input Embedding. For each sensor channel $c$, we first retrieve the combined sensor embeddings $\hat{o}_{s}^{c}$. These embeddings are then concatenated across all sensor channels, along with their corresponding statistical information (mean and variance) as prompts, to form the final input embedding $\hat{z}=\hat{o}_{s}^{1}\oplus\hat{o}_{s}^{2}\oplus\dots\oplus\hat{o}_{s}^{C}\oplus\hat{z}_{\text{stat}}$, where $\hat{z}_{\text{stat}}$ represents the statistical information of each channel, and $C$ is the total number of channels. This ensures the model integrates both temporal and statistical characteristics of the sensor data, providing a comprehensive input for HAR.
Loss Function. The input token sequence is passed through the LLM to extract a latent representation $\mathbf{H}\in\mathbb{R}^{K\times D}$, where $K$ is the number of tokens and $D$ is the embedding dimension. Due to causal masking, the model pools the final hidden state by selecting the representation of the last token, $\mathbf{h}=\mathbf{H}_{K}$, which encodes information from all preceding tokens. This pooled vector, $\mathbf{h}$, is then passed through a fully connected layer, which maps it to a vector of size $M$, where $M$ is the number of activity classes. The softmax function is applied to obtain the predicted class probabilities $\hat{y}_{i}$, which represent the likelihood of each activity class. The model then is optimized using the cross-entropy loss, defined as: $\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i}$, where $y_{i}$ is the ground truth label.
4 Datasets
Figure 3: Proportion of human activity labels in each training dataset.
To evaluate the effectiveness of SensorLLM, we utilized four publicly available HAR datasets.
USC Human Activity Dataset (USC-HAD). USC-HAD (Zhang & Sawchuk, 2012) consists of six sensor readings from body-worn 3-axis accelerometers and gyroscopes, collected from 14 subjects. The data is sampled at 100 Hz across six channels and includes 12 activity class labels. For evaluation, we use data from subjects 13 and 14 as the test set, while the remaining subjects’ data are used for training. A window size $w\in[5,200]$ is used in alignment stage, and $w=200$ is used in HAR.
UCI Human Activity Recognition Dataset (UCI-HAR). UCI-HAR (Anguita et al., 2013) includes data collected from 30 volunteers performing six activities while wearing a smartphone on their waist. The embedded accelerometer and gyroscope sensors sampled data at 50 Hz across six channels. The dataset was partitioned into 70% for training and 30% for testing. A window size $w\in[5,200]$ is used in alignment stage, and $w=128$ is used in HAR.
Physical Activity Monitoring Dataset (PAMAP2). PAMAP2 (Reiss & Stricker, 2012) includes data from nine subjects wearing IMUs on their chest, hands, and ankles. IMUs capture the acceleration, gyroscope, and magnetometer data across 27 channels and include 12 activity class labels. For our experiments, data from subjects 105 and 106 are used as the test set, with the remaining subjects’ data used for training. The sample rate is downsampled from 100 Hz to 50 Hz. A window size $w\in[5,100]$ is used in alignment stage, and $w=100$ in HAR.
Mobile Health Dataset (MHealth). MHealth (Baños et al., 2014) contains body motion and vital sign recordings from ten volunteers. Sensors were placed on the chest, right wrist, and left ankle of each subject. For our experiments, we used acceleration data from the chest, left ankle, and right lower arm, along with gyroscope data from the left ankle and right lower arm, resulting in a total of 15 channels. The data is sampled at 50 Hz and includes 12 activity class labels. Data from subjects 1, 3, and 6 is used as the test set, while the remaining subjects’ data are used for training. We use a window size $w\in[5,100]$ in alignment stage and $w=100$ in HAR.
The activities associated with each dataset are shown in Figure 3.
5 Experiments
In this section we investigate the effectiveness of SensorLLM in enabling LLM to interpret and reason across temporal-numerical (i.e., time-series) data, and in performing HAR tasks. All experiments were conducted on NVIDIA A100-80G GPUs. To ensure a fair evaluation, we used the same training and testing subjects across both stages. This setup guarantees that the test data used in Task-Aware Tuning Stage was not exposed during the alignment training in Sensor-Language Alignment Stage, ensuring that the task-aware results reflect the LLM’s true ability to learn and interpret the semantic information from the input data. We chose Chronos as the TS embedder for our experiments, as it has not been pre-trained on any sensor data, making it an ideal candidate for evaluating the robustness of our approach in learning from raw sensor inputs.
Table 1: Evaluation of Sensor Data Trend Analysis Tasks for SensorLLM and GPT-4o. The assessment includes human and GPT-4o ratings (from 1 to 5, with 5 being the highest), as well as BLEU-1, ROUGE-1, ROUGE-L, METEOR, SBERT, and SimCSE (in %). The column GPT-4o refers to the trend analysis generated by GPT-4o itself, while the row GPT-4o refers to GPT-4o’s evaluation of the generated outputs.
Metric USC-HAD UCI-HAR PAMAP2 MHealth
GPT-4o Ours GPT-4o Ours GPT-4o Ours GPT-4o Ours
BLEU-1 41.43 57.68 37.97 56.78 46.35 60.20 49.97 61.38
ROUGE-1 54.92 68.32 51.24 67.63 58.08 69.92 61.11 71.20
ROUGE-L 49.00 64.17 44.88 63.05 50.30 66.25 51.99 67.83
METEOR 30.51 45.95 26.93 45.81 37.17 52.21 38.50 51.73
SBERT 77.22 86.09 76.05 85.01 82.71 87.31 83.15 86.66
SimCSE 86.96 93.09 90.23 92.51 89.64 93.82 92.10 93.38
GPT-4o 1.67 3.11 1.61 3.2 1.90 3.77 1.69 3.69
Human 2.1 4.16 1.94 4.04 2.38 4.7 1.74 4.56
5.1 Sensor Data Understanding
Setups. For all four datasets, we trained with the following parameters in the Sensor-Language Alignment Stage: learning rate of 2e-3, 8 epochs, batch size of 4, gradient accumulation steps of 8, and maximum sequence length of 4096.
Metrics. We evaluate the effectiveness of SensorLLM in understanding and generating trend descriptions from sensor data, compared to the robust GPT-4o111gpt-4o-2024-08-06(OpenAI, 2024). This evaluation primarily focuses on assessing the success of our Sensor-Language Alignment. Using a predefined prompt A.3, GPT-4o generates answers based on the input sensor data, following our text template. To comprehensively assess performance, we employed three different evaluation methods.
  • NLP Metric Evaluation: We use standard NLP metrics including BLEU-1 (Papineni et al., 2002), ROUGE-1, ROUGE-L (Lin, 2004), and METEOR (Banerjee & Lavie, 2005) to measure surface-level similarity and n-gram overlap. Besides, we also report Sentent-BERT (SBERT) (Reimers & Gurevych, 2019) and SimCSE (Gao et al., 2021), which compute the semantic similarity between sentence embeddings of the model output and ground truth. For a fair comparison, we used SimCSE’s tokenizer across all metrics, except for SBERT, to better handle numerical content like timestamps.
  • GPT-4o Evaluation: We task GPT-4o with evaluating the generated text based on its alignment with ground truth. GPT-4o assessed the correctness of the trend descriptions and rated each model’s output on a scale of 1 to 5 (with 5 being the highest) with accompanying explanations based on predefined criteria. As one of the most advanced LLMs, GPT-4o’s semantic evaluation ensures a more accurate assessment of how well the models capture trend-related information.
  • Human Evaluation: Five human experts were tasked with evaluating the generated trend descriptions. Using the same scoring criteria as GPT-4o, they assessed the correctness of the descriptions, providing a human-centred evaluation of the model outputs.
A detailed introduction to the metrics and scoring criteria is provided in Appendix A.5. To ensure comprehensive evaluation, we randomly selected 200 data samples from each dataset for both SensorLLM and GPT-4o to generate descriptive and summary text. The evaluation results were averaged and reported. Due to the length of many descriptive texts, we selected 20 samples from each dataset with sequence lengths of 50 time steps or fewer for human evaluation.
Results. Table 1 compares SensorLLM and GPT-4o on the Sensor Data Trend Analysis tasks, showing that our model consistently outperforms GPT-4o across all metrics. BLEU-1, ROUGE-1, ROUGE-L, and METEOR primarily focus on surface-level lexical or n-gram overlaps and SBERT and SimCSE can capture factual correctness or deeper semantic similarities. Results from all metrics confirm SensorLLM’s superior ability to generate trend descriptions that are closely aligned with the ground truth. The GPT-4o evaluation offers a more nuanced assessment, examining the semantic depth, detail, and coherence of the generated descriptions. These results illustrate that SensorLLM has a stronger capacity to interpret and process sensor data, whereas even GPT-4o struggle with tasks involving numerical data and complex trend observations (Yehudai et al., 2024). Human evaluations also show higher accuracy for SensorLLM when processing shorter input sequences. Overall, these results demonstrate the effectiveness of our Sensor-Language Alignment method, validating its ability to enhance LLMs’ understanding of sensor data. Appendix A.8 shows the examples generated by GPT-4o and SensorLLM.
(a) USC-HAD
(b) UCI-HAR
(c) PAMAP2
(d) MHealth
Figure 4: Experimental results of Task-Aware Tuning Stage from five random runs. The red dashed line indicates the highest-performing result across the experiments.
5.2 Human Activity Recognition
Setups. In this section, we evaluate the performance of SensorLLM on HAR tasks. For all datasets, experiments were run five times with the following parameters: 8 training epochs, a batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 4096. We report the macro F1 score A.7 to account for class imbalance across different activity categories.
Baselines. To benchmark our model, we compare it against two sets of baselines. The first set includes state-of-the-art time-series (TS) models: Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021), NS-Transformer (Liu et al., 2022), PatchTST (Nie et al., 2023), TimesNet (Wu et al., 2023a), and iTransformer (Liu et al., 2024c). The second set consists of state-of-the-art HAR models: DeepConvLSTM (Ordóñez & Roggen, 2016), DeepConvLSTM + Attention (DeepConvLSTMAttn) (Murahari & Plötz, 2018), and Attend (Abedin et al., 2021). Detailed introduction of the baselines can be found in the Appendix A.6.
Results. As shown in Figure 4, SensorLLM outperforms the TS baselines across all datasets and surpasses the HAR baselines on USC-HAD, MHealth, and PAMAP2.
On USC-HAD, SensorLLM improves the median performance of the top TS baseline NS-Transformer by 8% and matches the top HAR baseline Attend, with a 3% gain in the best results. For UCI-HAR, SensorLLM beats NS-Transformer by 3% in median performance but trails 2% short of Attend. On PAMAP2, SensorLLM surpasses PatchTST by 4% in median performance and exceeds Attend by 1%, with a 2% improvement in best results. For MHealth, SensorLLM outperforms both PatchTST and iTransformer by 9% and improves on Attend by 6%, with a 10% gain in best results.
The fact that the TS embedder Chronos was not pre-trained on sensor data further highlights SensorLLM’s potential to understand sensor data and achieve superior performance in HAR and other related tasks.
6 Ablation Studies
In this section, we present our ablation studies to further demonstrate the effectiveness of our proposed method.
Figure 5: Experiments demonstrating the effectiveness of our Sensor-Language Alignment. The results shown are the averages from five random runs.
Impact of Alignment on HAR Performance Due to the limitations of Chronos’ pretraining sequence length, we were unable to use Chronos directly for HAR as a baseline for comparison. However, we tested the impact of bypassing Sensor-Language Alignment Stage by using Task-Aware Tuning Stage alone, directly perform HAR tasks using Chronos embeddings and the LLM. The results, shown in Figure 5, indicate that across all four datasets, SensorLLM outperforms the model that skips this step (Task-only). The performance of the Task-only model is either on par with or below the TS baselines, further highlighting the importance of our sensor-language alignment. These findings validate the effectiveness of aligning sensor-text modalities, as the integration of temporal and semantic features in the embeddings improves the LLM’s understanding of sensor data and significantly enhances HAR outcomes.
Table 2: The results for SensorLLM trained with/without text prompts. Task-only refers to conducting HAR directly bypassing sensor-language alignment. We report the F1-macro results in % with a mean and standard deviation of 5 repetitions.
Dataset Task-only (w/o prompts) Task-only (w/ prompts) SensorLLM (w/o prompts) SensorLLM (w/ prompts)
USC-HAD 43.4±$\pm 2.88$ 45.0±$\pm 1.58$ 49.6±$\pm 1.67$ 61.2±$\mathbf{\pm 3.56}$
UCI-HAR 80.0±$\pm 2.12$ 82.0±$\pm 1.58$ 89.2±$\pm 1.10$ 91.2±$\mathbf{\pm 1.48}$
PAMAP2 74.2±$\pm 2.28$ 75.4±$\pm 3.05$ 83.0±$\pm 0.71$ 86.2±$\mathbf{\pm 1.48}$
MHealth 76.6±$\pm 1.34$ 77.4±$\pm 3.13$ 86.6±$\pm 1.14$ 89.4±$\mathbf{\pm 3.85}$
Impact of Prompts To evaluate the contribution of additional textual information (statistical information for each channel’s sensor data) in Task-Aware Tuning Stage, we conducted experiments comparing SensorLLM’s performance with and without the inclusion of prompts. As shown in Table 2, the inclusion of prompts consistently improves F1-macro scores across all datasets, with a more pronounced effect when used in the full SensorLLM architecture. This indicates that the model effectively integrates both sensor and textual data, allowing it to better understand complex temporal patterns. The results highlight the advantages of leveraging multimodal inputs, which enrich the model’s representation of sensor data and lead to more accurate human activity recognition. Overall, the ability to process both sensor data and textual prompts not only boosts classification performance but also paves the way for LLMs to tackle more complex sensor-driven tasks in future applications.
Table 3: Cross-dataset experiments. We report the F1-macro results in % with a mean and standard deviation of 5 repetitions.
Stage 1 Stage 2 Results
USC-HAD UCI-HAR 91.0±$\mathbf{\pm 1.41}$
UCI-HAR USC-HAD 61.6±$\mathbf{\pm 2.07}$
Cross-Dataset Generalization To further assess the robustness of SensorLLM, we conducted cross-dataset experiments, where Sensor-Language Alignment Stage was trained on the USC-HAD dataset and Task-Aware Tuning Stage was trained on the UCI-HAR dataset (and vice versa). We chose these two datasets because they share the same sensor channels, although they differ in sample rates. As shown in Table 3, SensorLLM maintained strong performance, achieving results similar to those obtained when using the same dataset for both stages in Figure 4. These findings demonstrate that once the model effectively understands sensor data through modality alignment, retraining on new datasets for alignment becomes unnecessary. SensorLLM can directly perform downstream tasks across diverse datasets, laying the foundation for future TS-LLM models capable of generalizing across a wide range of time-series tasks.
7 Conclusion and future work
In this study, we introduced SensorLLM, a novel multimodal framework for aligning sensor data with automatically generated text. And unlike previous alignment methods which operate at the machine level, our approach aligns sensor data and text at the human perception level. SensorLLM thus can capture and interpret complex sensor data patterns—typically challenging for pre-trained LLMs—while outperforming or matching state-of-the-art models in HAR tasks. The results of cross-datasets experiments, without the need for dataset-specific alignment, also demonstrate SensorLLM’s strong generalizability and robustness. We have open-sourced our code and data generation methods, providing a valuable resource for the community and opening new avenues for aligning time-series and text data, particularly in domains with limited textual labels. Future work could explore more advanced analyses, such as chain-of-thought reasoning, sensor data generation, and zero-shot or few-shot learning, further enhancing LLMs’ potential in sensor-based tasks.
Appendix A Appendix
A.1 More related work
Deep learning in human activity recognition. Over the last decade, HAR has transitioned from hand-crafted feature extraction to deep learning models capable of automatic feature learning. Early work by Kwapisz et al. (2011) utilized machine learning techniques, such as decision trees and MLPs, to classify activities using features extracted from wearable sensor data. Later, Haresamudram et al. (2019) demonstrated that optimized feature extraction within the Activity Recognition Chain (ARC) could rival or outperform end-to-end deep learning models. Deep learning models, particularly CNNs and LSTMs, have since become dominant in HAR. Bevilacqua et al. (2019) developed a CNN-based model for HAR, while Ha & Choi (2016) introduced CNN-pf and CNN-pff architectures that apply partial and full weight sharing for better feature extraction. Other notable works include Perception-Net Kasnesis et al. (2019), which leverages 2D convolutions for multimodal sensor data, and InnoHAR (Xu et al., 2019), which combines Inception CNN and GRUs for multiscale temporal feature learning. A dual-stream network utilizing convolutional layers and LSTM units, known as ConvLSTM, was employed by Yuki et al. (2018) to analyze complex temporal hierarchies with streams handling different time lengths. The combination of attention mechanisms with recurrent networks to enhance the computation of weights for hidden state outputs has also been demonstrated by DeepConvLSTM Kasnesis et al. (2019) in capturing spatial-temporal features.
Large Language Models for Time-Series Forecasting. LLMs have achieved remarkable success in text-related tasks, and their utility has expanded into time-series forecasting. Xue & Salim (2023) presents PromptCast, which redefines time-series forecasting as a natural language generation task by transforming numerical inputs into textual prompts, enabling pre-trained language models to handle forecasting tasks with superior generalization in zero-shot settings. Gruver et al. (2023) explores encoding time-series as numerical strings, allowing LLMs like GPT-3 and LLaMA-2 to perform zero-shot forecasting, matching or surpassing the performance of specialized models, while highlighting challenges in uncertainty calibration due to model modifications like RLHF. Zhou et al. (2023a) demonstrates that pre-trained language and image models, such as a Frozen Pretrained Transformer (FPT), can be adapted for diverse time-series tasks like classification, forecasting, and anomaly detection, leveraging self-attention mechanisms to bridge the gap between different data types and achieving state-of-the-art performance across various tasks. Jin et al. (2024b) highlights the transformative potential of LLMs for time-series analysis by integrating language models with traditional analytical methods. Jin et al. (2024a) introduces a reprogramming framework that aligns time-series data with natural language processing capabilities, enabling LLMs to perform time-series forecasting without altering the core model structure. Cao et al. (2024) presents TEMPO, a generative transformer framework based on prompt tuning, which adapts pre-trained models for time-series forecasting by decomposing trends, seasonality, and residual information. Sun et al. (2024a) proposes TEST, an innovative embedding technique that integrates time-series data with LLMs through instance-wise, feature-wise, and text-prototype-aligned contrast, yielding improved or comparable results across various applications. Chang et al. (2024) develops a framework that enhances pre-trained LLMs for multivariate time-series forecasting through a two-stage fine-tuning process and a novel multi-scale temporal aggregation method, outperforming traditional models in both full-shot and few-shot scenarios. Finally, Liu et al. (2024b) introduces UniTime, a unified model that leverages language instructions and a Language-TS Transformer to handle multivariate time series across different domains, demonstrating enhanced forecasting performance and zero-shot transferability.
LLMs for Human Activity Recognition. While LLMs like ChatGPT have demonstrated remarkable performance in various NLP tasks, their effectiveness in HAR remains limited due to challenges in interpreting sensor data. These models often struggle to distinguish between activities that share similar objects, requiring more advanced prompt engineering to highlight activity-specific details. Xia et al. (2023) proposed an unsupervised approach to HAR using ChatGPT, leveraging two-stage prompts to infer activities from object sequences without manual descriptions. The method demonstrates superior performance on three benchmark datasets, marking a significant advancement in applying language models to activity recognition tasks. Similarly, Ji et al. (2024) explored LLMs for zero-shot HAR using raw IMU data, showing that GPT-4 can outperform both traditional and deep learning models in simple HAR tasks without domain-specific adaptations, highlighting LLMs’ potential in sensor-based systems.
A.2 Data Generation
In this section, we outline the process for generating text data from sensor readings. Our approach utilizes a set of predefined sentence templates, which are randomly selected to create diverse question-answer (QA) pairs. To enhance the variability, we use GPT-4o to generate synonymous versions of the sentences within the templates. Each sentence contains placeholders for relevant numerical values (e.g., timestamps, sensor readings) or textual information. We also choose a variety of synonym expressions for each trend. These values are then inserted into the placeholders to form coherent QA pairs corresponding to the sensor data inputs.
Table 4: Question templates examples.
Description: Kindly provide a detailed analysis of the trend changes observed in the {data}. Please offer a comprehensive description of how the trends in the {data} have evolved. I would appreciate a thorough explanation of the trend fluctuations that occurred within the {data}. Could you please examine the {data} in depth and explain the trend shifts observed step by step? Detail the {data}’s trend transitions. Could you kindly assess the {data} and provide a description of the trend transformations that took place step by step? Could you analyze the trends observed in the {data} over the specified period step by step? Can you dissect the {data} and explain the trend changes in a detailed manner? What trend changes can be seen in the {data}?
Summary: Could you provide a summary of the main features of the input {data} and the distribution of the trends? Please give an overview of the essential attributes of the input {data} and the spread of the trends. Describe the salient features and trend distribution within the {data}. Give a summary of the {data}’s main elements and trend apportionment. Summarize the {data}’s core features and trend dissemination. Outline the principal aspects and trend allocation of the {data}. Summarize the key features and trend distribution of the {data}. I need a summary of {data}’s main elements and their trend distributions.
Table 5: Answer templates examples.
Description: {start_time}s to {end_time}s: {trend} {start_time} seconds to {end_time} seconds: {trend} {start_time} to {end_time} seconds: {trend} {start_time}-{end_time} seconds: {trend} {start_time}-{end_time}s: {trend} {start_time}s-{end_time}s: {trend}
Summary 1: Number of {trend} trends: {num} Count of {trend} trends: {num} Number of {trend} segments: {num} Count of {trend} segments: {num}
Summary 2: The given {data_name} representing the {sensor_name} sensor readings from {start_time}s to {end_time}s. The {data_name} represents readings taken from the {sensor_name} sensor between {start_time} and {end_time} seconds. The {sensor_name} sensor readings recorded within the {start_time} to {end_time} second timeframe are presented in this {data_name}.
Summary 3: The data exhibits {trend_num} distinct trends, with a total of {change_num} changes in trend observed. Across {trend_num} trends, the data shows {change_num} occurrences of trend shifts. {trend_num} trends are present in the data, with {change_num} instances of trend changes.
Summary 4: To sum up, the data exhibited a {trend_type} trend for a cumulative period of {total_time} seconds. Overall, the data showed a {trend_type} trend over {total_time} seconds. To conclude, the trend was {trend_type} over a period of {total_time} seconds.
Summary 5: The overall trend is {overall_trend}. The primary trend detected is {overall_trend}. Looking at the big picture, the trend is {overall_trend}.
The system prompt guides the model on how to respond to the questions generated, incorporating characteristics specific to each dataset, such as sensor frequency and other dataset-specific attributes. These tailored prompts ensure the responses are aligned with the unique properties of the sensor data. Below is the template of the system prompts (instructions) used for each dataset:
  • A dialogue between a curious researcher and an AI assistant. The AI analyzes a sensor time-series dataset (N points, {sample_rate}Hz sampling rate) to answer specific questions. This interaction demonstrates the AI’s data analysis skills and the potential of human-AI collaboration in interpreting complex data.
A.3 GPT-4o Prompt for Sensor Data Trend Analysis
Table 6 outlines the system prompt used to generate trend-descriptive texts from sensor data, providing a structured format for GPT-4o to analyze and respond to specific questions about the data. This prompt serves as a standardized way for GPT-4o to examine the time-series data and produce outputs that can then be directly compared to the descriptions generated by our SensorLLM model. By using this prompt, we aim to evaluate GPT-4o’s ability to interpret numerical sensor data and generate accurate trend descriptions. The responses generated by GPT-4o are assessed against both human evaluations and NLP metrics to benchmark its performance against SensorLLM, allowing us to critically analyze the differences in how both models understand and interpret time-series data trends. This comparison helps demonstrate the effectiveness of our Sensor-Language Alignment Stage of SensorLLM.
Table 6: Prompt for GPT-4o to generate descriptive texts based on the given numerical sensor data.
Prompt A dialogue between a curious researcher and an AI assistant. The AI analyzes a sensor time-series dataset (N points, {sr}Hz sampling rate) to answer specific questions.
Please output your answer in the format like this example:
{example from ground-truth}
Now, analyze the following:
Input: {sensor_data} How trends in the given sensor data evolve?
Output:
A.4 Chronos
Chronos (Ansari et al., 2024) is an innovative framework for pretrained probabilistic time series models. It tokenizes time series data into a fixed vocabulary through scaling and quantization, then trains T5-based (Raffel et al., 2020) language models on these tokenized sequences using cross-entropy loss. Pretrained on a diverse corpus of public and synthetic datasets, Chronos models demonstrate remarkable performance, outperforming other methods on familiar datasets and showing competitive zero-shot capabilities on unseen tasks. This approach leverages data from various domains to enhance forecasting accuracy, potentially simplifying time series analysis pipelines across different fields.
Time-series Tokenization and Quantization. Chronos tokenizes real-valued time-series data through a two-step process: normalization and quantization. First, mean scaling is applied for normalization, where each observation is divided by the mean of its absolute values:
$\tilde{x}=\frac{x}{\text{mean}(\lvert x\rvert)}.$
This ensures that the data is scaled consistently across different series. After normalization, the values are quantized into discrete tokens. Quantization is performed by selecting $B$ bin centers $c_{1}<\dots<c_{B}$ along with $B-1$ edges $b_{i}$ separating them. Each real value is mapped to a token based on the quantization function $q(x)$, defined as:
$q(x)=\begin{cases}1&\text{if}\ -\infty\leq x<b_{1},\\ 2&\text{if}\ b_{1}\leq x<b_{2},\\ \vdots\\ B&\text{if}\ b_{B-1}\leq x<\infty,\end{cases}$
where each token represents a quantized range of values. This quantization ensures that the continuous time-series data can be represented as discrete tokens suitable for language models.
To further facilitate the modeling, special tokens such as PAD and EOS are added to the time-series vocabulary $V_{ts}$ to handle sequence padding and denote the end of sequences. This allows Chronos to efficiently process variable-length sequences and integrate them with language modeling tasks, treating the time-series data as token sequences.
Objective Function. In Chronos, the model uses a categorical distribution over the quantized time-series tokens from the vocabulary $V_{ts}$. The objective is to minimize the cross-entropy between the predicted categorical distribution and the ground-truth labels, including EOS tokens. The loss function for a single time series is defined as:
$\ell(\theta)=-\sum_{h=1}^{H+1}\sum_{i=1}^{|V_{ts}|}{\bf\mathbf{1}}(z_{C+h+1}=i)\log p_{\theta}(z_{C+h+1}=i\mid z_{1:C+h})$
where $C$ represents historical context of length and $H$ represents a forecast horizon, and $p_{\theta}(z_{C+h+1}=i\mid z_{1:C+h})$ is the predicted distribution. The loss is averaged over batches of time series.
This categorical output distribution offers two key advantages: (i) it requires no modification to the language model, allowing easy integration with existing architectures; and (ii) it supports learning arbitrary distributions, making it flexible for diverse time-series datasets.
A.5 Evaluation Metrics for Sensor-Language Alignment Stage
In this section, we describe the various evaluation metrics used to assess the performance of SensorLLM in generating trend descriptions from sensor data. Each metric offers a distinct perspective on model performance, ranging from surface-level textual similarity to more complex semantic alignment.
BLEU-1 (Papineni et al., 2002). BLEU (Bilingual Evaluation Understudy) is a precision-based metric commonly used to evaluate machine-generated text by comparing it to reference texts. BLEU-1 focuses on unigram (single-word) overlap, assessing the lexical similarity between the generated and reference text. While useful for measuring word-level matches, BLEU-1 does not capture deeper semantic meaning, making it most effective for surface-level alignment.
ROUGE-1 and ROUGE-L (Lin, 2004). ROUGE (Recall-Oriented Understudy for Gisting Evaluation) evaluates the recall-oriented overlap between generated text and reference text. ROUGE-1 focuses on unigram recall, similar to BLEU-1 but emphasizing how much of the reference text is captured. ROUGE-L measures the longest common subsequence, assessing both precision and recall in terms of structure and content overlap, though it does not evaluate semantic accuracy.
METEOR (Banerjee & Lavie, 2005). METEOR (Metric for Evaluation of Translation with Explicit Ordering)combines precision and recall, with additional alignment techniques such as stemming and synonym matching. Unlike BLEU and ROUGE, METEOR accounts for some degree of semantic similarity. However, its emphasis is still on word-level alignment rather than factual accuracy or meaning.
SBERT (Reimers & Gurevych, 2019). SBERT (Sentence-BERT)  [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 is a metric that generates sentence embeddings using the BERT architecture. It computes cosine similarity between embeddings of the generated and reference texts, providing a deeper assessment of semantic similarity beyond lexical matches.
SimCSE (Gao et al., 2021). SimCSE (Simple Contrastive Sentence Embedding)  [^3] https://huggingface.co/princeton-nlp/sup-simcse-roberta-large introduces a contrastive learning approach to fine-tune language models for sentence embeddings. By applying different dropout masks to the same sentence, it generates positive examples, encouraging similar embeddings for semantically identical sentences while distinguishing different ones.
GPT-4o Evaluation. In addition to the NLP metrics, we also employed GPT-4o as a human-like evaluator. Given its strong reasoning and comprehension abilities, GPT-4o was tasked with scoring the generated text based on its alignment with the ground truth. GPT-4o evaluated the correctness, completeness, and coherence of the trend descriptions and assigned a score from 1 to 5, accompanied by an explanation (see Table 7). This type of evaluation provides insights into how well the generated outputs capture the nuances of sensor data trends in a manner similar to human understanding.
Table 7: Prompt and output examples for GPT-4o in evaluating model-generated texts and ground-truth.
Prompt Please evaluate the model-generated trend descriptions against the ground truth. Rate each pair based on the degree of accuracy, using a scale from 1 to 5, where 1 represents the lowest correctness and 5 represents the highest. Deduct 1 point for minor errors in the trend description, and 2-3 points for moderate errors.
Provide your score (1-5) and a brief explanation in the format: ”score#reason” (e.g., 4#The description of trend changes slightly differs from the ground truth).
Now, please proceed to score the following:
Model: {model_output}
Human: {ground_truth}
Output:
Output example 1: 2#Significant discrepancies in segment durations and trend counts compared to ground-truth.
Output example 2: 5#The model’s description matches the human-generated text accurately.
Human Evaluation. Finally, five human experts assessed the correctness and quality of the generated trend descriptions. Following the same criteria as GPT-4o, they rated the outputs on a scale from 1 to 5, focusing on the factual accuracy and coherence of the descriptions. This manual evaluation serves as an important benchmark for the model’s performance from a human perspective, ensuring that the generated outputs are not only technically correct but also practically useful for human interpretation.
A.6 Baselines for Task-Aware Tuning Stage
In Task-Aware Tuning Stage, we compare SensorLLM against several state-of-the-art baseline models for time-series classification and human activity recognition (HAR). These models were selected for their strong performance in relevant tasks, providing a thorough benchmark for evaluating SensorLLM’s effectiveness.
Transformer (Vaswani et al., 2017). The Transformer model is a widely-used architecture in various tasks, including time-series forecasting and classification. It uses self-attention mechanisms to capture long-range dependencies in sequential data, making it highly effective for modeling complex temporal relationships.
Informer (Zhou et al., 2021). Informer is a transformer-based model designed for long sequence time-series data. It addresses key limitations of standard Transformers, such as high time complexity and memory usage, through three innovations: ProbSparse self-attention, which reduces time complexity; self-attention distilling, which enhances efficiency by focusing on dominant patterns; and a generative decoder that predicts entire sequences in a single forward pass.
NS-Transformer (Liu et al., 2022). Non-stationary Transformers (NS-Transformer) tackles the issue of over-stationarization in time-series by balancing series predictability and model capability. It introduces Series Stationarization to normalize inputs and De-stationary Attention to restore intrinsic non-stationary information into temporal dependencies.
PatchTST (Nie et al., 2023). PatchTST is a Transformer-based model for multivariate time series tasks, using subseries-level patches as input tokens and a channel-independent approach to reduce computation and improve efficiency. This design retains local semantics and allows for longer historical context, significantly improving long-term forecasting accuracy.
TimesNet (Wu et al., 2023a). TimesNet is a versatile backbone for time series analysis that transforms 1D time series into 2D tensors to better capture intraperiod and interperiod variations. This 2D transformation allows for more efficient modeling using 2D kernels. It also introduces TimesBlock to adaptively discovers multi-periodicity and extracts temporal features from transformed 2D tensors using a parameter-efficient inception block.
iTransformer (Liu et al., 2024c). iTransformer reimagines the Transformer architecture by applying attention and feed-forward networks to inverted dimensions. Time points of individual series are embedded as variate tokens, allowing the attention mechanism to capture multivariate correlations, while the feed-forward network learns nonlinear representations for each token.
DeepConvLSTM (Ordóñez & Roggen, 2016). DeepConvLSTM integrates four consecutive convolutional layers followed by two LSTM layers to effectively capture both spatial and temporal dynamics in sensor data. The final output vector is passed through a fully connected layer, and the softmax function is applied to produce activity class probabilities as the model’s final output.
DeepConvLSTMAttn (Murahari & Plötz, 2018). DeepConvLSTMAttn enhances the original DeepConvLSTM by integrating an attention mechanism to improve temporal modeling in HAR tasks. Instead of using the last LSTM hidden state for classification, the attention mechanism is applied to the first 7 hidden states, representing historical temporal context. These states are transformed through linear layers to generate attention scores, which are passed through softmax to produce weights. The weighted sum of the hidden states is combined with the last hidden state to form the final embedding for classification.
Attend (Abedin et al., 2021). The Attend model use the latent relationships between multi-channel sensor modalities and specific activities, apply data-agnostic augmentation to regularize sensor data streams, and incorporate a classification loss criterion to minimize intra-class representation differences while maximizing inter-class separability. These innovations result in more discriminative activity representations, significantly improving HAR performance.
A.7 Evaluation Metrics for Task-Aware Tuning Stage
In our evaluation, we use the F1-macro score to assess the model’s performance across datasets. F1-macro is particularly suitable for datasets with imbalanced label distributions, which is common in Human Activity Recognition (HAR) tasks where certain activities are overrepresented while others have fewer samples. Unlike the micro F1 score, which emphasizes the performance on frequent classes, F1-macro treats each class equally by calculating the F1 score independently for each class and then averaging them.
The formula for the F1-macro score is:
[…]
The precision and recall for each class are defined as:
$\text{Precision}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}}$
$\text{Recall}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FN}_{i}}$
where $\text{TP}_{i}$, $\text{FP}_{i}$, and $\text{FN}_{i}$ represent the number of true positives, false positives, and false negatives for class $i$, respectively. This metric ensures that performance is evaluated fairly across all classes, regardless of the frequency of each label, making it a robust measure for imbalanced datasets.
A.8 Sensor-Language Alignment Stage Output Examples
Tables LABEL:tab:stage1_ouput_1 and  LABEL:tab:stage1_ouput_2 present two examples of the trend analysis results generated by SensorLLM and GPT-4o based on the input sensor data. From the results, it is evident that SensorLLM outperforms GPT-4o across both shorter and medium-length sequences. This demonstrates that our approach enables LLMs to better understand numerical variations, as well as accurately compute the time duration represented by the input sequences based on their length and the given sample rate. In contrast, current large language models struggle with directly interpreting numerical data, as their tokenization methods are not well-suited for tasks such as comparing numerical values or counting (Yehudai et al., 2024).
Table 8: Example 1 of SensorLLM and GPT-4o Generated Descriptions with Ground-Truth and GPT-4o Evaluation
Sensor readings: [-9.8237, -9.4551, -10.007, -11.273, -11.258, -11.677, -11.774, -11.638, -11.195, -11.087, -10.833, -11.044, -11.393, -11.943, -12.168, -15.455, -12.967, -12.326, -12.515, -13.195, -12.634, -11.873, -12.002, -11.583, -10.859, -10.349, -9.831, -9.1622, -8.2721, -6.9299, -6.255, -5.5998]
Ground-truth: 0.0 seconds to 0.02 seconds: growing
0.02 seconds to 0.06 seconds: declining
0.06 seconds to 0.08 seconds: growing
0.08 seconds to 0.12 seconds: declining
0.12 seconds to 0.2 seconds: growing
0.2 seconds to 0.3 seconds: declining
0.3 seconds to 0.34 seconds: growing
0.34 seconds to 0.38 seconds: declining
0.38 seconds to 0.42 seconds: growing
0.42 seconds to 0.44 seconds: declining
0.44 seconds to 0.62 seconds: growing
Total growing trends: 6
Total declining trends: 5
From 0.0s to 0.62s, normalized left-ankle y-axis accelerometer data is showcased in the sensor data. Examining the data, we notice 2 clear trend characteristics, with the trend fluctuating a total of eleven times. The analysis reveals that the data’s declining inclination persisted for a total of 0.24 seconds, and a growing trend within a span of 0.38 seconds. The general trend observed is growing.
SensorLLM: 0.0 to 0.02 seconds: ascending
0.02 to 0.08 seconds: descending
0.08 to 0.1 seconds: ascending
0.1 to 0.12 seconds: descending
0.12 to 0.2 seconds: ascending
0.2 to 0.28 seconds: descending
0.28 to 0.32 seconds: ascending
0.32 to 0.36 seconds: descending
0.36 to 0.4 seconds: ascending
0.4 to 0.42 seconds: descending
0.42 to 0.62 seconds: ascending
Count of ascending segments: 6
Count of descending segments: 5
The time series data encapsulates normalized left-ankle y-axis accelerometer sensor readings from 0.0 seconds to 0.62 seconds. Two separate trends and nine trend shifts are observed in the data. The analysis reveals that the data’s descending inclination persisted for a total of 0.22 seconds, and an ascending trend for a sum of 0.40 seconds. The trend overall is ascending.
GPT-4o Evaluation: 3#Moderate errors in trend durations and shifts compared to ground truth.
GPT-4o’s output: 0.0s to 0.02s: rising
0.02s to 0.06s: falling
0.06s to 0.08s: rising
0.08s to 0.1s: falling
0.1s to 0.12s: rising
0.12s to 0.14s: falling
0.14s to 0.3s: rising
Total rising segments: 4
Total falling segments: 3
The normalized left-ankle y-axis accelerometer sensor readings recorded within the 0.0 to 0.3 second timeframe are presented in this sensor data. The input data displays three individual trends, with a comprehensive change count reaching 7. The examination reveals that the data’s falling inclination endured for an aggregate of 0.08 seconds, succeeded by a rising trend for a cumulative period of 0.22 seconds, and a steady pattern for a total of 0.00 seconds. The dominant trend is rising.
GPT-4o Evaluation: 2#Significant discrepancies in trend durations and counts between model and ground-truth.
Table 9: Example 2 of SensorLLM and GPT-4o Generated Descriptions with Ground-Truth and GPT-4o Evaluation
Sensor readings: [0.53137, 0.53137, 0.53137, 0.51176, 0.51176, 0.51176, 0.45098, 0.45098, 0.45098, 0.45098, 0.45882, 0.45882, 0.45882]
Ground-truth: 0.0 seconds to 0.04 seconds: stable
0.04 seconds to 0.06 seconds: decreasing
0.06 seconds to 0.1 seconds: stable
0.1 seconds to 0.12 seconds: decreasing
0.12 seconds to 0.18 seconds: stable
0.18 seconds to 0.2 seconds: increasing
0.2 seconds to 0.24 seconds: stable
Number of stable trends: 4
Number of decreasing trends: 2
Number of increasing trends: 1
The sensor data represents readings taken from a normalized right-lower-arm x-axis gyroscope sensor between 0.0 and 0.24 seconds. Analysis reveals three separate trends within the data, undergoing a cumulative total of seven shifts in direction. Encapsulating the outcomes, the data’s decreasing trend stretched across a total time of 0.04 seconds, came after an increasing pattern observed over 0.02 seconds, and a stable trend for 0.18 seconds in total. The dominant trend is decreasing.
SensorLLM: 0.0s to 0.04s: consistent
0.04s to 0.06s: downward
0.06s to 0.1s: consistent
0.1s to 0.12s: downward
0.12s to 0.18s: consistent
0.18s to 0.2s: upward
0.2s to 0.24s: consistent
Number of consistent segments: 4
Number of downward segments: 2
Number of upward segments: 1
The sensor data illustrates normalized right-lower-arm x-axis gyroscope sensor readings between 0.0 and 0.24 seconds. The input data displays three individual trends, with a comprehensive change count reaching 7. To encapsulate, the data’s downward trend spanned a combined duration of 0.04 seconds, and then an upward pattern for a sum of 0.02 seconds, and a consistent trend for an accumulated time of 0.18 seconds. The overarching trend is characterized as downward.
GPT-4o Evaluation: 5#The model’s description matches the ground truth accurately.
GPT-4o’s output: 0.0s to 0.04s: consistent
0.04s to 0.06s: downward
0.06s to 0.1s: consistent
0.1s to 0.12s: downward
0.12s to 0.18s: consistent
0.18s to 0.2s: upward
0.2s to 0.24s: consistent
Number of consistent segments: 4
Number of downward segments: 2
Number of upward segments: 1
The sensor data illustrates normalized right-lower-arm x-axis gyroscope sensor readings between 0.0 and 0.24 seconds. The input data displays three individual trends, with a comprehensive change count reaching 7. To encapsulate, the data’s downward trend spanned a combined duration of 0.04 seconds, and then an upward pattern for a sum of 0.02 seconds, and a consistent trend for an accumulated time of 0.18 seconds. The overarching trend is characterized as downward.
GPT-4o Evaluation: 2#Significant discrepancies in trend durations and counts compared to ground-truth.

### v2

@@
-Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim1
+Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim2
 Abstract
-In this work, we bridge the gap between wearable sensor technology and personalized AI assistants by enabling Large Language Models (LLMs) to understand time-series tasks like human activity recognition (HAR).
-Despite the strong reasoning and generalization capabilities of LLMs, leveraging them for sensor data tasks remains largely unexplored.
-This gap stems from challenges like the lack of semantic context in time-series data, computational limitations, and LLMs’ difficulty processing numerical inputs.
-To address these issues, we introduce SensorLLM, a two-stage framework to unlock LLMs’ potential for sensor data tasks.
-In the Sensor-Language Alignment Stage, we introduce special tokens for each sensor channel and automatically generate trend-descriptive text to align sensor data with textual inputs, enabling SensorLLM to capture numerical changes, channel-specific information, and sensor data of varying lengths—capabilities that existing LLMs typically struggle with, all without the need for human annotations.
-Next, in Task-Aware Tuning Stage, we refine the model for HAR classification using the frozen LLM and alignment module, achieving performance on par with or surpassing state-of-the-art models.
-We further demonstrate that SensorLLM evolves into an effective sensor learner, reasoner, and classifier through Sensor-Language Alignment, enabling it to generalize across diverse datasets for HAR tasks.
-We strongly believe our work lays the stepstone for future time-series and text alignment research, offering a path toward foundation models for sensor data.
+We introduce SensorLLM, a two-stage framework that enables Large Language Models (LLMs) to perform human activity recognition (HAR) from sensor data.
+Despite their strong reasoning and generalization capabilities, LLMs remain underutilized for motion sensor data due to the lack of semantic context in time-series, computational constraints, and challenges in processing numerical inputs.
+SensorLLM addresses these limitations through a Sensor-Language Alignment stage, where we introduce special tokens for each sensor channel and automatically generate textual trend descriptions.
+This alignment enables LLMs to capture numerical variations, channel-specific features, and data of varying durations—without requiring human annotations.
+In the subsequent Task-Aware Tuning stage, we refine the model for HAR classification, achieving performance that matches or surpasses state-of-the-art methods.
+Our results demonstrate that SensorLLM evolves into an effective sensor learner, reasoner, and classifier through Sensor-Language Alignment, generalizing across diverse HAR datasets.
+We believe this work establishes a foundation for future research on time-series and text alignment, paving the way for foundation models in sensor data analysis.
 Our codes are available at https://github.com/zechenli03/SensorLLM.
+[Uncaptioned image] SensorLLM: Aligning Large Language Models with Motion Sensors for Human Activity Recognition
+Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D.
+Salim2 1University of New South Wales, Sydney 2University of Tokyo {zechen.li, s.deldari, hao.xue1, flora.salim}@unsw.edu.au chen-linyao217@g.ecc.u-tokyo.ac.jp
 1 Introduction
-Human Activity Recognition (HAR) is a time-series (TS) classification task that involves mapping sensor data, such as accelerometer and gyroscope signals, to human activities.
-Models such as LSTM (Guan & Plötz, 2017; Hammerla et al., 2016) and DeepConvLSTM (Ordóñez & Roggen, 2016) have gained popularity in HAR due to their ability to learn high-level features.
-However, these models are typically task-specific and struggle to scale across varying sensor configurations and activity sets.
+Figure 1: SensorLLM can analyze and summarize trends in captured sensor data, facilitating human activity recognition tasks.
+Human Activity Recognition (HAR) is a time-series classification task that maps sensor signals, such as accelerometer and gyroscope data, to human activities.
+Traditional models like LSTM (Guan and Plötz, 2017; Hammerla et al., 2016) and DeepConvLSTM (Ordóñez and Roggen, 2016) learn high-level features but are task-specific and struggle to generalize across different sensor configurations and activity sets.
 In contrast, Large Language Models (LLMs) (Han et al., 2021) have shown remarkable success in integrating diverse data types (Wu et al., 2023b; Yin et al., 2023), including text and images.
-In this work, we explore how LLMs can unlock new possibilities for HAR by improving the interpretation of sensor data and offering a more scalable, flexible solution to this challenging task.
-To enable LLMs to understand effectively and process sensor data (Jin et al., 2023), there are two primary approaches.
-The first approach involves pretraining or fine-tuning LLMs on sensor data (Zhou et al., 2023b), optimizing them to handle time-series inputs directly.
-However, this method presents significant challenges, particularly related to computational resource requirements and data limitations.
-Pretraining LLMs on such data demands immense computational resources, and the limited availability of labeled sensor data, especially for rare activities, leads to class imbalance, making the training process difficult and inefficient.
-The second approach to enabling LLMs to understand sensor data focuses on using zero-shot and few-shot prompting, where sensor data is transformed into textual formats that LLMs can process without retraining (Ji et al., 2024).
-However, this introduces several challenges.
-One major issue is the numerical and temporal nature of data, which poses challenges due to how LLMs tokenize numerical inputs (Nate Gruver & Wilson, 2023).
-LLM tokenizers, designed for text, are not optimized for numerical values.
-They struggle with recognizing patterns and contextual nuances, treating consecutive values as independent tokens and ignoring temporal correlations.
-Sensor data, characterized by its continuous and extensive nature, often exceeds the maximum sequence length, resulting in truncation and potential loss of vital information.
-Another challenge lies in converting sensor data into text-like embeddings while preserving high-level time-series features, as this transformation risks losing the temporal relationships crucial for complex activity recognition (Spathis & Kawsar, 2024).
-Handling multi-channel sensor data further complicates this process, as encoding multi-dimensional information in a way that LLMs can interpret effectively is difficult cause the input for LLMs is univariate.
-Moreover, designing effective prompts that enable LLMs to interpret sensor readings, detect trends, and classify activities is complex, especially when dealing with intricate time-series patterns (Liu et al., 2023b).
-To address these challenges, we introduce a novel approach for aligning sensor data with natural language, allowing LLMs to interactively analyze sensor data through text-based commands (see Figure 1) without requiring any modifications to the LLM itself.
-A significant challenge in this context is managing the annotation bottleneck, which is further exacerbated by the inherent heterogeneity and complexity of sensor data.
-Unlike image-text pairs that can be more straightforwardly aligned, sensor data comprises multiple modalities with diverse characteristics, making the interpretation and alignment process considerably more difficult.
-Previous approaches (Jin et al., 2024a; Sun et al., 2024b) have often used condensed text prototypes for alignment.
-While these methods provide a solution, they face limitations in interpretability, as the mapping between numerical time-series data and abstract text prototypes is not always intuitive.
-Moreover, such methods often require multiple experiments to select appropriate text prototypes, adding complexity to the process.
-Figure 1: SensorLLM can analyze and summarize trends in captured sensor data, facilitating human activity recognition tasks.
-In this work, we propose a novel approach that automatically generates descriptive text from time-series data, removing the need for manual annotations.
-These descriptions are generated using statistical analyses or simple descriptions of numerical trends.
-We employ pre-defined templates to automate the generation of descriptive texts, offering a scalable, precise, and interpretable solution for aligning time-series data with text in our SensorLLM framework.
-SensorLLM introduces a two-stage framework process that has been proven effective across various domains  (Xu et al., 2024; Huang et al., 2023; Jiang et al., 2024).
-In Sensor-Language Alignment Stage, we generate automatic question-answer pairs to align sensor data with text, providing a scalable and interpretable solution.
-To preserve the temporal features of the sensor data, we use a pretrained encoder to generate sensor embeddings, which are then mapped into a space that the LLM can interpret.
-This avoids the issues caused by using a text-specific embedding model used by LLM.
-Furthermore, we introduce special tokens for each sensor channel, enabling the LLM to capture and interpret multi-channel data effectively.
-This ensures that both time-series characteristics and sensor-specific information are fully aligned with the LLM’s interpretative capabilities.
-In Task-Aware Tuning Stage, the aligned data is leveraged to address downstream tasks HAR, utilizing the LLM’s reasoning abilities while keeping its parameters frozen.
-This alignment framework enhances the capabilities of LLMs beyond their original pretraining.
-Our work demonstrates the effectiveness of SensorLLM, addressing the questions raised by Tan et al.
-(2024) regarding the broader applicability of LLMs for time-series data.
-To our knowledge, this is the first approach to incorporate sensor data into LLMs for sensor data analysis and HAR tasks.
-The key contributions of this work are as follows:
-  1. 1.
-    TS-Language Alignment: We introduce a human-intuitive approach for aligning time-series data with automatically generated text, removing the need for manual annotations.
-By leveraging text similarity metrics and human evaluations, we demonstrate that SensorLLM effectively captures both time-series patterns and channel-specific information, enabling robust multimodal understanding.
-  2. 2.
-    Task-aware Sensor-Language Model: SensorLLM achieves competitive performance across four HAR tasks, either surpassing or matching SOTA models.
-This success is driven by our modality alignment and the inclusion of task-specific prompts, both of which significantly enhance SensorLLM’s ability to interpret and classify sensor data effectively.
-  3. 3.
-    Cross-Dataset Generalization: We demonstrate that SensorLLM maintains superior performance in Task-Aware Tuning Stage, even when applied to datasets distinct from those used in alignment stage, underscoring the model’s robustness and generalizability in HAR tasks.
+Enabling LLMs to process sensor data (Jin et al., 2023) requires either (1) pretraining or fine-tuning on time-series data (Zhou et al., 2023a), which demands substantial computational resources and is hindered by limited and imbalanced labeled data, or (2) leveraging zero-shot and few-shot prompting by converting sensor data into text (Kim et al., 2024; Ji et al., 2024).
+The latter approach avoids retraining but introduces key challenges: (i) Numerical encoding issues—LLM tokenizers, designed for text, struggle with numerical values, treating consecutive numbers as independent tokens (Nate Gruver and Wilson, 2023) and failing to preserve temporal dependencies (Spathis and Kawsar, 2024).
+(ii) Sequence length constraints—sensor data often exceeds LLMs’ maximum context length, leading to truncation, information loss, and increased computational costs.
+(iii) Multi-channel complexity—LLMs process univariate inputs, making it difficult to encode multi-sensor data in a way that retains inter-channel dependencies.
+(iv) Prompt engineering challenges—designing effective prompts that enable LLMs to interpret numerical sensor readings, detect trends and classify activities remains a challenge (Liu et al., 2023b).
+To address these challenges, we propose SensorLLM, a framework that aligns sensor data with natural language, allowing LLMs to analyze sensor data through text-based interactions (see Figure 1) without modifying the LLM itself.
+A major challenge is the annotation bottleneck, further compounded by the complexity and heterogeneity of sensor data.
+Unlike image-text pairs, sensor data comprises multi-channel numerical signals with diverse characteristics, making direct interpretation and alignment difficult.
+Existing methods (Jin et al., 2024a; Sun et al., 2024a) have explored condensed text prototypes for alignment, but these approaches often lack interpretability and require extensive tuning to select suitable prototypes.
+In contrast, we propose an automatic text generation approach that aligns with human intuition by deriving descriptive trend-based text directly from time-series data using statistical analyses and predefined templates.
+This method is precise, scalable, and interpretable, eliminating the need for manual annotations while preserving essential sensor characteristics.
+SensorLLM follows a two-stage framework:
+Sensor-Language Alignment Stage.
+We automatically generate question-answer pairs to align sensor data with text while preserving temporal features using a pretrained encoder.
+The resulting embeddings are mapped into a space interpretable by the LLM, mitigating issues associated with text-specific tokenization.
+Additionally, we introduce special tokens for sensor channels, enabling LLMs to effectively capture multi-channel dependencies.
+Task-Aware Tuning Stage.
+The aligned embeddings are utilized for HAR, leveraging the LLM’s reasoning capabilities while keeping its parameters frozen.
+This design extends LLMs beyond their original training, addressing concerns raised by Tan et al.
+(2024) regarding their applicability to time-series data.
+To our knowledge, this is the first approach to integrate sensor data into LLMs for sensor data analysis and HAR tasks.
+To our knowledge, this is the first approach to integrate sensor data into LLMs for sensor data analysis and HAR tasks. The key contributions of this work are as follows:
+  • We propose a fully automated, human-intuitive approach for aligning time-series data with descriptive text, eliminating the need for manual annotations.
+Using text similarity metrics, human evaluations, and LLM-based assessments, we demonstrate that SensorLLM effectively captures temporal patterns and channel-specific features, enabling robust multimodal understanding.
+  • SensorLLM achieves competitive results across five HAR datasets, matching or surpassing state-of-the-art models.
+Experiments further validate that modality alignment and task-specific prompts significantly enhance the LLM’s ability to interpret and classify sensor data.
+  • We show that SensorLLM maintains strong performance in the Task-Aware Tuning Stage, even when applied to datasets distinct from those used during alignment, highlighting its robustness and generalizability in HAR tasks.
 2 Related Work
-In this section, we explore recent works that utilize LLMs for human-centric tasks, such as human activity recognition and mobile health sensing.
-Current approaches can be categorized based on two main aspects: (1) how they integrate numerical time-series data with LLMs, and (2) the specific end task, whether it is time-series forecasting, classification, etc.
-Zhang et al.
-(2024) provides a comprehensive review of the existing methodologies that leverage LLM capabilities in processing time-series data.
-Several works investigated using LLM for numerical data by treating time-series signals as raw text.
-They used the same tokenization and embedding models as text such as PromptCast  (Xue & Salim, 2023) and LLMTime (Gruver et al., 2024).
-However, despite LLM’s versatility, they are inherently designed for processing sequential text data and not for heterogeneous time-series numerical data.
-The challenge lies in enabling LLMs to interpret and learn patterns within multimodal time-series data effectively.
-As highlighted by Spathis & Kawsar (2024), one significant barrier is the absence of suitable tokenizers for numerical data.
-LLMs typically employ tokenizers that segment text into smaller units, but these tokenizers are not optimized for representing numerical values.
-Consequently, they may struggle to recognise repetitive patterns and context, often treating consecutive values as independent tokens and overlooking the crucial temporal relationships inherent in time-series data.
-To extract more meaningful insights from numerical data before feeding them to LLM, several works proposed the use of time-series encoders and tried to align the embedding space with text embedding models.
-Liu et al.
-(2024a) aligns ECG data with corresponding textual descriptions using contrastive learning techniques.
-Similar approaches have been investigated for aligning motion/IoT sensors with text (Zhou et al., 2023c; Haresamudram et al., 2024; Moon et al., 2023; Xia et al., 2024).
-Please refer to Appendix A.1 for a more in-depth review of existing works.
+In this section, we discuss recent developments in leveraging LLMs for time-series data, specifically focusing on two categories: (1) LLMs for time series as text and (2) Multimodal Large Language Models (MLLMs) for sensor data.
+A broader overview of other related works, including deep learning approaches to HAR and additional LLM-based forecasting methods, is provided in Appendix A.1.
+LLMs for Time Series as Text.
+While LLMs excel in processing natural language, applying them directly to time-series data poses unique challenges (Spathis and Kawsar, 2024).
+Certain methods address this by treating time-series signals as raw text, using the same tokenization as natural language.
+Notable examples include PromptCast (Xue and Salim, 2023), which transforms numeric inputs into textual prompts for zero-shot forecasting, and LLMTime (Gruver et al., 2024), which encodes time-series as numerical strings for GPT-like models.
+However, due to the lack of specialized tokenizers for numeric sequences, LLMs may fail to capture crucial temporal dependencies and repetitive patterns (Spathis and Kawsar, 2024).
+To mitigate these issues, several works employ time-series encoders before mapping the resulting embeddings to language model spaces (Liu et al., 2024a; Zhou et al., 2023c; Xia et al., 2024), thus aligning sensor embeddings with textual embeddings in a contrastive or supervised manner.
+MLLMs for Sensor Data.
+Extending LLMs to non-textual domains has gained traction, particularly through MLLMs that accept inputs beyond text, such as images or speech.
+For sensor data, the challenge lies in representing continuous signals effectively.
+Yoon et al.
+(2024) propose to ground MLLMs with sensor data via visual prompting.
+Sensor signals are first visualized as images, guiding the MLLM to analyze the visualized sensor traces alongside task descriptions, which also lower token costs compared to raw-text baselines.
+Similarly, Moon et al.
+(2023) introduce IMU2CLIP, which aligns inertial measurement unit streams with text and video in a joint representation space.
+This approach enables wearable AI applications like motion-based media search and LM-based multimodal reasoning, showcasing how sensor data can be integrated into broader multimodal frameworks.
 3 Methods
-In this work, we propose a novel approach to align sensor data with descriptive text using automatically generated question-answer pairs.
-The aim of this work is to develop a multimodal foundational model with reasoning capabilities to address challenges in analyzing wearable sensor data.
-As shown in Figure 2, our SensorLLM comprises three core components: 1) a pretrained LLM, 2) a pretrained TS embedder, and 3) an alignment module MLP.
-The three core modules cooperate in a two-stage framework.
+In this work, we propose SensorLLM, a framework that aligns sensor data with descriptive text through automatically generated question–answer pairs.
+Our goal is to develop a multimodal model with reasoning capabilities for analyzing wearable sensor data.
+As shown in Figure 2, SensorLLM comprises three core components: (1) a pretrained LLM, (2) a pretrained time-series (TS) embedder, and (3) an alignment module MLP, all operating in a two-stage framework.
 In the Sensor-Language Alignment Stage, a generative model aligns sensor readings with text based on user instructions (Liu et al., 2023a) and questions.
-In Task-Aware Tuning Stage, a classification model is used to perform HAR and maintain an end-to-end alignment.
-Notably, only the parameters of the MLP in both stages, as well as the classifier in the second stage, are updated during training, while the backbone LLM and TS embedder remain frozen.
-Consequently, only 5.67% of the parameters in stage 1 and 6% in stage 2 are trainable, ensuring a highly efficient and lightweight training process.
+In the Task-Aware Tuning Stage, a classifier is added on top of the LLM to perform HAR.
+Crucially, only the MLP in both stages and the classifier in the Task-Aware Tuning Stage are trainable, while the backbone LLM and TS embedder remain frozen.
+This design yields a highly efficient and lightweight training process, requiring only 5.67% (535.9M) of the parameters to be trainable in the alignment stage and 0.12% (10.5M) in the tuning stage.
 Figure 2: Our proposed SensorLLM framework: (a) Sensor-Language Alignment Stage, where a generative model aligns sensor readings with automatically generated text; (b) Task-Aware Tuning Stage, where a classification model leverages the aligned modalities to perform HAR.
 3.1 Sensor-Text Data Generation
-A key challenge in aligning time-series data with text for LLM-based tasks is effectively bridging these two modalities.
-Unlike other multimodal tasks, time-series data typically lacks rich semantic textual information beyond class labels, making manual annotation impractical  (Deldari et al., 2024; Haresamudram et al., 2024).
-While prior works have used predefined text prototypes for alignment (Sun et al., 2024a; Jin et al., 2024a), we aim to guide the model toward a more human-intended understanding of the data.
-We argue that time-series data inherently contains semantic information that can be expressed through descriptive text, from simple numerical trends to more complex insights derived from statistical analysis.
-Our approach focuses on automatically generating descriptive text based on observed trends and fluctuations in the input data.
-To achieve this, we use a variety of predefined templates to generate diverse question-answer (QA) pairs that describe trend changes in the sensor data, ensuring both accuracy and scalability, while avoiding the potential errors of manual annotation.
-These templates are randomly combined to produce diverse QA pairs.
+Aligning time-series data with text for LLM-based tasks is challenging due to the lack of rich semantic labels beyond class annotations, making manual annotation impractical (Deldari et al., 2024; Haresamudram et al., 2024).
+While prior works rely on predefined text prototypes (Sun et al., 2024b; Jin et al., 2024a), we aim for a more human-intuitive representation of sensor data.
+We argue that time-series data inherently contains semantic patterns that can be expressed through descriptive text, from simple numerical trends to statistical insights.
+To achieve this, we automatically generate descriptive text by analyzing observed trends and fluctuations in the data.
+Using predefined templates, we construct diverse question-answer (QA) pairs that capture trend changes while ensuring accuracy and scalability.
+These templates (Appendix A.2) are randomly combined to enhance diversity.
 For example:
@@
     To sum up, the data exhibited a <T> trend for a cumulative period of <$t_{t}$> seconds.
-where T and S represent placeholders for a specific trend and sensor type, respectively, and t correspond to numerical values.
-These automatically generated QA pairs offer a scalable and interpretable method for aligning sensor data and text in the latent space, facilitating more effective understanding and interaction with LLMs.
-See Appendix A.2 for detailed templates.
+where T and S denote specific trends and sensor types, and t corresponds to numerical values.
 3.2 Sensor-Language Alignment
-As shown in Figure 2 (a), Sensor-Language Alignment stage employs a generative model designed to create multi-modal sentences that combine single-channel sensor readings with textual descriptions.
-The sensor data is represented as a matrix $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ indicates the number of sensor channels, and $T$ represents the total number of time steps in the data sequence.
-Each individual channel’s data, denoted as $\mathbf{X}^{c}$ for channel $c$, is processed independently to maintain the uniqueness of information captured by each channel.
-The data is then divided into non-overlapping segments, $\mathbf{X}_{S}^{c}$, where $S$ represents the total number of segments.
-Each segment $x_{s}$ is assigned a random length $l$ within a predefined range.
-The variation in the length of segments allows the model to effectively learn from a variety of temporal patterns and trend variations across different segments.
-By learning from segments with different lengths, the generative model becomes more adept at handling the diverse characteristics of sensor data, ensuring that both long-term trends and short-term fluctuations are adequately represented in the generated multi-modal sentences.
-We employ the Chronos (Ansari et al., 2024) as TS embedder, which generates the segment embedding $\hat{x}_{s}\in\mathbb{R}^{(l+1)\times d_{ts}}$, where $d_{ts}$ is the feature dimension, and the $(l+1)$ accounts for the [EOS] token appended during Chronos tokenization.
-Unlike the original Chronos implementation, which uses mean normalization A.4, our approach utilizes instance normalization $\tilde{x}_{s}=\frac{x_{s}-\text{mean}(x_{s})}{\text{std}(x_{s})}$ for improved performance.
-We use Llama3-8b (Touvron et al., 2023) as our LLM backbone.
+As shown in Figure 2 (a), the Sensor-Language Alignment stage employs a generative model to create multimodal sentences that combine single-channel sensor readings with textual descriptions.
+The sensor data is represented as a matrix $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ is the number of sensor channels and $T$ is the sequence length.
+Each channel’s data, denoted as $\mathbf{X}^{c}$ for channel $c$, is processed independently to preserve its unique characteristics.
+The data is then divided into non-overlapping segments, $\mathbf{X}_{S}^{c}$, where $S$ is the total number of segments.
+Each segment $x_{s}$ is assigned a random length $l$ within a predefined range, allowing the model to learn from varying temporal patterns and trend variations.
+This segmentation strategy ensures that both long-term trends and short-term fluctuations are effectively captured in the generated multimodal sentences.
+We use Chronos (Ansari et al., 2024) as the TS embedder, generating segment embeddings $\hat{x}_{s}\in\mathbb{R}^{(l+1)\times d_{ts}}$, where $d_{ts}$ is the feature dimension, and $(l+1)$ accounts for the [EOS] token appended during Chronos tokenization.
+Prior to inputting sensor segments into Chronos (Appendix A.3), we apply instance normalization $\tilde{x}_{s}=\frac{x_{s}-\text{mean}(x_{s})}{\text{std}(x_{s})}$ to standardize the data.
+Llama3-8B (Touvron et al., 2023) serves as our LLM backbone.
 Alignment Module.
-To transform the time-series embeddings $\hat{x}{s}$ into text embeddings $\hat{a}_{s}\in\mathbb{R}^{(l+1)\times D}$ for further downstream tasks, we introduce an alignment projection module.
-This module is implemented as a multi-layer perceptron (MLP), which first maps the sensor embeddings to an intermediate latent space $d_{m}$ and then projects them to the final target dimension $D$.
-Formally, the projection can be written as: $\hat{a}_{s}=\mathbf{W}_{2}\cdot\sigma(\mathbf{W}_{1}\hat{x}_{s}+\mathbf{b}_{1})+\mathbf{b}_{2}$ where $\mathbf{W}1\in\mathbb{R}^{d_{m}\times d_{ts}}$ and $\mathbf{W}2\in\mathbb{R}^{D\times d_{m}}$ are learnable weights, $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are biases, and $\sigma$ is GELU activation function (Hendrycks & Gimpel, 2016).
-This projection ensures that the transformed embeddings $\hat{a}_{s}$ are semantically aligned with the text embedding space, making them suitable for subsequent tasks such as generation or classification.
+To transform TS embeddings $\hat{x}_{s}$ into text-aligned embeddings $\hat{a}_{s}\in\mathbb{R}^{(l+1)\times D}$ for downstream tasks, we introduce an alignment projection module.
+This module, implemented as a multi-layer perceptron (MLP), first maps sensor embeddings to an intermediate space of dimension $d_{m}$ and then projects them to the target dimension $D$.
+Formally,
+$\hat{a}_{s}=\mathbf{W}_{2}\cdot\sigma(\mathbf{W}_{1}\hat{x}_{s}+\mathbf{b}_{1})+\mathbf{b}_{2},$ (1)
+where $\mathbf{W}_{1}\in\mathbb{R}^{d_{m}\times d_{ts}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{D\times d_{m}}$ are learnable weights, $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are biases, and $\sigma$ is the GELU activation function (Hendrycks and Gimpel, 2016).
+This projection ensures that the transformed embeddings $\hat{a}_{s}$ are semantically aligned with the text embedding space, making them suitable for tasks such as text generation and classification.
 Input Embedding.
-To enable the LLM to process sensor channel information and align sensor features with text tokens, we introduce two special tokens for each channel (e.g.,<x_acc_start> and <x_acc_end> for the x-axis accelerometer), extending the LLM’s embedding matrix from $\mathbf{E}\in\mathbb{R}^{V\times D}$ to $\mathbf{E}\in\mathbb{R}^{V^{\prime}\times D}$, where $V^{\prime}=V+2c$ and $V$ is the vocabulary size, and $c$ is the number of channels.
-Then the special token embeddings are concatenated with the aligned sensor embeddings.
-Finally, the combined sensor embeddings $\hat{o}_{s}\in\mathbb{R}^{(l+3)\times D}$ are concatenated with instruction and question embeddings to form the final input $\hat{z}\in\mathbb{R}^{k\times D}$, where $k$ is the total number of tokens.
+To integrate sensor data into the LLM, we introduce two special tokens per sensor channel (e.g., <x_acc_start> and <x_acc_end> for the x-axis accelerometer), extending the LLM’s embedding matrix from $\mathbf{E}\in\mathbb{R}^{V\times D}$ to $\mathbf{E}\in\mathbb{R}^{V^{\prime}\times D}$, where $V^{\prime}=V+2c$, with $V$ as the vocabulary size and $c$ as the number of channels.
+These special token embeddings are concatenated with the aligned sensor embeddings.
+The final combined sensor representation $\hat{o}_{s}\in\mathbb{R}^{(l+3)\times D}$ is then concatenated with instruction and question embeddings to form the full input sequence $\hat{z}\in\mathbb{R}^{k\times D}$, where $k$ is the total number of tokens.
 Loss Function.
-Our SensorLLM processes a mixed sequence of tokens, $\mathbf{Z}_{s}=\{z_{s}^{i}\}_{i=1}^{K}$, consisting of both sensor and text embeddings, and generates an output sequence, $\mathbf{Z}_{t}=\{z_{t}^{i}\}_{i=1}^{N}$, where $z_{s}^{i}$ and $z_{t}^{i}\in V^{\prime}$, and $K$ and $N$ represent the number of input and output tokens, respectively.
-The generative model is trained using a causal language modeling objective, where the next token in the sequence is predicted based on all preceding tokens.
-To optimize this, we minimize the negative log-likelihood of the generated text tokens ($\mathcal{L}_{gen}=-\sum_{i=0}^{N-1}\log P(z_{t}^{i}|Z_{t}^{<i},z_{s})$).
-The loss is computed solely for output tokens, excluding those corresponding to instructions, questions, channel special tokens, and sensor data.
-This ensures that SensorLLM effectively integrates sensor and text embeddings, enabling the generation of coherent and contextually appropriate responses.
+SensorLLM processes an input sequence $\mathbf{Z}_{s}=\{z_{s}^{i}\}_{i=1}^{K}$ consisting of sensor and text embeddings and generates an output sequence $\mathbf{Z}_{t}=\{z_{t}^{i}\}_{i=1}^{N}$, where $z_{s}^{i},z_{t}^{i}\in V^{\prime}$, and $K$ and $N$ represent the number of input and output tokens, respectively.
+The model is trained using a causal language modeling objective, predicting the next token based on previous ones.
+The optimization minimizes the negative log-likelihood:
+$\mathcal{L}_{gen}=-\sum_{i=0}^{N-1}\log P(z_{t}^{i}|Z_{t}^{<i},z_{s}).$ (2)
+Loss is computed only on generated tokens, ensuring SensorLLM effectively integrates sensor and text embeddings to produce coherent, contextually appropriate responses.
 3.3 Task-Aware Tuning
-As shown in Figure 2 (b), the Task-Aware Tuning stage employs a classification model to align the multimodal sensor-text embeddings with human activity labels.
-This stage further integrates multi-channel sensor readings with activity labels for HAR.
-The input sensor data $\mathbf{X}$ is divided into segments of window size $L$ with 50% overlap (Li et al., 2018), forming segments $\mathbf{X}_{S}\in\mathbb{R}^{S\times C\times L}$, where $S$ is the number of segments and $C$ is the number of channels.
-The pretrained alignment module from first stage is used to map sensor data to activity labels, learning the relationship between temporal patterns and human activities while preserving inter-channel dependencies.
+As shown in Figure 2 (b), the Task-Aware Tuning stage refines the multimodal sensor-text embeddings for HAR.
+This stage integrates multi-channel sensor readings with activity labels, aligning temporal patterns with human activities.
+The input sensor data $\mathbf{X}$ is segmented into overlapping windows of size $L$ with a 50% overlap (Li et al., 2018), forming segments $\mathbf{X}_{S}\in\mathbb{R}^{S\times C\times L}$, where $S$ is the number of segments and $C$ is the number of channels.
+The pretrained alignment module from the first stage maps sensor data to activity labels, preserving inter-channel dependencies while learning activity-related patterns.
 Input Embedding.
-For each sensor channel $c$, we first retrieve the combined sensor embeddings $\hat{o}_{s}^{c}$.
-These embeddings are then concatenated across all sensor channels, along with their corresponding statistical information (mean and variance) as prompts, to form the final input embedding $\hat{z}=\hat{o}_{s}^{1}\oplus\hat{o}_{s}^{2}\oplus\dots\oplus\hat{o}_{s}^{C}\oplus\hat{z}_{\text{stat}}$, where $\hat{z}_{\text{stat}}$ represents the statistical information of each channel, and $C$ is the total number of channels.
-This ensures the model integrates both temporal and statistical characteristics of the sensor data, providing a comprehensive input for HAR.
+For each sensor channel $c$, we retrieve its aligned sensor embeddings $\hat{o}_{s}^{c}$.
+These embeddings are then concatenated across all channels, along with their corresponding statistical features (mean and variance), to form the final input embedding:
+$\hat{z}=\hat{o}_{s}^{1}\oplus\hat{o}_{s}^{2}\oplus\dots\oplus\hat{o}_{s}^{C}\oplus\hat{z}_{\text{stat}},$ (3)
+where $\hat{z}_{\text{stat}}$ represents the statistical information, and $C$ is the number of channels. This ensures the model integrates both temporal and statistical characteristics for HAR.
 Loss Function.
-The input token sequence is passed through the LLM to extract a latent representation $\mathbf{H}\in\mathbb{R}^{K\times D}$, where $K$ is the number of tokens and $D$ is the embedding dimension.
-Due to causal masking, the model pools the final hidden state by selecting the representation of the last token, $\mathbf{h}=\mathbf{H}_{K}$, which encodes information from all preceding tokens.
-This pooled vector, $\mathbf{h}$, is then passed through a fully connected layer, which maps it to a vector of size $M$, where $M$ is the number of activity classes.
-The softmax function is applied to obtain the predicted class probabilities $\hat{y}_{i}$, which represent the likelihood of each activity class.
-The model then is optimized using the cross-entropy loss, defined as: $\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i}$, where $y_{i}$ is the ground truth label.
-4 Datasets
-Figure 3: Proportion of human activity labels in each training dataset.
-To evaluate the effectiveness of SensorLLM, we utilized four publicly available HAR datasets.
-USC Human Activity Dataset (USC-HAD).
-USC-HAD (Zhang & Sawchuk, 2012) consists of six sensor readings from body-worn 3-axis accelerometers and gyroscopes, collected from 14 subjects.
-The data is sampled at 100 Hz across six channels and includes 12 activity class labels.
-For evaluation, we use data from subjects 13 and 14 as the test set, while the remaining subjects’ data are used for training.
-A window size $w\in[5,200]$ is used in alignment stage, and $w=200$ is used in HAR.
-UCI Human Activity Recognition Dataset (UCI-HAR).
-UCI-HAR (Anguita et al., 2013) includes data collected from 30 volunteers performing six activities while wearing a smartphone on their waist.
-The embedded accelerometer and gyroscope sensors sampled data at 50 Hz across six channels.
-The dataset was partitioned into 70% for training and 30% for testing.
-A window size $w\in[5,200]$ is used in alignment stage, and $w=128$ is used in HAR.
-Physical Activity Monitoring Dataset (PAMAP2).
-PAMAP2 (Reiss & Stricker, 2012) includes data from nine subjects wearing IMUs on their chest, hands, and ankles.
-IMUs capture the acceleration, gyroscope, and magnetometer data across 27 channels and include 12 activity class labels.
-For our experiments, data from subjects 105 and 106 are used as the test set, with the remaining subjects’ data used for training.
-The sample rate is downsampled from 100 Hz to 50 Hz.
-A window size $w\in[5,100]$ is used in alignment stage, and $w=100$ in HAR.
-Mobile Health Dataset (MHealth).
-MHealth (Baños et al., 2014) contains body motion and vital sign recordings from ten volunteers.
-Sensors were placed on the chest, right wrist, and left ankle of each subject.
-For our experiments, we used acceleration data from the chest, left ankle, and right lower arm, along with gyroscope data from the left ankle and right lower arm, resulting in a total of 15 channels.
-The data is sampled at 50 Hz and includes 12 activity class labels.
-Data from subjects 1, 3, and 6 is used as the test set, while the remaining subjects’ data are used for training.
-We use a window size $w\in[5,100]$ in alignment stage and $w=100$ in HAR.
-The activities associated with each dataset are shown in Figure 3.
+The input token sequence is processed by the LLM, yielding a latent representation $\mathbf{H}\in\mathbb{R}^{K\times D}$, where $K$ is the number of tokens and $D$ is the embedding dimension.
+Due to causal masking, we extract the final hidden state, $\mathbf{h}=\mathbf{H}_{K}$, which encodes all preceding token information.
+This pooled vector is passed through a fully connected layer to produce a prediction vector of size $M$, where $M$ is the number of activity classes.
+The final class probabilities $\hat{y}_{i}$ are obtained via the softmax function, and the model is optimized using cross-entropy loss:
+$\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i},$ (4)
+where $y_{i}$ is the ground truth label.
 4 Experiments
-In this section we investigate the effectiveness of SensorLLM in enabling LLM to interpret and reason across temporal-numerical (i.e., time-series) data, and in performing HAR tasks.
-All experiments were conducted on NVIDIA A100-80G GPUs.
-To ensure a fair evaluation, we used the same training and testing subjects across both stages.
-This setup guarantees that the test data used in Task-Aware Tuning Stage was not exposed during the alignment training in Sensor-Language Alignment Stage, ensuring that the task-aware results reflect the LLM’s true ability to learn and interpret the semantic information from the input data.
-We chose Chronos as the TS embedder for our experiments, as it has not been pre-trained on any sensor data, making it an ideal candidate for evaluating the robustness of our approach in learning from raw sensor inputs.
+In this section, we evaluate SensorLLM in enabling LLMs to interpret, reason about, and classify sensor data for HAR tasks.
+All experiments are conducted on NVIDIA A100-80G GPUs.
+To assess the LLM’s ability to learn and generalize from raw sensor inputs, we ensure that the same training and testing subjects are used in both the Sensor-Language Alignment and Task-Aware Tuning stages.
+This guarantees that test data in the second stage remains unseen during alignment, ensuring a fair evaluation of generalization.
+We select Chronos as the TS embedder because it has not been pre-trained on motion sensor data, making it an ideal candidate for evaluating our approach’s robustness in learning directly from unprocessed sensor signals.
+4.1 Datasets
+To evaluate the effectiveness and generalizability of SensorLLM, we conduct experiments on five publicly available HAR datasets: USC-HAD (Zhang and Sawchuk, 2012), UCI-HAR (Anguita et al., 2013), PAMAP2 (Reiss and Stricker, 2012), MHealth (Baños et al., 2014), and CAPTURE-24 (Chan et al., 2024).
+These datasets differ in sensor placement, sampling rates, channel configurations, and activity types, covering both controlled laboratory conditions and free-living environments.
+Additionally, they vary in scale, with USC-HAD, UCI-HAR, PAMAP2, and MHealth collected from a limited number of subjects, whereas CAPTURE-24 features a large-scale dataset with sensor recordings from 151 participants in real-world settings.
+Full dataset details, including subject count, sensor configurations, data splits, activity classes, preprocessing steps, and windowing strategies, are provided in Appendix A.6.
+Metric USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
+GPT-4o Ours GPT-4o Ours GPT-4o Ours GPT-4o Ours GPT-4o Ours
+BLEU-1 41.43 57.68 37.97 56.78 46.35 60.20 49.97 61.38 46.58 57.10
+ROUGE-1 54.92 68.32 51.24 67.63 58.08 69.92 61.11 71.20 58.21 68.11
+ROUGE-L 49.00 64.17 44.88 63.05 50.30 66.25 51.99 67.83 48.88 60.90
+METEOR 30.51 45.95 26.93 45.81 37.17 52.21 38.50 51.73 31.16 40.51
+SBERT 77.22 86.09 76.05 85.01 82.71 87.31 83.15 86.66 83.11 84.83
+SimCSE 86.96 93.09 90.23 92.51 89.64 93.82 92.10 93.38 90.10 92.20
+GPT-4o 1.67 3.11 1.61 3.20 1.90 3.77 1.69 3.69 1.70 2.32
+Human 2.10 4.16 1.94 4.04 2.38 4.70 1.74 4.56 2.30 3.10
 Table 1: Evaluation of Sensor Data Trend Analysis Tasks for SensorLLM and GPT-4o.
@@
 The column GPT-4o refers to the trend analysis generated by GPT-4o itself, while the row GPT-4o refers to GPT-4o’s evaluation of the generated outputs.
-Metric USC-HAD UCI-HAR PAMAP2 MHealth
-GPT-4o Ours GPT-4o Ours GPT-4o Ours GPT-4o Ours
-BLEU-1 41.43 57.68 37.97 56.78 46.35 60.20 49.97 61.38
-ROUGE-1 54.92 68.32 51.24 67.63 58.08 69.92 61.11 71.20
-ROUGE-L 49.00 64.17 44.88 63.05 50.30 66.25 51.99 67.83
-METEOR 30.51 45.95 26.93 45.81 37.17 52.21 38.50 51.73
-SBERT 77.22 86.09 76.05 85.01 82.71 87.31 83.15 86.66
-SimCSE 86.96 93.09 90.23 92.51 89.64 93.82 92.10 93.38
-GPT-4o 1.67 3.11 1.61 3.2 1.90 3.77 1.69 3.69
-Human 2.1 4.16 1.94 4.04 2.38 4.7 1.74 4.56
 4.2 Sensor Data Understanding
-Setups.
-For all four datasets, we trained with the following parameters in the Sensor-Language Alignment Stage: learning rate of 2e-3, 8 epochs, batch size of 4, gradient accumulation steps of 8, and maximum sequence length of 4096.
-Metrics.
-We evaluate the effectiveness of SensorLLM in understanding and generating trend descriptions from sensor data, compared to the robust GPT-4o111gpt-4o-2024-08-06(OpenAI, 2024).
-This evaluation primarily focuses on assessing the success of our Sensor-Language Alignment.
-Using a predefined prompt A.3, GPT-4o generates answers based on the input sensor data, following our text template.
-To comprehensively assess performance, we employed three different evaluation methods.
-  • NLP Metric Evaluation: We use standard NLP metrics including BLEU-1 (Papineni et al., 2002), ROUGE-1, ROUGE-L (Lin, 2004), and METEOR (Banerjee & Lavie, 2005) to measure surface-level similarity and n-gram overlap.
-Besides, we also report Sentent-BERT (SBERT) (Reimers & Gurevych, 2019) and SimCSE (Gao et al., 2021), which compute the semantic similarity between sentence embeddings of the model output and ground truth.
-For a fair comparison, we used SimCSE’s tokenizer across all metrics, except for SBERT, to better handle numerical content like timestamps.
-  • GPT-4o Evaluation: We task GPT-4o with evaluating the generated text based on its alignment with ground truth.
-GPT-4o assessed the correctness of the trend descriptions and rated each model’s output on a scale of 1 to 5 (with 5 being the highest) with accompanying explanations based on predefined criteria.
-As one of the most advanced LLMs, GPT-4o’s semantic evaluation ensures a more accurate assessment of how well the models capture trend-related information.
-  • Human Evaluation: Five human experts were tasked with evaluating the generated trend descriptions.
-Using the same scoring criteria as GPT-4o, they assessed the correctness of the descriptions, providing a human-centred evaluation of the model outputs.
-A detailed introduction to the metrics and scoring criteria is provided in Appendix A.5.
-To ensure comprehensive evaluation, we randomly selected 200 data samples from each dataset for both SensorLLM and GPT-4o to generate descriptive and summary text.
-The evaluation results were averaged and reported.
-Due to the length of many descriptive texts, we selected 20 samples from each dataset with sequence lengths of 50 time steps or fewer for human evaluation.
+Setup.
+We train SensorLLM on all five datasets using the same following parameters in the Sensor-Language Alignment Stage: a learning rate of 2e-3, 8 epochs, batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 8192 for CAPTURE-24 and 4096 for the other datasets.
+Evaluation Metrics.
+We assess SensorLLM’s ability to generate trend descriptions from sensor data, comparing it with the advanced GPT-4o 111gpt-4o-2024-08-06(OpenAI, 2024) to evaluate Sensor-Language Alignment.
+GPT-4o generates responses based on a predefined prompt (Appendix A.4), following our text template.
+We employ three evaluation methods:
+  • NLP Metrics.
+We measure surface-level similarity and n-gram overlap using BLEU-1 (Papineni et al., 2002), ROUGE-1, ROUGE-L (Lin, 2004), and METEOR (Banerjee and Lavie, 2005).
+We also use SBERT (Reimers and Gurevych, 2019) and SimCSE (Gao et al., 2021) to evaluate semantic similarity between model outputs and ground truth.
+  • GPT-4o Evaluation.
+GPT-4o rates the generated trend descriptions on a scale of 1 to 5 (with 5 being the highest) by comparing each output to ground truth and providing explanatory feedback.
+As an advanced LLM, its evaluation ensures a semantic assessment of trend comprehension.
+  • Human Evaluation.
+Five time-series experts (PhD students, postdocs, and academics) score accuracy and quality using the same criteria as GPT-4o, providing a human-centered perspective on the model’s outputs.
+Appendix A.5 details all metrics and scoring criteria.
+We randomly sample 200 instances per dataset for both SensorLLM and GPT-4o, then average the results for comparison.
+Because reading and comparing lengthy sequences is difficult for human annotators, we conduct human evaluation on 20 shorter sequences per dataset (each containing at most 50 time steps).
 Results.
@@
 BLEU-1, ROUGE-1, ROUGE-L, and METEOR primarily focus on surface-level lexical or n-gram overlaps and SBERT and SimCSE can capture factual correctness or deeper semantic similarities.
-Results from all metrics confirm SensorLLM’s superior ability to generate trend descriptions that are closely aligned with the ground truth.
-The GPT-4o evaluation offers a more nuanced assessment, examining the semantic depth, detail, and coherence of the generated descriptions.
-These results illustrate that SensorLLM has a stronger capacity to interpret and process sensor data, whereas even GPT-4o struggle with tasks involving numerical data and complex trend observations (Yehudai et al., 2024).
-Human evaluations also show higher accuracy for SensorLLM when processing shorter input sequences.
-Overall, these results demonstrate the effectiveness of our Sensor-Language Alignment method, validating its ability to enhance LLMs’ understanding of sensor data.
-Appendix A.8 shows the examples generated by GPT-4o and SensorLLM.
-(a) USC-HAD
-(b) UCI-HAR
-(c) PAMAP2
-(d) MHealth
-Figure 4: Experimental results of Task-Aware Tuning Stage from five random runs. The red dashed line indicates the highest-performing result across the experiments.
+Across all metrics, SensorLLM generates trend descriptions more closely aligned with the ground truth.
+GPT-4o evaluations further highlight SensorLLM’s superior ability to capture trend details and coherence, whereas GPT-4o struggles with complex numerical data and trend observations (Yehudai et al., 2024).
+Human evaluation also favors SensorLLM, particularly for shorter sequences.
+CAPTURE-24 results are weaker compared to other datasets, likely due to its longer sequences being trained with the same parameters.
+Overall, these findings validate the effectiveness of our Sensor-Language Alignment method in enhancing LLMs’ ability to interpret complex numerical sequence.
+Appendix A.9 provides qualitative examples of outputs from both models.
+Method USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
+PatchTST 45.2±$\pm 1.48$ 86.8±$\pm 0.84$ 82.0±$\pm 0.71$ 80.0±$\pm 1.58$ 35.6±$\pm 0.89$
+Ns-Transformer 52.6±$\pm 2.30$ 88.0±$\pm 0.71$ 78.8±$\pm 0.84$ 77.2±$\pm 1.48$ 34.8±$\pm 1.10$
+Informer 51.2±$\pm 1.30$ 86.6±$\pm 1.14$ 78.0±$\pm 1.58$ 74.0±$\pm 0.71$ 35.6±$\pm 0.55$
+Transformer 49.6±$\pm 1.67$ 85.4±$\pm 0.89$ 77.0±$\pm 0.71$ 75.2±$\pm 1.30$ 32.8±$\pm 0.84$
+iTransformer 48.4±$\pm 1.82$ 81.8±$\pm 0.84$ 76.6±$\pm 0.55$ 80.4±$\pm 1.14$ 19.8±$\pm 0.84$
+TimesNet 52.2±$\pm 2.39$ 87.4±$\pm 1.14$ 76.2±$\pm 1.92$ 78.4±$\pm 1.52$ 34.8±$\pm 0.84$
+GPT4TS 54.2±$\pm 2.05$ 88.2±$\pm 0.84$ 80.4±$\pm 0.89$ 76.4±$\pm 1.14$ 32.8±$\pm 1.10$
+Chronos+MLP 44.2±$\pm 1.30$ 82.2±$\pm 0.84$ 79.8±$\pm 0.45$ 83.0±$\pm 0.71$ 38.0±$\pm 0.71$
+DeepConvLSTM 48.8±$\pm 2.39$ 89.2±$\pm 0.84$ 78.4±$\pm 1.52$ 75.0±$\pm 1.87$ 40.4±$\pm 0.89$
+DeepConvLSTMAtt 54.0±$\pm 2.12$ 89.6 ±$\pm 1.14$ 79.2±$\pm 1.30$ 77.4±$\pm 2.19$ 41.4±$\pm 0.55$
+Attend 60.2±$\pm 2.17$ 93.2±$\pm 0.84$ 84.6±$\pm 1.14$ 83.4±$\pm 1.14$ 43.6±$\pm 0.55$
+SensorLLM 61.2±$\pm 3.56$ 91.2±$\pm 1.48$ 86.2±$\pm 1.48$ 89.4±$\pm 3.85$ 48.6±$\pm 1.14$
+Table 2: F1-macro results (%) for the Task-Aware Tuning Stage, presented as the mean and standard deviation over 5 random repetitions.
+The top results for each dataset are highlighted as follows: Bold for the best and underline for the second-best.
 4.3 Human Activity Recognition
-Setups.
+Setup.
 In this section, we evaluate the performance of SensorLLM on HAR tasks.
-For all datasets, experiments were run five times with the following parameters: 8 training epochs, a batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 4096.
-We report the macro F1 score A.7 to account for class imbalance across different activity categories.
+Each experiment runs for five trials, using 8 training epochs, a batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 4096.
+We report the F1 macro score A.8 to account for class imbalance across different activity categories.
 Baselines.
-To benchmark our model, we compare it against two sets of baselines.
-The first set includes state-of-the-art time-series (TS) models: Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021), NS-Transformer (Liu et al., 2022), PatchTST (Nie et al., 2023), TimesNet (Wu et al., 2023a), and iTransformer (Liu et al., 2024c).
-The second set consists of state-of-the-art HAR models: DeepConvLSTM (Ordóñez & Roggen, 2016), DeepConvLSTM + Attention (DeepConvLSTMAttn) (Murahari & Plötz, 2018), and Attend (Abedin et al., 2021).
-Detailed introduction of the baselines can be found in the Appendix A.6.
-Results. As shown in Figure 4, SensorLLM outperforms the TS baselines across all datasets and surpasses the HAR baselines on USC-HAD, MHealth, and PAMAP2.
-On USC-HAD, SensorLLM improves the median performance of the top TS baseline NS-Transformer by 8% and matches the top HAR baseline Attend, with a 3% gain in the best results.
-For UCI-HAR, SensorLLM beats NS-Transformer by 3% in median performance but trails 2% short of Attend.
-On PAMAP2, SensorLLM surpasses PatchTST by 4% in median performance and exceeds Attend by 1%, with a 2% improvement in best results.
-For MHealth, SensorLLM outperforms both PatchTST and iTransformer by 9% and improves on Attend by 6%, with a 10% gain in best results.
-The fact that the TS embedder Chronos was not pre-trained on sensor data further highlights SensorLLM’s potential to understand sensor data and achieve superior performance in HAR and other related tasks.
+We benchmark SensorLLM against 11 baselines across two categories: (i) TS models—Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021), NS-Transformer (Liu et al., 2022), PatchTST (Nie et al., 2023), TimesNet (Wu et al., 2023a), and iTransformer (Liu et al., 2024c); (ii) HAR models—DeepConvLSTM (Ordóñez and Roggen, 2016), DeepConvLSTMAttn (Murahari and Plötz, 2018), and Attend (Abedin et al., 2021).
+We also include Chronos+MLP and GPT4TS (Zhou et al., 2023a) for a more comprehensive comparison.
+Full baseline details are in Appendix A.7.
+Results.
+Table 2 reports the macro F1 scores (%) averaged over five random runs.
+SensorLLM achieves the highest performance on four datasets (USC-HAD, PAMAP2, MHealth, CAPTURE-24) and ranks second on UCI-HAR, demonstrating its effectiveness in handling diverse sensor data.
+Notably, SensorLLM achieves a notable improvement on the CAPTURE-24 dataset, surpassing all baselines by a significant margin with a mean F1-macro score of 48.6%, which is 5.0% higher than Attend (43.6%).
+On USC-HAD, SensorLLM achieves the highest score of 61.2%, outperforming Attend, the second-best baseline, by 1.0%.
+Similarly, on PAMAP2, SensorLLM achieves a score of 86.2%, exceeding Attend (84.6%) by 1.6%.
+On MHealth, SensorLLM sets a new state-of-the-art with a score of 89.4%, surpassing Attend (83.4%) by 6.0%.
+These results highlight SensorLLM’s ability to consistently outperform existing methods across diverse datasets.
+For UCI-HAR, SensorLLM achieves the second-best score (91.2%), slightly trailing Attend (93.2%).
+In contrast, Chronos+MLP shows only a slight improvement over iTransformer, the lowest-performing baseline (82.2% vs.
+81.8%), indicating that Chronos embeddings alone have limited utility for HAR on this dataset.
+However, our framework significantly enhances their effectiveness, highlighting the robustness of our alignment approach.
 5 Ablation Studies
-In this section, we present our ablation studies to further demonstrate the effectiveness of our proposed method.
-Figure 5: Experiments demonstrating the effectiveness of our Sensor-Language Alignment. The results shown are the averages from five random runs.
-Impact of Alignment on HAR Performance Due to the limitations of Chronos’ pretraining sequence length, we were unable to use Chronos directly for HAR as a baseline for comparison.
-However, we tested the impact of bypassing Sensor-Language Alignment Stage by using Task-Aware Tuning Stage alone, directly perform HAR tasks using Chronos embeddings and the LLM.
-The results, shown in Figure 5, indicate that across all four datasets, SensorLLM outperforms the model that skips this step (Task-only).
-The performance of the Task-only model is either on par with or below the TS baselines, further highlighting the importance of our sensor-language alignment.
-These findings validate the effectiveness of aligning sensor-text modalities, as the integration of temporal and semantic features in the embeddings improves the LLM’s understanding of sensor data and significantly enhances HAR outcomes.
-Table 2: The results for SensorLLM trained with/without text prompts.
-Task-only refers to conducting HAR directly bypassing sensor-language alignment.
-We report the F1-macro results in % with a mean and standard deviation of 5 repetitions.
-Dataset Task-only (w/o prompts) Task-only (w/ prompts) SensorLLM (w/o prompts) SensorLLM (w/ prompts)
+Dataset Task-only SensorLLM
+w/o prompts w/ prompts w/o prompts w/ prompts
 USC-HAD 43.4±$\pm 2.88$ 45.0±$\pm 1.58$ 49.6±$\pm 1.67$ 61.2±$\mathbf{\pm 3.56}$
@@
 MHealth 76.6±$\pm 1.34$ 77.4±$\pm 3.13$ 86.6±$\pm 1.14$ 89.4±$\mathbf{\pm 3.85}$
-Impact of Prompts To evaluate the contribution of additional textual information (statistical information for each channel’s sensor data) in Task-Aware Tuning Stage, we conducted experiments comparing SensorLLM’s performance with and without the inclusion of prompts.
-As shown in Table 2, the inclusion of prompts consistently improves F1-macro scores across all datasets, with a more pronounced effect when used in the full SensorLLM architecture.
-This indicates that the model effectively integrates both sensor and textual data, allowing it to better understand complex temporal patterns.
-The results highlight the advantages of leveraging multimodal inputs, which enrich the model’s representation of sensor data and lead to more accurate human activity recognition.
-Overall, the ability to process both sensor data and textual prompts not only boosts classification performance but also paves the way for LLMs to tackle more complex sensor-driven tasks in future applications.
-Table 3: Cross-dataset experiments. We report the F1-macro results in % with a mean and standard deviation of 5 repetitions.
+CAPTURE-24 44.8±$\pm 0.84$ 46.0±$\pm 0.71$ 47.2±$\pm 0.84$ 48.6±$\mathbf{\pm 1.14}$
+Table 3: The results for SensorLLM trained with/without text prompts. Task-only refers to conducting HAR directly bypassing sensor-language alignment.
+Impact of Alignment.
+To evaluate the impact of alignment, we included the Chronos+MLP baseline in Section 4.3 to show that SensorLLM’s performance is not solely driven by Chronos embeddings.
+Additionally, we compared SensorLLM with a Task-only model, which skips the Sensor-Language Alignment Stage and directly applies Chronos embeddings and the LLM for HAR.
+As shown in Table 3, SensorLLM consistently outperforms the Task-only model across all five datasets, regardless of additional textual input.
+Notably, the Task-only model often performs on par with or worse than traditional TS baselines, underscoring the importance of our alignment method.
+These results confirm that Chronos embeddings alone are insufficient for optimal HAR performance and that alignment is essential for enabling LLMs to effectively understand sensor data.
+Impact of Prompts.
+To assess the role of additional textual information (e.g., statistical features for each sensor channel) in the Task-Aware Tuning Stage, we compared SensorLLM’s performance with and without prompts.
+As shown in Table 3, incorporating prompts consistently improves F1-macro scores across all datasets, with a more pronounced effect in the full SensorLLM architecture.
+This demonstrates that the model effectively integrates sensor and textual data, enhancing its ability to capture complex temporal patterns.
+The results highlight the benefits of multimodal inputs, which enrich sensor data representations and improve HAR accuracy.
+More broadly, the ability to process both sensor signals and textual prompts not only enhances classification performance but also extends LLMs’ potential for tackling complex sensor-driven tasks in future applications.
+Figure 3: Effect of the number of alignment module layers.
+Alignment Module Layers.
+We investigate the effect of hidden layer depth in the alignment module MLP on UCI-HAR, PAMAP2, and MHealth.
+As shown in Figure 3, increasing the number of hidden layers from one ($1024\to 2048\to 4096$) to two ($1024\to 2048\to 3072\to 4096$) led to mixed results.
+The average F1-macro scores improved on UCI-HAR (91.2% → 92.0%) and MHealth (89.4% → 90.2%), but slightly dropped on PAMAP2 (86.2% → 85.8%).
+These results suggest that the optimal number of hidden layers varies by dataset.
+Effect of Model Size on Performance.
+To assess the impact of model size, we tested SensorLLM-3b, a resource-efficient variant using Chronos-base and Llama3.2-3b.
+Experiments were conducted on USC-HAD, UCI-HAR, and MHealth.
+As shown in Figure 4, SensorLLM-3b achieves slightly lower performance than SensorLLM-8b across all datasets, illustrating the trade-off between model size and accuracy.
+However, SensorLLM-3b remains competitive, outperforming Attend on USC-HAD and MHealth, while trailing it only on UCI-HAR.
+These results indicate that SensorLLM-3b offers a resource-efficient alternative, maintaining strong performance relative to other baselines.
+Figure 4: Effect of Model Size.
+Cross-Dataset Generalization.
+To assess the robustness of SensorLLM, we conducted cross-dataset experiments, training the Sensor-Language Alignment Stage on USC-HAD and the Task-Aware Tuning Stage on UCI-HAR, and vice versa.
+These datasets share the same sensor channels but differ in sampling rates.
+As shown in Table 4, SensorLLM achieves performance comparable to training both stages on the same dataset.
+This indicates that once modality alignment is established, retraining on new datasets is unnecessary.
+SensorLLM effectively generalizes to downstream tasks across diverse datasets, demonstrating that the alignment stage enables the LLM to truly understand sensor data, rather than memorizing dataset-specific features.
+These results highlight SensorLLM’s potential for broad cross-dataset generalization, laying the groundwork for future TS-LLM models.
 Stage 1 Stage 2 Results
@@
 UCI-HAR USC-HAD 61.6±$\mathbf{\pm 2.07}$
-Cross-Dataset Generalization To further assess the robustness of SensorLLM, we conducted cross-dataset experiments, where Sensor-Language Alignment Stage was trained on the USC-HAD dataset and Task-Aware Tuning Stage was trained on the UCI-HAR dataset (and vice versa).
-We chose these two datasets because they share the same sensor channels, although they differ in sample rates.
-As shown in Table 3, SensorLLM maintained strong performance, achieving results similar to those obtained when using the same dataset for both stages in Figure 4.
-These findings demonstrate that once the model effectively understands sensor data through modality alignment, retraining on new datasets for alignment becomes unnecessary.
-SensorLLM can directly perform downstream tasks across diverse datasets, laying the foundation for future TS-LLM models capable of generalizing across a wide range of time-series tasks.
-7 Conclusion and future work
-In this study, we introduced SensorLLM, a novel multimodal framework for aligning sensor data with automatically generated text.
-And unlike previous alignment methods which operate at the machine level, our approach aligns sensor data and text at the human perception level.
-SensorLLM thus can capture and interpret complex sensor data patterns—typically challenging for pre-trained LLMs—while outperforming or matching state-of-the-art models in HAR tasks.
-The results of cross-datasets experiments, without the need for dataset-specific alignment, also demonstrate SensorLLM’s strong generalizability and robustness.
-We have open-sourced our code and data generation methods, providing a valuable resource for the community and opening new avenues for aligning time-series and text data, particularly in domains with limited textual labels.
-Future work could explore more advanced analyses, such as chain-of-thought reasoning, sensor data generation, and zero-shot or few-shot learning, further enhancing LLMs’ potential in sensor-based tasks.
+Table 4: Cross-dataset experiments.
+6 Conclusions
+We introduced SensorLLM, a multimodal framework that aligns sensor data with automatically generated text at a human-perception level, moving beyond machine-level alignment.
+SensorLLM effectively captures complex sensor patterns, achieving superior performance in HAR tasks.
+Experiments across diverse datasets demonstrate its robustness in handling variable-length sequences, multi-channel inputs, and textual metadata.
+Cross-dataset results further highlight its strong generalizability without requiring dataset-specific alignment.
+This work establishes a foundation for Sensor-Text MLLMs, with potential applications for sensor data analysis.
+We release our code and data generation pipeline to facilitate future research on integrating time-series and text, particularly in low-resource domains.
+7 Limitations
+While SensorLLM demonstrates strong performance in aligning sensor data with LLMs, certain limitations remain, offering directions for future exploration.
+Classifier-Based Design.
+To ensure fair comparisons with existing HAR models (classification models), we adopt a classifier for downstream tasks rather than fully exploiting the LLM’s generative abilities.
+Although our results verify that the Sensor-Language Alignment Stage can generalize across datasets, relying on a fixed-class classifier may constrain the model’s adaptability to new activity categories.
+Future work could explore using the LLM in a generative or prompt-based capacity, enabling broader application scenarios such as activity discovery or open-set recognition.
+Scope of Sensor-Text Alignment.
+Our alignment focuses on mapping sensor data to trend-descriptive text, demonstrating clear benefits for LLM-based HAR.
+However, human-intuitive descriptions of sensor data extend beyond trend changes—incorporating frequency-domain features, periodicity, and higher-order patterns may further enhance an LLM’s ability to interpret time-series data.
+Future research could investigate whether aligning text with alternative sensor characteristics improves time-series reasoning.
+This could expand the potential of multimodal NLP applications in sensor-driven tasks beyond activity recognition.
+8 Acknowledgements
+This research includes computations using the computational cluster Wolfpack supported by School of Computer Science and Engineering at UNSW Sydney.
 Appendix A Appendix
@@
 Bevilacqua et al.
-(2019) developed a CNN-based model for HAR, while Ha & Choi (2016) introduced CNN-pf and CNN-pff architectures that apply partial and full weight sharing for better feature extraction.
+(2019) developed a CNN-based model for HAR, while Ha and Choi (2016) introduced CNN-pf and CNN-pff architectures that apply partial and full weight sharing for better feature extraction.
 Other notable works include Perception-Net Kasnesis et al.
@@
 LLMs have achieved remarkable success in text-related tasks, and their utility has expanded into time-series forecasting.
-Xue & Salim (2023) presents PromptCast, which redefines time-series forecasting as a natural language generation task by transforming numerical inputs into textual prompts, enabling pre-trained language models to handle forecasting tasks with superior generalization in zero-shot settings.
+Xue and Salim (2023) presents PromptCast, which redefines time-series forecasting as a natural language generation task by transforming numerical inputs into textual prompts, enabling pre-trained language models to handle forecasting tasks with superior generalization in zero-shot settings.
 Gruver et al.
@@
 Zhou et al.
-(2023a) demonstrates that pre-trained language and image models, such as a Frozen Pretrained Transformer (FPT), can be adapted for diverse time-series tasks like classification, forecasting, and anomaly detection, leveraging self-attention mechanisms to bridge the gap between different data types and achieving state-of-the-art performance across various tasks.
+(2023b) demonstrates that pre-trained language and image models, such as a Frozen Pretrained Transformer (FPT), can be adapted for diverse time-series tasks like classification, forecasting, and anomaly detection, leveraging self-attention mechanisms to bridge the gap between different data types and achieving state-of-the-art performance across various tasks.
 Jin et al.
@@
 Sun et al.
-(2024a) proposes TEST, an innovative embedding technique that integrates time-series data with LLMs through instance-wise, feature-wise, and text-prototype-aligned contrast, yielding improved or comparable results across various applications.
+(2024b) proposes TEST, an innovative embedding technique that integrates time-series data with LLMs through instance-wise, feature-wise, and text-prototype-aligned contrast, yielding improved or comparable results across various applications.
 Chang et al.
@@
 A.2 Data Generation
-In this section, we outline the process for generating text data from sensor readings.
-Our approach utilizes a set of predefined sentence templates, which are randomly selected to create diverse question-answer (QA) pairs.
-To enhance the variability, we use GPT-4o to generate synonymous versions of the sentences within the templates.
-Each sentence contains placeholders for relevant numerical values (e.g., timestamps, sensor readings) or textual information.
-We also choose a variety of synonym expressions for each trend.
-These values are then inserted into the placeholders to form coherent QA pairs corresponding to the sensor data inputs.
-Table 4: Question templates examples.
-Description: Kindly provide a detailed analysis of the trend changes observed in the {data}.
+We generate text data from sensor readings using predefined sentence templates (5, Tables 6,, 7).
+These templates are randomly selected to create diverse question-answer (QA) pairs.
+To enhance variability, we employ GPT-4o to generate synonymous variations.
+Each sentence contains placeholders for numerical values (e.g., timestamps, sensor readings) or textual information, which are dynamically replaced to produce coherent QA pairs aligned with the sensor data.
+Trend Description Templates
+{start_time}s to {end_time}s: {trend} {start_time} seconds to {end_time} seconds: {trend} {start_time} to {end_time} seconds: {trend} {start_time}-{end_time} seconds: {trend} {start_time}-{end_time}s: {trend} {start_time}s-{end_time}s: {trend}
+Table 5: Examples of answer templates used for trend descriptions.
+Trend Description Templates
+Kindly provide a detailed analysis of the trend changes observed in the {data}.
 Please offer a comprehensive description of how the trends in the {data} have evolved.
 I would appreciate a thorough explanation of the trend fluctuations that occurred within the {data}.
-Could you please examine the {data} in depth and explain the trend shifts observed step by step?
+Could you examine the {data} in depth and explain the trend shifts observed step by step?
 Detail the {data}’s trend transitions.
-Could you kindly assess the {data} and provide a description of the trend transformations that took place step by step?
+Could you assess the {data} and describe the trend transformations step by step?
 Could you analyze the trends observed in the {data} over the specified period step by step?
@@
 What trend changes can be seen in the {data}?
-Summary: Could you provide a summary of the main features of the input {data} and the distribution of the trends?
+Summary Templates
+Could you provide a summary of the main features of the input {data} and the distribution of the trends?
 Please give an overview of the essential attributes of the input {data} and the spread of the trends.
@@
 I need a summary of {data}’s main elements and their trend distributions.
-Table 5: Answer templates examples.
-Description: {start_time}s to {end_time}s: {trend} {start_time} seconds to {end_time} seconds: {trend} {start_time} to {end_time} seconds: {trend} {start_time}-{end_time} seconds: {trend} {start_time}-{end_time}s: {trend} {start_time}s-{end_time}s: {trend}
-Summary 1: Number of {trend} trends: {num} Count of {trend} trends: {num} Number of {trend} segments: {num} Count of {trend} segments: {num}
-Summary 2: The given {data_name} representing the {sensor_name} sensor readings from {start_time}s to {end_time}s.
-The {data_name} represents readings taken from the {sensor_name} sensor between {start_time} and {end_time} seconds.
-The {sensor_name} sensor readings recorded within the {start_time} to {end_time} second timeframe are presented in this {data_name}.
-Summary 3: The data exhibits {trend_num} distinct trends, with a total of {change_num} changes in trend observed.
+Table 6: Examples of question templates used for trend description and summary generation.
+Summary 1: Trend Count
+Number of {trend} trends: {num} Count of {trend} trends: {num} Number of {trend} segments: {num} Count of {trend} segments: {num}
+Summary 2: Sensor Data Context
+The given {data_name} represents {sensor_name} sensor readings from {start_time}s to {end_time}s.
+The {data_name} contains {sensor_name} sensor readings recorded between {start_time} and {end_time} seconds.
+The {sensor_name} sensor readings collected from {start_time} to {end_time} seconds are presented in this {data_name}.
+Summary 3: Trend Change Statistics
+The data exhibits {trend_num} distinct trends, with {change_num} trend changes observed.
 Across {trend_num} trends, the data shows {change_num} occurrences of trend shifts.
-{trend_num} trends are present in the data, with {change_num} instances of trend changes.
-Summary 4: To sum up, the data exhibited a {trend_type} trend for a cumulative period of {total_time} seconds.
-Overall, the data showed a {trend_type} trend over {total_time} seconds.
-To conclude, the trend was {trend_type} over a period of {total_time} seconds.
-Summary 5: The overall trend is {overall_trend}. The primary trend detected is {overall_trend}. Looking at the big picture, the trend is {overall_trend}.
-The system prompt guides the model on how to respond to the questions generated, incorporating characteristics specific to each dataset, such as sensor frequency and other dataset-specific attributes.
-These tailored prompts ensure the responses are aligned with the unique properties of the sensor data.
-Below is the template of the system prompts (instructions) used for each dataset:
-  • A dialogue between a curious researcher and an AI assistant.
-The AI analyzes a sensor time-series dataset (N points, {sample_rate}Hz sampling rate) to answer specific questions.
-This interaction demonstrates the AI’s data analysis skills and the potential of human-AI collaboration in interpreting complex data.
-A.3 GPT-4o Prompt for Sensor Data Trend Analysis
-Table 6 outlines the system prompt used to generate trend-descriptive texts from sensor data, providing a structured format for GPT-4o to analyze and respond to specific questions about the data.
-This prompt serves as a standardized way for GPT-4o to examine the time-series data and produce outputs that can then be directly compared to the descriptions generated by our SensorLLM model.
-By using this prompt, we aim to evaluate GPT-4o’s ability to interpret numerical sensor data and generate accurate trend descriptions.
-The responses generated by GPT-4o are assessed against both human evaluations and NLP metrics to benchmark its performance against SensorLLM, allowing us to critically analyze the differences in how both models understand and interpret time-series data trends.
-This comparison helps demonstrate the effectiveness of our Sensor-Language Alignment Stage of SensorLLM.
-Table 6: Prompt for GPT-4o to generate descriptive texts based on the given numerical sensor data.
+{trend_num} trends are present, with {change_num} instances of trend changes.
+Summary 4: Cumulative Trend Analysis
+To sum up, the data exhibited a {trend_type} trend for a total duration of {total_time} seconds.
+Overall, the data showed a {trend_type} trend spanning {total_time} seconds.
+In conclusion, the trend was {trend_type} over {total_time} seconds.
+Summary 5: Overall Trend Summary
+The overall trend is {overall_trend}. The primary trend detected is {overall_trend}. Looking at the broader pattern, the trend is {overall_trend}.
+Table 7: Examples of answer templates used for summaries.
+The system prompt instructs the model on how to respond to generated questions, incorporating dataset-specific attributes such as sensor frequency and sampling rate.
+These tailored prompts ensure responses align with the unique characteristics of each dataset.
+Below is the system prompt template used for all datasets:
+  • A dialogue between a researcher and an AI assistant.
+The AI analyzes a sensor time-series dataset (N points, sampled at {sample_rate}Hz) to answer specific questions, demonstrating its analytical capabilities and the potential for human-AI collaboration in interpreting sensor data.
+A.3 Chronos
+Chronos (Ansari et al., 2024) is a pretrained probabilistic time-series framework that tokenizes real-valued time-series data into discrete representations for language model training.
+It utilizes scaling and quantization to transform time-series data into a fixed vocabulary, enabling T5-based (Raffel et al., 2020) models to learn from tokenized sequences using cross-entropy loss.
+Pretrained on diverse public and synthetic datasets, Chronos surpasses existing models on familiar datasets and demonstrates strong zero-shot performance on unseen tasks, making it a versatile tool for time-series forecasting across domains.
+Time-Series Tokenization and Quantization.
+Chronos converts time-series data into discrete tokens through a two-step process: normalization and quantization. Mean scaling is first applied to ensure consistency across different time series:
+$\tilde{x}=\frac{x}{\text{mean}(\lvert x\rvert)}$ (5)
+Next, the normalized values are quantized using $B$ bin centers $c_{1},\dots,c_{B}$ and corresponding bin edges $b_{1},\dots,b_{B-1}$, mapping real values to discrete tokens via:
+$q(x)=\begin{cases}1&\text{if}\ -\infty\leq x<b_{1},\\ 2&\text{if}\ b_{1}\leq x<b_{2},\\ \vdots\\ B&\text{if}\ b_{B-1}\leq x<\infty.\end{cases}$ (6)
+Special tokens such as PAD and EOS are added to handle sequence padding and denote the end of sequences, allowing Chronos to process variable-length inputs efficiently within language models.
+Objective Function.
+Chronos models the tokenized time series using a categorical distribution over the vocabulary $V_{ts}$, minimizing the cross-entropy loss:
+$\begin{split}\ell(\theta)=-\sum_{h=1}^{H+1}\sum_{i=1}^{|V_{ts}|}\mathbf{1}(z_{C+h+1}=i)\\ \cdot\log p_{\theta}(z_{C+h+1}=i\mid z_{1:C+h})\end{split}$ (7)
+where $C$ is the historical context length, $H$ is the forecast horizon, and $p_{\theta}$ is the predicted token distribution.
+This approach offers two key advantages: (i) Seamless integration with language models, requiring no architectural modifications, and (ii) Flexible distribution learning, enabling robust generalization across diverse time-series datasets.
+A.4 GPT-4o Prompt for Sensor Data Trend Analysis
+Table 8 presents the system prompt used to generate trend-descriptive texts from sensor data, providing a structured framework for GPT-4o to analyze and respond to specific questions.
+This standardized prompt ensures consistency in GPT-4o’s interpretation of time-series data, allowing direct comparison with descriptions produced by SensorLLM.
 Prompt A dialogue between a curious researcher and an AI assistant. The AI analyzes a sensor time-series dataset (N points, {sr}Hz sampling rate) to answer specific questions.
@@
 Output:
-A.4 Chronos
-Chronos (Ansari et al., 2024) is an innovative framework for pretrained probabilistic time series models.
-It tokenizes time series data into a fixed vocabulary through scaling and quantization, then trains T5-based (Raffel et al., 2020) language models on these tokenized sequences using cross-entropy loss.
-Pretrained on a diverse corpus of public and synthetic datasets, Chronos models demonstrate remarkable performance, outperforming other methods on familiar datasets and showing competitive zero-shot capabilities on unseen tasks.
-This approach leverages data from various domains to enhance forecasting accuracy, potentially simplifying time series analysis pipelines across different fields.
-Time-series Tokenization and Quantization.
-Chronos tokenizes real-valued time-series data through a two-step process: normalization and quantization.
-First, mean scaling is applied for normalization, where each observation is divided by the mean of its absolute values:
-$\tilde{x}=\frac{x}{\text{mean}(\lvert x\rvert)}.$
-This ensures that the data is scaled consistently across different series.
-After normalization, the values are quantized into discrete tokens.
-Quantization is performed by selecting $B$ bin centers $c_{1}<\dots<c_{B}$ along with $B-1$ edges $b_{i}$ separating them.
-Each real value is mapped to a token based on the quantization function $q(x)$, defined as:
-$q(x)=\begin{cases}1&\text{if}\ -\infty\leq x<b_{1},\\ 2&\text{if}\ b_{1}\leq x<b_{2},\\ \vdots\\ B&\text{if}\ b_{B-1}\leq x<\infty,\end{cases}$
-where each token represents a quantized range of values. This quantization ensures that the continuous time-series data can be represented as discrete tokens suitable for language models.
-To further facilitate the modeling, special tokens such as PAD and EOS are added to the time-series vocabulary $V_{ts}$ to handle sequence padding and denote the end of sequences.
-This allows Chronos to efficiently process variable-length sequences and integrate them with language modeling tasks, treating the time-series data as token sequences.
-Objective Function.
-In Chronos, the model uses a categorical distribution over the quantized time-series tokens from the vocabulary $V_{ts}$.
-The objective is to minimize the cross-entropy between the predicted categorical distribution and the ground-truth labels, including EOS tokens.
-The loss function for a single time series is defined as:
-$\ell(\theta)=-\sum_{h=1}^{H+1}\sum_{i=1}^{|V_{ts}|}{\bf\mathbf{1}}(z_{C+h+1}=i)\log p_{\theta}(z_{C+h+1}=i\mid z_{1:C+h})$
-where $C$ represents historical context of length and $H$ represents a forecast horizon, and $p_{\theta}(z_{C+h+1}=i\mid z_{1:C+h})$ is the predicted distribution.
-The loss is averaged over batches of time series.
-This categorical output distribution offers two key advantages: (i) it requires no modification to the language model, allowing easy integration with existing architectures; and (ii) it supports learning arbitrary distributions, making it flexible for diverse time-series datasets.
+Table 8: Prompt for GPT-4o to generate descriptive texts based on the given numerical sensor data.
+We evaluate GPT-4o’s ability to interpret numerical sensor data by assessing its responses against human evaluations and NLP metrics.
+This comparison benchmarks GPT-4o’s performance against SensorLLM, highlighting differences in how both models process time-series data trends.
+The results demonstrate the effectiveness of SensorLLM’s Sensor-Language Alignment Stage.
 A.5 Evaluation Metrics for Sensor-Language Alignment Stage
@@
 ROUGE-L measures the longest common subsequence, assessing both precision and recall in terms of structure and content overlap, though it does not evaluate semantic accuracy.
-METEOR (Banerjee & Lavie, 2005).
+METEOR (Banerjee and Lavie, 2005).
 METEOR (Metric for Evaluation of Translation with Explicit Ordering)combines precision and recall, with additional alignment techniques such as stemming and synonym matching.
@@
 However, its emphasis is still on word-level alignment rather than factual accuracy or meaning.
-SBERT (Reimers & Gurevych, 2019).
+SBERT (Reimers and Gurevych, 2019).
 SBERT (Sentence-BERT)  [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 is a metric that generates sentence embeddings using the BERT architecture.
@@
 This type of evaluation provides insights into how well the generated outputs capture the nuances of sensor data trends in a manner similar to human understanding.
-Table 7: Prompt and output examples for GPT-4o in evaluating model-generated texts and ground-truth.
 Prompt Please evaluate the model-generated trend descriptions against the ground truth.
@@
 Deduct 1 point for minor errors in the trend description, and 2-3 points for moderate errors.
-Provide your score (1-5) and a brief explanation in the format: ”score#reason” (e.g., 4#The description of trend changes slightly differs from the ground truth).
+Provide your score (1-5) and a brief explanation in the format: "score#reason" (e.g., 4#The description of trend changes slightly differs from the ground truth).
 Now, please proceed to score the following:
@@
 Output example 2: 5#The model’s description matches the human-generated text accurately.
+Table 9: Prompt and output examples for GPT-4o in evaluating model-generated texts and ground-truth.
 Human Evaluation.
@@
 This manual evaluation serves as an important benchmark for the model’s performance from a human perspective, ensuring that the generated outputs are not only technically correct but also practically useful for human interpretation.
-A.6 Baselines for Task-Aware Tuning Stage
+A.6 Datasets
+We used five datasets in our study:
+USC Human Activity Dataset (USC-HAD).
+USC-HAD (Zhang and Sawchuk, 2012) consists of six sensor readings from body-worn 3-axis accelerometers and gyroscopes, collected from 14 subjects.
+The data is sampled at 100 Hz across six channels and includes 12 activity class labels.
+For evaluation, we use data from subjects 13 and 14 as the test set, while the remaining subjects’ data are used for training.
+A window size $w\in[5,200]$ is used in alignment stage, and $w=200$ with stride of $100$ are used in HAR.
+UCI Human Activity Recognition Dataset (UCI-HAR).
+UCI-HAR (Anguita et al., 2013) includes data collected from 30 volunteers performing six activities while wearing a smartphone on their waist.
+The embedded accelerometer and gyroscope sensors sampled data at 50 Hz across six channels.
+The dataset was partitioned into 70% for training and 30% for testing.
+A window size $w\in[5,200]$ is used in alignment stage, and $w=128$ with stride of $64$ is used in HAR.
+Physical Activity Monitoring Dataset (PAMAP2).
+PAMAP2 (Reiss and Stricker, 2012) includes data from nine subjects wearing IMUs on their chest, hands, and ankles.
+IMUs capture the acceleration, gyroscope, and magnetometer data across 27 channels and include 12 activity class labels.
+For our experiments, data from subjects 105 and 106 are used as the test set, with the remaining subjects’ data used for training.
+The sample rate is downsampled from 100 Hz to 50 Hz.
+A window size $w\in[5,100]$ is used in alignment stage, and $w=100$ with stride of $50$ in HAR.
+Mobile Health Dataset (MHealth).
+MHealth (Baños et al., 2014) contains body motion and vital sign recordings from ten volunteers.
+Sensors were placed on the chest, right wrist, and left ankle of each subject.
+For our experiments, we used acceleration data from the chest, left ankle, and right lower arm, along with gyroscope data from the left ankle and right lower arm, resulting in a total of 15 channels.
+The data is sampled at 50 Hz and includes 12 activity class labels.
+Data from subjects 1, 3, and 6 is used as the test set, while the remaining subjects’ data are used for training.
+We use a window size $w\in[5,100]$ in alignment stage and $w=100$ with stride of $50$ in HAR.
+CAPTURE-24.
+CAPTURE-24 (Chan et al., 2024) is a large-scale dataset featuring 3-channel wrist-worn accelerometer data collected in free-living settings for over 24 hours per participant.
+It includes annotated data from 151 participants, making it significantly larger than existing datasets.
+We used the first 100 participants as the training set and the remaining 51 as the test set.
+For each subject, sequences were windowed, and 5% of the data was randomly selected for training and testing.
+The sample rate was downsampled from 100 Hz to 50 Hz and it includes 10 activity class labels.
+During the alignment stage, we used a variable window size $w\in[10,500]$, while in the HAR, we fixed $w=500$ with a stride of $250$.
+Each dataset includes multiple activity classes, and the proportion of each class in the dataset is shown in Table 10.
+Dataset # Classes Classes Proportions (%)
+USC-HAD 12
+Sleeping, Sitting, Elevator down,
+Elevator up, Standing, Jumping,
+Walking downstairs, Walking right,
+Walking forward, Running forward,
+Walking upstairs, Walking left
+12.97, 9.06, 6.04, 5.94, 8.6, 3.62,
+7.61, 9.81, 13.15, 5.72, 8.22, 9.25
+UCI-HAR 6
+Standing, Sitting, Laying,
+Walking, Walking downstairs,
+Walking upstairs
+18.69, 17.49, 19.14,
+16.68, 13.41, 14.59
+PAMAP2 12
+Lying, Sitting, Standing,
+Ironing, Vacuum cleaning,
+Ascending stairs, Descending stairs,
+Walking, Nordic walking, Cycling,
+Running, Rope jumping
+10.25, 9.52, 10.11,
+11.82, 9.14, 6.3,
+5.67, 12.77, 9.52,
+8.42, 3.57, 2.91
+MHealth 12
+Climbing stairs, Standing still,
+Sitting and relaxing, Lying down,
+Walking, Waist bends forward,
+Frontal elevation of arms,
+Knees bending (crouching),
+Jogging, Running, Jump front
+& back, Cycling
+8.91, 8.95, 8.95,
+8.95, 8.95, 8.26,
+8.7, 8.53, 8.95,
+8.95, 2.96, 8.95
+CAPTURE-24 10
+Sleep, Household-chores, Walking,
+Vehicle, Standing, Mixed-activity,
+Sitting, Bicycling, Sports,
+Manual-work
+37.45, 6.5, 6.16,
+3.83, 3.25, 3.49,
+37.07, 1.03, 0.43, 0.79
+Table 10: Dataset classes and Proportions
+A.7 Baselines for Task-Aware Tuning Stage
 In Task-Aware Tuning Stage, we compare SensorLLM against several state-of-the-art baseline models for time-series classification and human activity recognition (HAR).
@@
 Time points of individual series are embedded as variate tokens, allowing the attention mechanism to capture multivariate correlations, while the feed-forward network learns nonlinear representations for each token.
-DeepConvLSTM (Ordóñez & Roggen, 2016).
+DeepConvLSTM (Ordóñez and Roggen, 2016).
 DeepConvLSTM integrates four consecutive convolutional layers followed by two LSTM layers to effectively capture both spatial and temporal dynamics in sensor data.
 The final output vector is passed through a fully connected layer, and the softmax function is applied to produce activity class probabilities as the model’s final output.
-DeepConvLSTMAttn (Murahari & Plötz, 2018).
+DeepConvLSTMAttn (Murahari and Plötz, 2018).
 DeepConvLSTMAttn enhances the original DeepConvLSTM by integrating an attention mechanism to improve temporal modeling in HAR tasks.
@@
 These innovations result in more discriminative activity representations, significantly improving HAR performance.
-A.7 Evaluation Metrics for Task-Aware Tuning Stage
+Chronos+MLP.
+Chronos (Ansari et al., 2024)+MLP is a baseline designed to evaluate whether the performance gains in SensorLLM are solely attributable to Chronos and the MLP.
+In SensorLLM, Chronos is used to generate sensor embeddings, which are then mapped by the MLP for input into the LLM to perform HAR.
+Since Chronos does not natively support classification tasks and only processes single-channel data, we adapt it for HAR by inputting each channel’s data separately into Chronos.
+The resulting sensor embeddings for all channels are then concatenated and fed into an MLP, which acts as a classifier.
+This setup allows us to benchmark against a simpler framework and validate the unique contributions of SensorLLM’s design.
+GPT4TS (Zhou et al., 2023a).
+GPT4TS is a unified framework that leverages a frozen pre-trained language model (e.g., GPT-2 (Radford et al., 2019)) to achieve state-of-the-art or comparable performance across various time-series analysis tasks, including classification, forecasting (short/long-term), imputation, anomaly detection, and few-shot/zero-sample forecasting.
+The authors also found that self-attention functions similarly to PCA, providing a theoretical explanation for the versatility of transformers.
+A.8 Evaluation Metrics for Task-Aware Tuning Stage
 In our evaluation, we use the F1-macro score to assess the model’s performance across datasets.
@@
 The precision and recall for each class are defined as:
-$\text{Precision}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}}$
-$\text{Recall}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FN}_{i}}$
+$\text{Precision}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}}$ (10)
+$\text{Recall}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FN}_{i}}$ (11)
 where $\text{TP}_{i}$, $\text{FP}_{i}$, and $\text{FN}_{i}$ represent the number of true positives, false positives, and false negatives for class $i$, respectively.
 This metric ensures that performance is evaluated fairly across all classes, regardless of the frequency of each label, making it a robust measure for imbalanced datasets.
-A.8 Sensor-Language Alignment Stage Output Examples
+A.9 Sensor-Language Alignment Stage Output Examples
 Tables LABEL:tab:stage1_ouput_1 and  LABEL:tab:stage1_ouput_2 present two examples of the trend analysis results generated by SensorLLM and GPT-4o based on the input sensor data.
@@
 GPT-4o Evaluation: 5#The model’s description matches the ground truth accurately.
-GPT-4o’s output: 0.0s to 0.04s: consistent
-0.04s to 0.06s: downward
-0.06s to 0.1s: consistent
-0.1s to 0.12s: downward
-0.12s to 0.18s: consistent
-0.18s to 0.2s: upward
-0.2s to 0.24s: consistent
-Number of consistent segments: 4
-Number of downward segments: 2
-Number of upward segments: 1
-The sensor data illustrates normalized right-lower-arm x-axis gyroscope sensor readings between 0.0 and 0.24 seconds.
-The input data displays three individual trends, with a comprehensive change count reaching 7.
-To encapsulate, the data’s downward trend spanned a combined duration of 0.04 seconds, and then an upward pattern for a sum of 0.02 seconds, and a consistent trend for an accumulated time of 0.18 seconds.
-The overarching trend is characterized as downward.
+GPT-4o’s output: 0.0s to 0.06s: steady
+0.06s to 0.12s: falling
+0.12s to 0.14s: steady
+0.14s to 0.16s: rising
+0.16s to 0.18s: steady
+Total steady segments: 3
+Total rising segments: 1
+Total falling segments: 1
+The normalized right-lower-arm x-axis gyroscope sensor readings recorded within the 0.0 to 0.18 second timeframe are presented in this sensor data.
+The input data displays three individual trends, with a comprehensive change count reaching 5.
+The examination reveals that the data’s falling inclination endured for an aggregate of 0.06 seconds, succeeded by a rising trend for a cumulative period of 0.02 seconds, and a steady pattern for a total of 0.10 seconds.
+The dominant trend is steady.
 GPT-4o Evaluation: 2#Significant discrepancies in trend durations and counts compared to ground-truth.

### v3

@@
 Abstract
-We introduce SensorLLM, a two-stage framework that enables Large Language Models (LLMs) to perform human activity recognition (HAR) from sensor data.
-Despite their strong reasoning and generalization capabilities, LLMs remain underutilized for motion sensor data due to the lack of semantic context in time-series, computational constraints, and challenges in processing numerical inputs.
-SensorLLM addresses these limitations through a Sensor-Language Alignment stage, where we introduce special tokens for each sensor channel and automatically generate textual trend descriptions.
-This alignment enables LLMs to capture numerical variations, channel-specific features, and data of varying durations—without requiring human annotations.
-In the subsequent Task-Aware Tuning stage, we refine the model for HAR classification, achieving performance that matches or surpasses state-of-the-art methods.
-Our results demonstrate that SensorLLM evolves into an effective sensor learner, reasoner, and classifier through Sensor-Language Alignment, generalizing across diverse HAR datasets.
-We believe this work establishes a foundation for future research on time-series and text alignment, paving the way for foundation models in sensor data analysis.
+We introduce SensorLLM, a two-stage framework that enables Large Language Models (LLMs) to perform human activity recognition (HAR) from wearable sensor data.
+While LLMs excel at reasoning and generalization, they struggle with time-series inputs due to limited semantic context, numerical complexity, and sequence variability.
+To address these challenges, we construct SensorQA, a question-answering dataset of human-intuitive sensor-text pairs spanning diverse HAR scenarios.
+It supervises the Sensor-Language Alignment stage, where the model aligns sensor inputs with trend descriptions.
+Special tokens are introduced to mark channel boundaries.
+This alignment enables LLMs to interpret numerical patterns, channel-specific signals, and variable-length inputs—without requiring human annotation.
+In the subsequent Task-Aware Tuning stage, we adapt the model for multivariate HAR classification, achieving performance that matches or exceeds state-of-the-art methods.
+Our results show that, guided by human-intuitive alignment, SensorLLM becomes an effective sensor learner, reasoner, and classifier—generalizing across varied HAR settings and paving the way for foundation model research in time-series analysis.
 Our codes are available at https://github.com/zechenli03/SensorLLM.
-[Uncaptioned image] SensorLLM: Aligning Large Language Models with Motion Sensors for Human Activity Recognition
+[Uncaptioned image] SensorLLM: Human-Intuitive Alignment of Multivariate Sensor Data with LLMs for Activity Recognition
 Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D.
@@
 Figure 1: SensorLLM can analyze and summarize trends in captured sensor data, facilitating human activity recognition tasks.
-Human Activity Recognition (HAR) is a time-series classification task that maps sensor signals, such as accelerometer and gyroscope data, to human activities.
+Human Activity Recognition (HAR) is a time-series (TS) classification task that maps sensor signals, such as accelerometer and gyroscope data, to human activities.
 Traditional models like LSTM (Guan and Plötz, 2017; Hammerla et al., 2016) and DeepConvLSTM (Ordóñez and Roggen, 2016) learn high-level features but are task-specific and struggle to generalize across different sensor configurations and activity sets.
-In contrast, Large Language Models (LLMs) (Han et al., 2021) have shown remarkable success in integrating diverse data types (Wu et al., 2023b; Yin et al., 2023), including text and images.
-Enabling LLMs to process sensor data (Jin et al., 2023) requires either (1) pretraining or fine-tuning on time-series data (Zhou et al., 2023a), which demands substantial computational resources and is hindered by limited and imbalanced labeled data, or (2) leveraging zero-shot and few-shot prompting by converting sensor data into text (Kim et al., 2024; Ji et al., 2024).
-The latter approach avoids retraining but introduces key challenges: (i) Numerical encoding issues—LLM tokenizers, designed for text, struggle with numerical values, treating consecutive numbers as independent tokens (Nate Gruver and Wilson, 2023) and failing to preserve temporal dependencies (Spathis and Kawsar, 2024).
-(ii) Sequence length constraints—sensor data often exceeds LLMs’ maximum context length, leading to truncation, information loss, and increased computational costs.
-(iii) Multi-channel complexity—LLMs process univariate inputs, making it difficult to encode multi-sensor data in a way that retains inter-channel dependencies.
+In contrast, Large Language Models (LLMs) (Han et al., 2021) have shown remarkable success in integrating diverse data types (Liu et al., 2023a; Wu et al., 2023b; Yin et al., 2023), including text and images.
+Enabling LLMs to process wearable sensor data (Jin et al., 2023) requires either (1) pretraining or fine-tuning on TS data (Zhou et al., 2023a), which demands substantial computational resources and is hindered by limited and imbalanced labeled data, or (2) leveraging zero-shot and few-shot prompting by converting sensor data into text (Kim et al., 2024; Ji et al., 2024).
+The latter approach avoids retraining but introduces key challenges: (i) Numerical encoding issues—language model tokenizers, designed for text, struggle with numerical values, treating consecutive numbers as independent tokens (Nate Gruver and Wilson, 2023) and failing to preserve temporal dependencies (Spathis and Kawsar, 2024).
+(ii) Sequence length constraints—complex numerical sequence often exceeds LLMs’ maximum context length, leading to truncation, information loss, and increased computational costs.
+(iii) Multivariate complexity—LLMs process univariate inputs, making it difficult to encode multivariate sensor data in a way that retains inter-channel dependencies.
 (iv) Prompt engineering challenges—designing effective prompts that enable LLMs to interpret numerical sensor readings, detect trends and classify activities remains a challenge (Liu et al., 2023b).
-To address these challenges, we propose SensorLLM, a framework that aligns sensor data with natural language, allowing LLMs to analyze sensor data through text-based interactions (see Figure 1) without modifying the LLM itself.
-A major challenge is the annotation bottleneck, further compounded by the complexity and heterogeneity of sensor data.
-Unlike image-text pairs, sensor data comprises multi-channel numerical signals with diverse characteristics, making direct interpretation and alignment difficult.
+To address these challenges, we introduce SensorLLM, a human-intuition inspired framework that aligns numerical sensor data with natural language.
+Unlike image-text pairs, sensor data consists of multivariate signals with complex and varied patterns, making annotation particularly difficult.
+To tackle this, we release SensorQA, a question–answer (QA) dataset of aligned sensor-text pairs spanning diverse HAR tasks.
+SensorLLM is trained on this dataset to learn intuitive mappings between sensor readings and descriptive language, enabling LLMs to interpret sensor data through natural interactions—without any modification to the LLM architecture (see Figure 1).
 Existing methods (Jin et al., 2024a; Sun et al., 2024a) have explored condensed text prototypes for alignment, but these approaches often lack interpretability and require extensive tuning to select suitable prototypes.
-In contrast, we propose an automatic text generation approach that aligns with human intuition by deriving descriptive trend-based text directly from time-series data using statistical analyses and predefined templates.
+In contrast, we propose an automatic text generation approach that aligns with human intuition by deriving descriptive trend-based text directly from TS data using statistical analyses and predefined templates.
 This method is precise, scalable, and interpretable, eliminating the need for manual annotations while preserving essential sensor characteristics.
@@
 Sensor-Language Alignment Stage.
-We automatically generate question-answer pairs to align sensor data with text while preserving temporal features using a pretrained encoder.
-The resulting embeddings are mapped into a space interpretable by the LLM, mitigating issues associated with text-specific tokenization.
-Additionally, we introduce special tokens for sensor channels, enabling LLMs to effectively capture multi-channel dependencies.
+Our SensorQA pairs each sensor window with QA pairs to align sensor features with natural language in a structured format.
+A pretrained TS encoder extracts temporal features from the sensor data, which are then projected into a space interpretable by the LLM, mitigating text-specific tokenization issues.
+We also introduce special tokens for each sensor channel, enabling LLMs to effectively capture multivariate (multi-channel) dependencies.
 Task-Aware Tuning Stage.
-The aligned embeddings are utilized for HAR, leveraging the LLM’s reasoning capabilities while keeping its parameters frozen.
-This design extends LLMs beyond their original training, addressing concerns raised by Tan et al.
-(2024) regarding their applicability to time-series data.
-To our knowledge, this is the first approach to integrate sensor data into LLMs for sensor data analysis and HAR tasks.
-To our knowledge, this is the first approach to integrate sensor data into LLMs for sensor data analysis and HAR tasks. The key contributions of this work are as follows:
-  • We propose a fully automated, human-intuitive approach for aligning time-series data with descriptive text, eliminating the need for manual annotations.
-Using text similarity metrics, human evaluations, and LLM-based assessments, we demonstrate that SensorLLM effectively captures temporal patterns and channel-specific features, enabling robust multimodal understanding.
+The aligned embeddings are used for HAR, leveraging the LLM’s reasoning capabilities while keeping its parameters frozen.
+Importantly, our framework naturally supports sensor inputs of varying sequence lengths and arbitrary numbers of channels (i.e., multivariate TS data), a flexibility that prior approaches have struggled to accommodate.
+To the best of our knowledge, this is the first approach to integrate sensor data into LLMs for sensor-based analysis and activity recognition. The key contributions of this work are:
+  • We release SensorQA, a novel human-intuitive dataset of aligned sensor-text QA pairs covering diverse HAR scenarios.
+We evaluate SensorLLM through text similarity metrics, human judgments, and LLM-based assessments, confirming its ability to capture temporal patterns for robust multimodal understanding.
+SensorQA and SensorLLM support sensor inputs with varying sequence lengths and channel configurations, allowing broad and realistic HAR evaluation.
   • SensorLLM achieves competitive results across five HAR datasets, matching or surpassing state-of-the-art models.
@@
 3 Methods
-In this work, we propose SensorLLM, a framework that aligns sensor data with descriptive text through automatically generated question–answer pairs.
-Our goal is to develop a multimodal model with reasoning capabilities for analyzing wearable sensor data.
-As shown in Figure 2, SensorLLM comprises three core components: (1) a pretrained LLM, (2) a pretrained time-series (TS) embedder, and (3) an alignment module MLP, all operating in a two-stage framework.
-In the Sensor-Language Alignment Stage, a generative model aligns sensor readings with text based on user instructions (Liu et al., 2023a) and questions.
-In the Task-Aware Tuning Stage, a classifier is added on top of the LLM to perform HAR.
-Crucially, only the MLP in both stages and the classifier in the Task-Aware Tuning Stage are trainable, while the backbone LLM and TS embedder remain frozen.
-This design yields a highly efficient and lightweight training process, requiring only 5.67% (535.9M) of the parameters to be trainable in the alignment stage and 0.12% (10.5M) in the tuning stage.
+In this work, we propose SensorLLM, a two-stage framework that aligns wearable sensor data with descriptive text by a high-precision corpus of question-answering pairs created without any human annotations and tailored for wearable sensor reasoning.
+Our aim is to build a multimodal model capable of interpreting and reasoning over time-series (TS) signals.
+As shown in Figure 2, SensorLLM consists of three core components: (1) a pretrained LLM, (2) a pretrained TS embedder, and (3) a lightweight MLP alignment module.
+In the Sensor–Language Alignment stage, a generative model aligns sensor readings with text, and in the Task–Aware Tuning stage, a lightweight classifier is added on top of the LLM to perform HAR.
+Crucially, only the alignment MLP and this classifier are trainable—both the backbone LLM and the TS embedder remain frozen—resulting in just 5.67% (535.9 M) of parameters being fine-tuned in the first stage and 0.12% (10.5 M) in the second, making training extremely efficient.
 Figure 2: Our proposed SensorLLM framework: (a) Sensor-Language Alignment Stage, where a generative model aligns sensor readings with automatically generated text; (b) Task-Aware Tuning Stage, where a classification model leverages the aligned modalities to perform HAR.
-3.1 Sensor-Text Data Generation
-Aligning time-series data with text for LLM-based tasks is challenging due to the lack of rich semantic labels beyond class annotations, making manual annotation impractical (Deldari et al., 2024; Haresamudram et al., 2024).
-While prior works rely on predefined text prototypes (Sun et al., 2024b; Jin et al., 2024a), we aim for a more human-intuitive representation of sensor data.
-We argue that time-series data inherently contains semantic patterns that can be expressed through descriptive text, from simple numerical trends to statistical insights.
-To achieve this, we automatically generate descriptive text by analyzing observed trends and fluctuations in the data.
-Using predefined templates, we construct diverse question-answer (QA) pairs that capture trend changes while ensuring accuracy and scalability.
-These templates (Appendix A.2) are randomly combined to enhance diversity.
+3.1 SensorQA Dataset
+Aligning TS data with natural language is challenging due to the lack of rich semantic annotations beyond class labels, making manual labeling costly and impractical (Deldari et al., 2024; Haresamudram et al., 2024).
+While prior works often rely on fixed text prototypes (Sun et al., 2024b; Jin et al., 2024a), we introduce SensorQA —a scalable and human-intuitive dataset designed to bridge sensor data and language through structured QA pairs.
+SensorQA is built on the idea that TS signals naturally exhibit semantic patterns—such as trends and statistical behaviors—that can be described in natural language.
+Using predefined templates (Appendix A.2), descriptive QA pairs are automatically generated for each sensor window without any human annotations.
+Each pair includes information such as sensor type, time range, and observed trends.
+Templates are randomly combined to enhance diversity.
 For example:
@@
 3.2 Sensor-Language Alignment
-As shown in Figure 2 (a), the Sensor-Language Alignment stage employs a generative model to create multimodal sentences that combine single-channel sensor readings with textual descriptions.
-The sensor data is represented as a matrix $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ is the number of sensor channels and $T$ is the sequence length.
-Each channel’s data, denoted as $\mathbf{X}^{c}$ for channel $c$, is processed independently to preserve its unique characteristics.
-The data is then divided into non-overlapping segments, $\mathbf{X}_{S}^{c}$, where $S$ is the total number of segments.
-Each segment $x_{s}$ is assigned a random length $l$ within a predefined range, allowing the model to learn from varying temporal patterns and trend variations.
-This segmentation strategy ensures that both long-term trends and short-term fluctuations are effectively captured in the generated multimodal sentences.
-We use Chronos (Ansari et al., 2024) as the TS embedder, generating segment embeddings $\hat{x}_{s}\in\mathbb{R}^{(l+1)\times d_{ts}}$, where $d_{ts}$ is the feature dimension, and $(l+1)$ accounts for the [EOS] token appended during Chronos tokenization.
-Prior to inputting sensor segments into Chronos (Appendix A.3), we apply instance normalization $\tilde{x}_{s}=\frac{x_{s}-\text{mean}(x_{s})}{\text{std}(x_{s})}$ to standardize the data.
-Llama3-8B (Touvron et al., 2023) serves as our LLM backbone.
+As shown in Figure 2 (a), the Sensor–Language Alignment stage uses a generative model to create multimodal sentences by combining single-channel sensor readings with text from our SensorQA dataset.
+The sensor data is represented as a matrix $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels and $T$ is the sequence length.
+Each channel’s data, denoted as $\mathbf{X}^{c}$, is processed independently to retain channel-specific characteristics.
+The data is segmented into non-overlapping segments $\mathbf{X}_{S}^{c}$, where $S$ is the total number of segments.
+Each segment $x_{s}$ is assigned a random length $l$ within a predefined range, encouraging the model to learn from both short-term fluctuations and long-term trends.
+We use Chronos (Ansari et al., 2024) as the TS encoder to generate segment embeddings $\hat{x}_{s}\in\mathbb{R}^{(l+1)\times d{ts}}$, where $d_{ts}$ is the embedding dimension and $(l+1)$ accounts for the [EOS] token added during Chronos tokenization (Appendix A.3).
+Before feeding segments into Chronos, we apply instance normalization: $\tilde{x}_{s}=\frac{x_{s}-\text{mean}(x_{s})}{\text{std}(x_{s})}$.
+For the language backbone, we use LLaMA3-8B (Touvron et al., 2023).
 Alignment Module.
@@
 Loss is computed only on generated tokens, ensuring SensorLLM effectively integrates sensor and text embeddings to produce coherent, contextually appropriate responses.
-3.3 Task-Aware Tuning
-As shown in Figure 2 (b), the Task-Aware Tuning stage refines the multimodal sensor-text embeddings for HAR.
-This stage integrates multi-channel sensor readings with activity labels, aligning temporal patterns with human activities.
-The input sensor data $\mathbf{X}$ is segmented into overlapping windows of size $L$ with a 50% overlap (Li et al., 2018), forming segments $\mathbf{X}_{S}\in\mathbb{R}^{S\times C\times L}$, where $S$ is the number of segments and $C$ is the number of channels.
-The pretrained alignment module from the first stage maps sensor data to activity labels, preserving inter-channel dependencies while learning activity-related patterns.
-Input Embedding.
-For each sensor channel $c$, we retrieve its aligned sensor embeddings $\hat{o}_{s}^{c}$.
-These embeddings are then concatenated across all channels, along with their corresponding statistical features (mean and variance), to form the final input embedding:
-$\hat{z}=\hat{o}_{s}^{1}\oplus\hat{o}_{s}^{2}\oplus\dots\oplus\hat{o}_{s}^{C}\oplus\hat{z}_{\text{stat}},$ (3)
-where $\hat{z}_{\text{stat}}$ represents the statistical information, and $C$ is the number of channels. This ensures the model integrates both temporal and statistical characteristics for HAR.
-Loss Function.
-The input token sequence is processed by the LLM, yielding a latent representation $\mathbf{H}\in\mathbb{R}^{K\times D}$, where $K$ is the number of tokens and $D$ is the embedding dimension.
-Due to causal masking, we extract the final hidden state, $\mathbf{h}=\mathbf{H}_{K}$, which encodes all preceding token information.
-This pooled vector is passed through a fully connected layer to produce a prediction vector of size $M$, where $M$ is the number of activity classes.
-The final class probabilities $\hat{y}_{i}$ are obtained via the softmax function, and the model is optimized using cross-entropy loss:
-$\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i},$ (4)
-where $y_{i}$ is the ground truth label.
-4 Experiments
-In this section, we evaluate SensorLLM in enabling LLMs to interpret, reason about, and classify sensor data for HAR tasks.
-All experiments are conducted on NVIDIA A100-80G GPUs.
-To assess the LLM’s ability to learn and generalize from raw sensor inputs, we ensure that the same training and testing subjects are used in both the Sensor-Language Alignment and Task-Aware Tuning stages.
-This guarantees that test data in the second stage remains unseen during alignment, ensuring a fair evaluation of generalization.
-We select Chronos as the TS embedder because it has not been pre-trained on motion sensor data, making it an ideal candidate for evaluating our approach’s robustness in learning directly from unprocessed sensor signals.
-4.1 Datasets
-To evaluate the effectiveness and generalizability of SensorLLM, we conduct experiments on five publicly available HAR datasets: USC-HAD (Zhang and Sawchuk, 2012), UCI-HAR (Anguita et al., 2013), PAMAP2 (Reiss and Stricker, 2012), MHealth (Baños et al., 2014), and CAPTURE-24 (Chan et al., 2024).
-These datasets differ in sensor placement, sampling rates, channel configurations, and activity types, covering both controlled laboratory conditions and free-living environments.
-Additionally, they vary in scale, with USC-HAD, UCI-HAR, PAMAP2, and MHealth collected from a limited number of subjects, whereas CAPTURE-24 features a large-scale dataset with sensor recordings from 151 participants in real-world settings.
-Full dataset details, including subject count, sensor configurations, data splits, activity classes, preprocessing steps, and windowing strategies, are provided in Appendix A.6.
 Metric USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
@@
 Human 2.10 4.16 1.94 4.04 2.38 4.70 1.74 4.56 2.30 3.10
-Table 1: Evaluation of Sensor Data Trend Analysis Tasks for SensorLLM and GPT-4o.
-The assessment includes human and GPT-4o ratings (from 1 to 5, with 5 being the highest), as well as BLEU-1, ROUGE-1, ROUGE-L, METEOR, SBERT, and SimCSE (in %).
-The column GPT-4o refers to the trend analysis generated by GPT-4o itself, while the row GPT-4o refers to GPT-4o’s evaluation of the generated outputs.
+Table 1: Evaluation of Sensor Data Understanding tasks.
+The column GPT-4o denotes trend descriptions generated by GPT-4o, while the row GPT-4o indicates evaluations conducted by GPT-4o on the model outputs.
+3.3 Task-Aware Tuning
+As shown in Figure 2 (b), the Task-Aware Tuning stage refines the multimodal sensor-text embeddings for HAR.
+This stage integrates multi-channel sensor readings with activity labels, aligning temporal patterns with human activities.
+The input sensor data $\mathbf{X}$ is segmented into overlapping windows of size $L$ with a 50% overlap (Li et al., 2018), forming segments $\mathbf{X}_{S}\in\mathbb{R}^{S\times C\times L}$, where $S$ is the number of segments and $C$ is the number of channels.
+The pretrained alignment module from the first stage maps sensor data to activity labels, preserving inter-channel dependencies while learning activity-related patterns.
+Input Embedding.
+For each sensor channel $c$, we retrieve its aligned embeddings $\hat{o}_{s}^{c}$ from the pretrained alignment module.
+These are concatenated across all channels, along with statistical features (mean and variance) from our SensorQA dataset, to form the final input embedding:
+$\hat{z}=\hat{o}_{s}^{1}\oplus\hat{o}_{s}^{2}\oplus\dots\oplus\hat{o}_{s}^{C}\oplus\hat{z}_{\text{stat}},$ (3)
+where $\hat{z}_{\text{stat}}$ represents the statistical information. This ensures the model integrates both temporal and statistical characteristics for HAR.
+Loss Function.
+The input token sequence is processed by the LLM, yielding a latent representation $\mathbf{H}\in\mathbb{R}^{K\times D}$, where $K$ is the number of tokens and $D$ is the embedding dimension.
+Due to causal masking, we extract the final hidden state, $\mathbf{h}=\mathbf{H}_{K}$, which encodes all preceding token information.
+This pooled vector is passed through a fully connected layer to produce a prediction vector of size $M$, where $M$ is the number of activity classes.
+The final class probabilities $\hat{y}_{i}$ are obtained via the softmax function, and the model is optimized using cross-entropy loss:
+$\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i},$ (4)
+where $y_{i}$ is the ground truth label.
+Method USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
+F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy
+PatchTST 45.2±$\pm 1.48$ 45.6±$\pm 2.19$ 86.8±$\pm 0.84$ 86.0±$\pm 0.71$ 82.0±$\pm 0.71$ 81.2±$\pm 0.84$ 80.0±$\pm 1.58$ 79.4±$\pm 1.34$ 35.6±$\pm 0.89$ 66.2±$\pm 1.10$
+Ns-Transformer 52.6±$\pm 2.30$ 51.8±$\pm 2.86$ 88.0±$\pm 0.71$ 87.4±$\pm 0.55$ 78.8±$\pm 0.84$ 78.8±$\pm 0.84$ 77.2±$\pm 1.48$ 75.8±$\pm 1.48$ 34.8±$\pm 1.10$ 65.4±$\pm 0.55$
+Informer 51.2±$\pm 1.30$ 51.6±$\pm 1.52$ 86.6±$\pm 1.14$ 86.4±$\pm 0.89$ 78.0±$\pm 1.58$ 78.6±$\pm 1.34$ 74.0±$\pm 0.71$ 72.8±$\pm 0.84$ 35.6±$\pm 0.55$ 66.8±$\pm 0.84$
+Transformer 49.6±$\pm 1.67$ 50.6±$\pm 0.55$ 85.4±$\pm 0.89$ 85.2±$\pm 1.10$ 77.0±$\pm 0.71$ 77.6±$\pm 0.89$ 75.2±$\pm 1.30$ 74.6±$\pm 1.34$ 32.8±$\pm 0.84$ 65.4±$\pm 0.89$
+iTransformer 48.4±$\pm 1.82$ 49.6±$\pm 1.67$ 81.8±$\pm 0.84$ 81.8±$\pm 0.84$ 76.6±$\pm 0.55$ 75.8±$\pm 0.45$ 80.4±$\pm 1.14$ 80.0±$\pm 1.22$ 19.8±$\pm 0.84$ 62.4±$\pm 0.89$
+TimesNet 52.2±$\pm 2.39$ 52.6±$\pm 2.07$ 87.4±$\pm 1.14$ 86.6±$\pm 1.14$ 76.2±$\pm 1.92$ 77.4±$\pm 1.14$ 78.4±$\pm 1.52$ 77.2±$\pm 1.48$ 34.8±$\pm 0.84$ 65.8±$\pm 1.79$
+GPT4TS 54.2±$\pm 2.05$ 56.0±$\pm 1.58$ 88.2±$\pm 0.84$ 87.6±$\pm 0.55$ 80.4±$\pm 0.89$ 79.8±$\pm 0.45$ 76.4±$\pm 1.14$ 75.4±$\pm 1.14$ 32.8±$\pm 1.10$ 62.2±$\pm 1.92$
+Chronos+MLP 44.2±$\pm 1.30$ 44.0±$\pm 0.71$ 82.2±$\pm 0.84$ 81.2±$\pm 0.84$ 79.8±$\pm 0.45$ 79.8±$\pm 0.45$ 83.0±$\pm 0.71$ 82.0±$\pm 0.71$ 38.0±$\pm 0.71$ 68.2±$\pm 0.84$
+DeepConvLSTM 48.8±$\pm 2.39$ 50.6±$\pm 2.41$ 89.2±$\pm 0.84$ 89.2±$\pm 0.84$ 78.4±$\pm 1.52$ 78.2±$\pm 1.10$ 75.0±$\pm 1.87$ 76.0±$\pm 1.00$ 40.4±$\pm 0.89$ 69.4±$\pm 1.14$
+DeepConvLSTMAtt 54.0±$\pm 2.12$ 54.4±$\pm 3.21$ 89.6±$\pm 1.14$ 89.4±$\pm 1.14$ 79.2±$\pm 1.30$ 79.6±$\pm 1.14$ 77.4±$\pm 2.19$ 76.8±$\pm 1.48$ 41.4±$\pm 0.55$ 70.4±$\pm 0.55$
+Attend 60.2±$\pm 2.17$ 60.8±$\pm 1.92$ 93.2±$\pm 0.84$ 92.8±$\pm 0.45$ 84.6±$\pm 1.14$ 85.0±$\pm 0.71$ 83.4±$\pm 1.14$ 82.6±$\pm 1.14$ 43.6±$\pm 0.55$ $\pm 0.71$
+SensorLLM 61.2±$\pm 3.56$ 62.6±$\pm 3.36$ 91.2±$\pm 1.48$ 90.8±$\pm 1.30$ 86.2±$\pm 1.48$ 87.2±$\pm 0.84$ 89.4±$\pm 3.85$ 89.0±$\pm 3.54$ 48.6±$\pm 1.14$ 72.0±$\pm 0.71$
+Table 2: F1-macro and accuracy scores (%) for the Human Activity Recognition tasks, presented as the mean and standard deviation over 5 random repetitions.
+Bold for the best and underline for the second-best.
+4 Experiments
+In this section, we evaluate SensorLLM in enabling LLMs to interpret, reason about, and classify sensor data for HAR tasks.
+All experiments are conducted on NVIDIA A100-80G GPUs.
+To assess the LLM’s ability to learn and generalize from raw sensor inputs, we ensure that the same training and testing subjects are used in both the Sensor-Language Alignment and Task-Aware Tuning stages.
+This guarantees that test data in the second stage remains unseen during alignment, ensuring a fair evaluation of generalization.
+We select Chronos as the TS embedder since it has not been pre-trained on HAR-specific data, allowing us to evaluate the robustness of our approach in adapting to raw, domain-agnostic sensor signals.
+4.1 Datasets
+To evaluate the effectiveness and generalizability of SensorLLM, we conduct experiments on five publicly available HAR datasets: USC-HAD (Zhang and Sawchuk, 2012), UCI-HAR (Anguita et al., 2013), PAMAP2 (Reiss and Stricker, 2012), MHealth (Baños et al., 2014), and CAPTURE-24 (Chan et al., 2024).
+These datasets vary widely in subject counts, sensor placement, sampling rates, channel configurations, and activity types, covering both controlled laboratory conditions and free-living environments.
+SensorQA (see Appendix A.7) is built on these five benchmarks and supports both alignment and activity classification.
+All datasets are publicly available, containing no personally identifiable information, thus posing minimal ethical or privacy concerns.
+We use subject-independent splits for all datasets except UCI-HAR, which comes with a fixed split.
+In all other datasets, training and test sets come from different subjects, ensuring the model is evaluated on unseen users.
+Full dataset details, including subject count, sensor configurations, data splits, activity classes, preprocessing steps, and windowing strategies, are provided in Appendix A.6.
 4.2 Sensor Data Understanding
 Setup.
-We train SensorLLM on all five datasets using the same following parameters in the Sensor-Language Alignment Stage: a learning rate of 2e-3, 8 epochs, batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 8192 for CAPTURE-24 and 4096 for the other datasets.
+All datasets are trained using the same parameters in the Sensor-Language Alignment Stage: a learning rate of 2e-3, 8 epochs, batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 8192 for CAPTURE-24 and 4096 for others.
 Evaluation Metrics.
-We assess SensorLLM’s ability to generate trend descriptions from sensor data, comparing it with the advanced GPT-4o 111gpt-4o-2024-08-06(OpenAI, 2024) to evaluate Sensor-Language Alignment.
-GPT-4o generates responses based on a predefined prompt (Appendix A.4), following our text template.
-We employ three evaluation methods:
+We assess the performance of SensorLLM in the sensor–language alignment stage by comparing its ability to generate trend descriptions from sensor data with that of the advanced GPT-4o 111gpt-4o-2024-08-06 (OpenAI, 2024).
+GPT-4o generates responses using a predefined prompt (Appendix A.4).
+We adopt three evaluation methods:
   • NLP Metrics.
-We measure surface-level similarity and n-gram overlap using BLEU-1 (Papineni et al., 2002), ROUGE-1, ROUGE-L (Lin, 2004), and METEOR (Banerjee and Lavie, 2005).
-We also use SBERT (Reimers and Gurevych, 2019) and SimCSE (Gao et al., 2021) to evaluate semantic similarity between model outputs and ground truth.
+We use BLEU-1 (Papineni et al., 2002), ROUGE-1, ROUGE-L (Lin, 2004), and METEOR (Banerjee and Lavie, 2005) to measure surface-level similarity and n-gram overlap.
+For deeper semantic alignment and factual correctness, we adopt SBERT (Reimers and Gurevych, 2019) and SimCSE (Gao et al., 2021).
   • GPT-4o Evaluation.
@@
 Results.
-Table 1 compares SensorLLM and GPT-4o on the Sensor Data Trend Analysis tasks, showing that our model consistently outperforms GPT-4o across all metrics.
-BLEU-1, ROUGE-1, ROUGE-L, and METEOR primarily focus on surface-level lexical or n-gram overlaps and SBERT and SimCSE can capture factual correctness or deeper semantic similarities.
-Across all metrics, SensorLLM generates trend descriptions more closely aligned with the ground truth.
-GPT-4o evaluations further highlight SensorLLM’s superior ability to capture trend details and coherence, whereas GPT-4o struggles with complex numerical data and trend observations (Yehudai et al., 2024).
-Human evaluation also favors SensorLLM, particularly for shorter sequences.
-CAPTURE-24 results are weaker compared to other datasets, likely due to its longer sequences being trained with the same parameters.
-Overall, these findings validate the effectiveness of our Sensor-Language Alignment method in enhancing LLMs’ ability to interpret complex numerical sequence.
-Appendix A.9 provides qualitative examples of outputs from both models.
-Method USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
-PatchTST 45.2±$\pm 1.48$ 86.8±$\pm 0.84$ 82.0±$\pm 0.71$ 80.0±$\pm 1.58$ 35.6±$\pm 0.89$
-Ns-Transformer 52.6±$\pm 2.30$ 88.0±$\pm 0.71$ 78.8±$\pm 0.84$ 77.2±$\pm 1.48$ 34.8±$\pm 1.10$
-Informer 51.2±$\pm 1.30$ 86.6±$\pm 1.14$ 78.0±$\pm 1.58$ 74.0±$\pm 0.71$ 35.6±$\pm 0.55$
-Transformer 49.6±$\pm 1.67$ 85.4±$\pm 0.89$ 77.0±$\pm 0.71$ 75.2±$\pm 1.30$ 32.8±$\pm 0.84$
-iTransformer 48.4±$\pm 1.82$ 81.8±$\pm 0.84$ 76.6±$\pm 0.55$ 80.4±$\pm 1.14$ 19.8±$\pm 0.84$
-TimesNet 52.2±$\pm 2.39$ 87.4±$\pm 1.14$ 76.2±$\pm 1.92$ 78.4±$\pm 1.52$ 34.8±$\pm 0.84$
-GPT4TS 54.2±$\pm 2.05$ 88.2±$\pm 0.84$ 80.4±$\pm 0.89$ 76.4±$\pm 1.14$ 32.8±$\pm 1.10$
-Chronos+MLP 44.2±$\pm 1.30$ 82.2±$\pm 0.84$ 79.8±$\pm 0.45$ 83.0±$\pm 0.71$ 38.0±$\pm 0.71$
-DeepConvLSTM 48.8±$\pm 2.39$ 89.2±$\pm 0.84$ 78.4±$\pm 1.52$ 75.0±$\pm 1.87$ 40.4±$\pm 0.89$
-DeepConvLSTMAtt 54.0±$\pm 2.12$ 89.6 ±$\pm 1.14$ 79.2±$\pm 1.30$ 77.4±$\pm 2.19$ 41.4±$\pm 0.55$
-Attend 60.2±$\pm 2.17$ 93.2±$\pm 0.84$ 84.6±$\pm 1.14$ 83.4±$\pm 1.14$ 43.6±$\pm 0.55$
-SensorLLM 61.2±$\pm 3.56$ 91.2±$\pm 1.48$ 86.2±$\pm 1.48$ 89.4±$\pm 3.85$ 48.6±$\pm 1.14$
-Table 2: F1-macro results (%) for the Task-Aware Tuning Stage, presented as the mean and standard deviation over 5 random repetitions.
-The top results for each dataset are highlighted as follows: Bold for the best and underline for the second-best.
+Table 1 compares SensorLLM and GPT-4o on the Sensor Data Understanding task.
+SensorLLM consistently outperforms GPT-4o across all metrics, with outputs that align more closely with ground truth and exhibit stronger trend understanding and coherence.
+In contrast, GPT-4o often struggles with complex numerical data and trend detection (Yehudai et al., 2024).
+When serving as an evaluator, GPT-4o also indicates a preference for SensorLLM.
+Human evaluation, conducted on shorter sequences, likewise favors SensorLLM.
+Performance on CAPTURE-24 is comparatively lower, likely due to longer sequences being trained under fixed parameters.
+Overall, these results validate the effectiveness of our alignment method in enabling LLMs to interpret complex TS data.
+Qualitative examples are provided in Appendix A.10.
 4.3 Human Activity Recognition
@@
 In this section, we evaluate the performance of SensorLLM on HAR tasks.
-Each experiment runs for five trials, using 8 training epochs, a batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 4096.
-We report the F1 macro score A.8 to account for class imbalance across different activity categories.
+Each experiment is run for five trials using 8 training epochs, a batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 4096.
+We report both F1-macro (Appendix A.9) and accuracy to account for class imbalance and overall prediction performance across different activity categories.
 Baselines.
@@
 Results.
-Table 2 reports the macro F1 scores (%) averaged over five random runs.
-SensorLLM achieves the highest performance on four datasets (USC-HAD, PAMAP2, MHealth, CAPTURE-24) and ranks second on UCI-HAR, demonstrating its effectiveness in handling diverse sensor data.
-Notably, SensorLLM achieves a notable improvement on the CAPTURE-24 dataset, surpassing all baselines by a significant margin with a mean F1-macro score of 48.6%, which is 5.0% higher than Attend (43.6%).
-On USC-HAD, SensorLLM achieves the highest score of 61.2%, outperforming Attend, the second-best baseline, by 1.0%.
-Similarly, on PAMAP2, SensorLLM achieves a score of 86.2%, exceeding Attend (84.6%) by 1.6%.
-On MHealth, SensorLLM sets a new state-of-the-art with a score of 89.4%, surpassing Attend (83.4%) by 6.0%.
-These results highlight SensorLLM’s ability to consistently outperform existing methods across diverse datasets.
-For UCI-HAR, SensorLLM achieves the second-best score (91.2%), slightly trailing Attend (93.2%).
-In contrast, Chronos+MLP shows only a slight improvement over iTransformer, the lowest-performing baseline (82.2% vs.
-81.8%), indicating that Chronos embeddings alone have limited utility for HAR on this dataset.
-However, our framework significantly enhances their effectiveness, highlighting the robustness of our alignment approach.
+Table 2 reports F1-macro and accuracy scores (%) averaged over five runs.
+SensorLLM achieves the best performance on four out of five datasets (USC-HAD, PAMAP2, MHealth, CAPTURE-24), and ranks second on UCI-HAR, slightly behind Attend.
+It shows notable gains on challenging datasets such as CAPTURE-24 and MHealth, demonstrating strong performance in real-world and long-sequence settings.
+Compared to Chronos+MLP, which uses the same TS encoder, SensorLLM significantly improves both F1-macro and accuracy, highlighting the effectiveness of our alignment strategy in enabling LLMs to understand and classify sensor data.
+Strong results on both F1-macro and accuracy indicate that SensorLLM performs well in both overall prediction and per-class balance, showing robust generalization across diverse sensor configurations, activity types, and data collection environments.
 5 Ablation Studies
@@
 w/o prompts w/ prompts w/o prompts w/ prompts
-USC-HAD 43.4±$\pm 2.88$ 45.0±$\pm 1.58$ 49.6±$\pm 1.67$ 61.2±$\mathbf{\pm 3.56}$
-UCI-HAR 80.0±$\pm 2.12$ 82.0±$\pm 1.58$ 89.2±$\pm 1.10$ 91.2±$\mathbf{\pm 1.48}$
-PAMAP2 74.2±$\pm 2.28$ 75.4±$\pm 3.05$ 83.0±$\pm 0.71$ 86.2±$\mathbf{\pm 1.48}$
-MHealth 76.6±$\pm 1.34$ 77.4±$\pm 3.13$ 86.6±$\pm 1.14$ 89.4±$\mathbf{\pm 3.85}$
-CAPTURE-24 44.8±$\pm 0.84$ 46.0±$\pm 0.71$ 47.2±$\pm 0.84$ 48.6±$\mathbf{\pm 1.14}$
-Table 3: The results for SensorLLM trained with/without text prompts. Task-only refers to conducting HAR directly bypassing sensor-language alignment.
-Impact of Alignment.
-To evaluate the impact of alignment, we included the Chronos+MLP baseline in Section 4.3 to show that SensorLLM’s performance is not solely driven by Chronos embeddings.
-Additionally, we compared SensorLLM with a Task-only model, which skips the Sensor-Language Alignment Stage and directly applies Chronos embeddings and the LLM for HAR.
-As shown in Table 3, SensorLLM consistently outperforms the Task-only model across all five datasets, regardless of additional textual input.
-Notably, the Task-only model often performs on par with or worse than traditional TS baselines, underscoring the importance of our alignment method.
-These results confirm that Chronos embeddings alone are insufficient for optimal HAR performance and that alignment is essential for enabling LLMs to effectively understand sensor data.
-Impact of Prompts.
+USC-HAD 43.4±$\pm 2.88$ 45.0±$\pm 1.58$ 49.6±$\pm 1.67$ 61.2±$\pm 3.56$
+UCI-HAR 80.0±$\pm 2.12$ 82.0±$\pm 1.58$ 89.2±$\pm 1.10$ 91.2±$\pm 1.48$
+PAMAP2 74.2±$\pm 2.28$ 75.4±$\pm 3.05$ 83.0±$\pm 0.71$ 86.2±$\pm 1.48$
+MHealth 76.6±$\pm 1.34$ 77.4±$\pm 3.13$ 86.6±$\pm 1.14$ 89.4±$\pm 3.85$
+CAPTURE-24 44.8±$\pm 0.84$ 46.0±$\pm 0.71$ 47.2±$\pm 0.84$ 48.6±$\pm 1.14$
+Table 3: F1-macro scores for models trained with and without text prompts. Task-only refers to conducting Task-Aware Tuning directly bypassing the alignment stage.
+Removing Alignment Hurts.
+To assess the role of sensor–language alignment, we include the Chronos+MLP baseline (Section 4.3) to demonstrate that SensorLLM ’s performance is not solely due to the strength of the Chronos encoder.
+We further compare SensorLLM with a Task-only variant that skips the Sensor–Language Alignment stage and directly feeds Chronos embeddings into the LLM for HAR.
+As shown in Table 3, SensorLLM consistently outperforms the Task-only model across all five datasets, regardless of whether textual prompts are included.
+Notably, the Task-only model often performs comparably to or worse than traditional TS baselines, underscoring the critical role of alignment.
+These results confirm that Chronos embeddings alone are insufficient for optimal HAR performance, and that our alignment stage is essential for enabling the LLM to effectively interpret sensor data.
+Textual Prompts Enhance HAR.
 To assess the role of additional textual information (e.g., statistical features for each sensor channel) in the Task-Aware Tuning Stage, we compared SensorLLM ’s performance with and without prompts.
@@
 The results highlight the benefits of multimodal inputs, which enrich sensor data representations and improve HAR accuracy.
-More broadly, the ability to process both sensor signals and textual prompts not only enhances classification performance but also extends LLMs’ potential for tackling complex sensor-driven tasks in future applications.
+More broadly, the ability to jointly process sensor data and textual prompts underscores the potential of LLMs for more generalizable and interpretable sensor-driven applications.
 Figure 3: Effect of the number of alignment module layers.
-Alignment Module Layers.
-We investigate the effect of hidden layer depth in the alignment module MLP on UCI-HAR, PAMAP2, and MHealth.
-As shown in Figure 3, increasing the number of hidden layers from one ($1024\to 2048\to 4096$) to two ($1024\to 2048\to 3072\to 4096$) led to mixed results.
-The average F1-macro scores improved on UCI-HAR (91.2% → 92.0%) and MHealth (89.4% → 90.2%), but slightly dropped on PAMAP2 (86.2% → 85.8%).
-These results suggest that the optimal number of hidden layers varies by dataset.
-Effect of Model Size on Performance.
-To assess the impact of model size, we tested SensorLLM-3b, a resource-efficient variant using Chronos-base and Llama3.2-3b.
+MLP Depth Trade-offs.
+We examine how the depth of the alignment module MLP affects performance on UCI-HAR, PAMAP2, and MHealth.
+As shown in Figure 3, increasing the number of hidden layers from one ($1024\to 2048\to 4096$) to two ($1024\to 2048\to 3072\to 4096$) yields mixed results.
+F1-macro scores improve on UCI-HAR and MHealth, but slightly decrease on PAMAP2.
+These findings suggest that deeper MLPs do not always improve performance, and a single hidden layer offers a good balance between accuracy and efficiency.
+Smaller SensorLLM Still Compete.
+To address computational feasibility for deployment in resource-constrained environments, we evaluate SensorLLM-3b—a lighter variant built with Chronos-base and LLaMA3.2-3b.
 Experiments were conducted on USC-HAD, UCI-HAR, and MHealth.
-As shown in Figure 4, SensorLLM-3b achieves slightly lower performance than SensorLLM-8b across all datasets, illustrating the trade-off between model size and accuracy.
-However, SensorLLM-3b remains competitive, outperforming Attend on USC-HAD and MHealth, while trailing it only on UCI-HAR.
-These results indicate that SensorLLM-3b offers a resource-efficient alternative, maintaining strong performance relative to other baselines.
+As shown in Figure 4, SensorLLM-3b achieves slightly lower performance than SensorLLM-8b, reflecting the trade-off between model size and accuracy.
+Nevertheless, it remains competitive—outperforming Attend on USC-HAD and MHealth, and closely trailing it on UCI-HAR.
+These results suggest that SensorLLM-3b provides a strong balance between efficiency and performance, making it a viable choice for real-world, resource-limited applications.
 Figure 4: Effect of Model Size.
-Cross-Dataset Generalization.
-To assess the robustness of SensorLLM, we conducted cross-dataset experiments, training the Sensor-Language Alignment Stage on USC-HAD and the Task-Aware Tuning Stage on UCI-HAR, and vice versa.
-These datasets share the same sensor channels but differ in sampling rates.
-As shown in Table 4, SensorLLM achieves performance comparable to training both stages on the same dataset.
-This indicates that once modality alignment is established, retraining on new datasets is unnecessary.
-SensorLLM effectively generalizes to downstream tasks across diverse datasets, demonstrating that the alignment stage enables the LLM to truly understand sensor data, rather than memorizing dataset-specific features.
-These results highlight SensorLLM’s potential for broad cross-dataset generalization, laying the groundwork for future TS-LLM models.
-Stage 1 Stage 2 Results
-USC-HAD UCI-HAR 91.0±$\mathbf{\pm 1.41}$
-UCI-HAR USC-HAD 61.6±$\mathbf{\pm 2.07}$
-Table 4: Cross-dataset experiments.
+Dataset F1-macro # Channels
+w/o ST w/ ST
+MHealth 89.6±$\pm 2.70$ 90.2±$\pm 3.11$ 15
+PAMAP2 84.4±$\pm 1.14$ 85.8±$\pm 0.84$ 27
+Table 4: Effect of special tokens on HAR based on two-layer alignment MLP. ST refers to special tokens.
+Special Tokens Improve Performance.
+We investigate the role of special tokens in helping SensorLLM distinguish sensor data from text and identify different sensor channel types.
+Special tokens are added to the aligned embeddings of each sensor channel and act as learned identifiers.
+They provide structural cues that help the LLM model channel-wise dependencies and reduce modality confusion.
+We conduct experiments on PAMAP2 and MHealth, both of which contain multiple sensor channels.
+As shown in Table 4, removing special tokens leads to a slight drop in F1-macro scores, with the performance gap tending to widen as the number of sensor channels increases.
+This confirms their value in preserving positional and channel-level structure within a flat token sequence.
+Alignment Enables Generalization.
+To assess the robustness of SensorLLM, we conduct cross-dataset experiments by training the Sensor–Language Alignment Stage on USC-HAD and the Task-Aware Tuning Stage on UCI-HAR, and vice versa.
+While these datasets share the same sensor channels, they differ in sensor wearing position, sampling rates and activity distributions.
+As shown in Table 5, SensorLLM achieves performance comparable to models trained entirely on the same dataset.
+This suggests that once modality alignment is learned, it can be transferred across datasets without retraining.
+These results indicate that SensorLLM does not overfit to dataset-specific patterns but learns generalizable sensor-language representations, demonstrating strong cross-dataset adaptability and paving the way for more universal TS–LLM frameworks.
+Stage 2 Stage 1 F1-macro
+USC-HAD UCI-HAR 61.6±$\pm 2.07$
+USC-HAD 61.2±$\pm 3.56$
+UCI-HAR UCI-HAR 91.2±$\pm 1.48$
+USC-HAD 91.0±$\pm 1.41$
+Table 5: F1-macro scores for cross-dataset experiments.
 6 Conclusions
-We introduced SensorLLM, a multimodal framework that aligns sensor data with automatically generated text at a human-perception level, moving beyond machine-level alignment.
-SensorLLM effectively captures complex sensor patterns, achieving superior performance in HAR tasks.
-Experiments across diverse datasets demonstrate its robustness in handling variable-length sequences, multi-channel inputs, and textual metadata.
-Cross-dataset results further highlight its strong generalizability without requiring dataset-specific alignment.
-This work establishes a foundation for Sensor-Text MLLMs, with potential applications for sensor data analysis.
-We release our code and data generation pipeline to facilitate future research on integrating time-series and text, particularly in low-resource domains.
+We present SensorLLM, a multimodal framework that aligns sensor data with natural language through a QA format at a human-perception level, moving beyond machine-level alignment.
+It effectively captures complex sensor patterns, achieves strong performance on HAR tasks, and generalizes well without requiring dataset-specific alignment.
+Experiments demonstrate its robustness across variable-length sequences, multivariate inputs, and textual metadata.
+To support future research, we release our code and the SensorQA dataset, constructed from five public HAR benchmarks, to advance time-series and language integration, particularly in low-resource domains.
 7 Limitations
@@
 Classifier-Based Design.
-To ensure fair comparisons with existing HAR models (classification models), we adopt a classifier for downstream tasks rather than fully exploiting the LLM’s generative abilities.
-Although our results verify that the Sensor-Language Alignment Stage can generalize across datasets, relying on a fixed-class classifier may constrain the model’s adaptability to new activity categories.
-Future work could explore using the LLM in a generative or prompt-based capacity, enabling broader application scenarios such as activity discovery or open-set recognition.
+To ensure fair comparisons with existing HAR models, we adopt a classifier for downstream tasks rather than fully leveraging the LLM’s generative capabilities.
+While our results demonstrate that the Sensor–Language Alignment Stage can generalize across datasets, relying on a fixed-class classifier may limit adaptability to new activity categories.
+Although zero-shot adaptability is a valuable direction, we did not explore it here due to the lack of comparable baselines.
+To the best of our knowledge, no prior work supports generalization to unseen activity classes under variable-length and variable-channel sensor input, as our framework does, making fair comparison on zero-shot settings infeasible at this stage.
+Future work could explore generative or prompt-based approaches to support broader applications such as activity discovery or open-set recognition.
 Scope of Sensor-Text Alignment.
@@
 (2024) explored LLMs for zero-shot HAR using raw IMU data, showing that GPT-4 can outperform both traditional and deep learning models in simple HAR tasks without domain-specific adaptations, highlighting LLMs’ potential in sensor-based systems.
-A.2 Data Generation
-We generate text data from sensor readings using predefined sentence templates (5, Tables 6,, 7).
+A.2 SensorQA Generation
+For SensorQA, we generate text data from sensor readings using predefined sentence templates (Tables 6, 7, 8).
 These templates are randomly selected to create diverse question-answer (QA) pairs.
@@
 Table 11: Dataset classes and Proportions
-A.7 Baselines for Task-Aware Tuning Stage
+A.7 SensorQA
+We introduce SensorQA, a novel question-answering dataset designed to align time-series sensor data with human-interpretable natural language.
+Each sample pairs a segment of sensor input with natural language questions and answers, capturing trends, patterns, or activity-level semantics.
+Sensor-Language Alignment Stage
+of SensorQA focuses on aligning uni-variate sensor sequence of variable length with descriptive textual responses and includes two types of QA tasks:
+  • Trend Analysis QA, which describes how the signal changes within the window.
+  • Trend Summary QA, which summarizes the overall behavior across a window in a concise natural language phrase.
+Task-Aware Tuning Stage
+focuses on using multi-variate sensor sequences to perform human activity classification, leveraging the aligned representations learned in the alignment stage.
+This stage of SensorQA contains statistical information from each sensor channel as part of the input representation.
+Dataset Stage 1 Stage 2
+Train Test Train Test
+USC-HAD 300,744 58,704 22,790 4,555
+UCI-HAR 128,292 25,932 7,352 2,947
+PAMAP2 738,666 271,674 14,163 5,210
+MHealth 283,020 60,780 4771 2,039
+CAPTURE-24 72,714 35,688 61,327 30,138
+Table 12: Training and testing sample counts for Stage 1 and Stage 2 across datasets of SensorQA.
+The distribution of training and testing data across both stages is summarized in Table 12.
+A.8 Baselines for Task-Aware Tuning Stage
 In Task-Aware Tuning Stage, we compare SensorLLM against several state-of-the-art baseline models for time-series classification and human activity recognition (HAR).
@@
 The authors also found that self-attention functions similarly to PCA, providing a theoretical explanation for the versatility of transformers.
-A.8 Evaluation Metrics for Task-Aware Tuning Stage
+A.9 Evaluation Metrics for Task-Aware Tuning Stage
 In our evaluation, we use the F1-macro score to assess the model’s performance across datasets.
@@
 This metric ensures that performance is evaluated fairly across all classes, regardless of the frequency of each label, making it a robust measure for imbalanced datasets.
-A.9 Sensor-Language Alignment Stage Output Examples
+A.10 Sensor-Language Alignment Stage Output Examples
 Tables LABEL:tab:stage1_ouput_1 and  LABEL:tab:stage1_ouput_2 present two examples of the trend analysis results generated by SensorLLM and GPT-4o based on the input sensor data.

### v4

@@
-Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim2
+Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim1
 Abstract
-We introduce SensorLLM, a two-stage framework that enables Large Language Models (LLMs) to perform human activity recognition (HAR) from wearable sensor data.
-While LLMs excel at reasoning and generalization, they struggle with time-series inputs due to limited semantic context, numerical complexity, and sequence variability.
-To address these challenges, we construct SensorQA, a question-answering dataset of human-intuitive sensor-text pairs spanning diverse HAR scenarios.
-It supervises the Sensor-Language Alignment stage, where the model aligns sensor inputs with trend descriptions.
+We introduce SensorLLM, a two-stage framework that enables Large Language Models (LLMs) to perform human activity recognition (HAR) from sensor time-series data.
+Despite their strong reasoning and generalization capabilities, LLMs remain underutilized for motion sensor data due to the lack of semantic context in time-series, computational constraints, and challenges in processing numerical inputs.
+SensorLLM addresses these limitations through a Sensor-Language Alignment stage, where the model aligns sensor inputs with trend descriptions.
 Special tokens are introduced to mark channel boundaries.
-This alignment enables LLMs to interpret numerical patterns, channel-specific signals, and variable-length inputs—without requiring human annotation.
-In the subsequent Task-Aware Tuning stage, we adapt the model for multivariate HAR classification, achieving performance that matches or exceeds state-of-the-art methods.
-Our results show that, guided by human-intuitive alignment, SensorLLM becomes an effective sensor learner, reasoner, and classifier—generalizing across varied HAR settings and paving the way for foundation model research in time-series analysis.
+This alignment enables LLMs to capture numerical variations, channel-specific features, and data of varying durations, without requiring human annotations.
+In the subsequent Task-Aware Tuning stage, we refine the model for HAR classification, achieving performance that matches or surpasses state-of-the-art methods.
+Our results demonstrate that SensorLLM evolves into an effective sensor learner, reasoner, and classifier through human-intuitive Sensor-Language Alignment, generalizing across diverse HAR datasets.
+We believe this work establishes a foundation for future research on time-series and text alignment, paving the way for foundation models in sensor data analysis.
 Our codes are available at https://github.com/zechenli03/SensorLLM.
-[Uncaptioned image] SensorLLM: Human-Intuitive Alignment of Multivariate Sensor Data with LLMs for Activity Recognition
+[Uncaptioned image] SensorLLM: Aligning Large Language Models with Motion Sensors for Human Activity Recognition
 Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D.
-Salim2 1University of New South Wales, Sydney 2University of Tokyo {zechen.li, s.deldari, hao.xue1, flora.salim}@unsw.edu.au chen-linyao217@g.ecc.u-tokyo.ac.jp
+Salim1 1University of New South Wales, Sydney 2University of Tokyo {zechen.li, s.deldari, hao.xue1, flora.salim}@unsw.edu.au chen-linyao217@g.ecc.u-tokyo.ac.jp
 1 Introduction
 Figure 1: SensorLLM can analyze and summarize trends in captured sensor data, facilitating human activity recognition tasks.
-Human Activity Recognition (HAR) is a time-series (TS) classification task that maps sensor signals, such as accelerometer and gyroscope data, to human activities.
+Human Activity Recognition (HAR) is a time-series classification task that maps sensor signals, such as accelerometer and gyroscope data, to human activities.
 Traditional models like LSTM (Guan and Plötz, 2017; Hammerla et al., 2016) and DeepConvLSTM (Ordóñez and Roggen, 2016) learn high-level features but are task-specific and struggle to generalize across different sensor configurations and activity sets.
-In contrast, Large Language Models (LLMs) (Han et al., 2021) have shown remarkable success in integrating diverse data types (Liu et al., 2023a; Wu et al., 2023b; Yin et al., 2023), including text and images.
+In contrast, Large Language Models (LLMs) (Han et al., 2021) have shown remarkable success in integrating diverse data types (Liu et al., 2023a; Wu et al., 2024; Yin et al., 2024), including text and images.
 Enabling LLMs to process wearable sensor data (Jin et al., 2023) requires either (1) pretraining or fine-tuning on TS data (Zhou et al., 2023a), which demands substantial computational resources and is hindered by limited and imbalanced labeled data, or (2) leveraging zero-shot and few-shot prompting by converting sensor data into text (Kim et al., 2024; Ji et al., 2024).
-The latter approach avoids retraining but introduces key challenges: (i) Numerical encoding issues—language model tokenizers, designed for text, struggle with numerical values, treating consecutive numbers as independent tokens (Nate Gruver and Wilson, 2023) and failing to preserve temporal dependencies (Spathis and Kawsar, 2024).
-(ii) Sequence length constraints—complex numerical sequence often exceeds LLMs’ maximum context length, leading to truncation, information loss, and increased computational costs.
-(iii) Multivariate complexity—LLMs process univariate inputs, making it difficult to encode multivariate sensor data in a way that retains inter-channel dependencies.
-(iv) Prompt engineering challenges—designing effective prompts that enable LLMs to interpret numerical sensor readings, detect trends and classify activities remains a challenge (Liu et al., 2023b).
-To address these challenges, we introduce SensorLLM, a human-intuition inspired framework that aligns numerical sensor data with natural language.
-Unlike image-text pairs, sensor data consists of multivariate signals with complex and varied patterns, making annotation particularly difficult.
-To tackle this, we release SensorQA, a question–answer (QA) dataset of aligned sensor-text pairs spanning diverse HAR tasks.
-SensorLLM is trained on this dataset to learn intuitive mappings between sensor readings and descriptive language, enabling LLMs to interpret sensor data through natural interactions—without any modification to the LLM architecture (see Figure 1).
-Existing methods (Jin et al., 2024a; Sun et al., 2024a) have explored condensed text prototypes for alignment, but these approaches often lack interpretability and require extensive tuning to select suitable prototypes.
-In contrast, we propose an automatic text generation approach that aligns with human intuition by deriving descriptive trend-based text directly from TS data using statistical analyses and predefined templates.
+The latter approach avoids retraining but introduces key challenges: (i) Numerical encoding issues.
+Language model tokenizers, designed for text, struggle with numerical values, treating consecutive numbers as independent tokens (Nate Gruver and Wilson, 2023) and failing to preserve temporal dependencies (Spathis and Kawsar, 2023).
+(ii) Sequence length constraints.
+Complex numerical sequence often exceeds LLMs’ maximum context length, leading to truncation, information loss, and increased computational costs.
+(iii) Multivariate complexity.
+LLMs process univariate inputs, making it difficult to encode multivariate sensor data in a way that retains inter-channel dependencies.
+(iv) Prompt engineering challenges.
+Designing effective prompts that enable LLMs to interpret numerical sensor readings, detect trends and classify activities remains a challenge (Liu et al., 2023b).
+Unlike image-text pairs, sensor data comprises multi-channel numerical signals with diverse characteristics, making direct interpretation and annotation particularly difficult.
+Existing methods (Jin et al., 2024a; Sun et al., 2024) have explored condensed text prototypes for alignment, but these approaches often lack interpretability and require extensive tuning to select suitable prototypes.
+To address these challenges, we propose SensorLLM, a human-intuition inspired framework that aligns wearable sensor data with natural language without modifying the LLM itself.
+We propose an automatic text generation approach that aligns with human intuition by deriving descriptive trend-based text directly from time-series data using statistical analyses and predefined templates (see Figure 1).
 This method is precise, scalable, and interpretable, eliminating the need for manual annotations while preserving essential sensor characteristics.
@@
 Sensor-Language Alignment Stage.
-Our SensorQA pairs each sensor window with QA pairs to align sensor features with natural language in a structured format.
-A pretrained TS encoder extracts temporal features from the sensor data, which are then projected into a space interpretable by the LLM, mitigating text-specific tokenization issues.
-We also introduce special tokens for each sensor channel, enabling LLMs to effectively capture multivariate (multi-channel) dependencies.
+We automatically generate question-answer pairs to align sensor time-series with text while preserving temporal features using a pretrained encoder.
+The resulting embeddings are mapped into a space interpretable by the LLM, mitigating issues associated with text-specific tokenization.
+Additionally, we introduce special tokens for sensor channels, enabling LLMs to effectively capture multi-channel dependencies.
 Task-Aware Tuning Stage.
 The aligned embeddings are used for HAR, leveraging the LLM’s reasoning capabilities while keeping its parameters frozen.
-Importantly, our framework naturally supports sensor inputs of varying sequence lengths and arbitrary numbers of channels (i.e., multivariate TS data), a flexibility that prior approaches have struggled to accommodate.
+This design extends LLMs beyond their original training, addressing concerns raised by Tan et al.
+(2024) regarding their applicability to time-series data.
+Importantly, our framework naturally supports sensor inputs of varying sequence lengths and arbitrary numbers of channels (i.e., multivariate time-series), a flexibility that prior approaches have struggled to accommodate.
 To the best of our knowledge, this is the first approach to integrate sensor data into LLMs for sensor-based analysis and activity recognition. The key contributions of this work are:
-  • We release SensorQA, a novel human-intuitive dataset of aligned sensor-text QA pairs covering diverse HAR scenarios.
+  • We propose a human-intuitive approach for aligning sensor data with descriptive text, eliminating the need for manual annotations.
+SensorLLM support time-series inputs with varying sequence lengths and channel configurations, allowing broad and realistic HAR evaluation.
 We evaluate SensorLLM through text similarity metrics, human judgments, and LLM-based assessments, confirming its ability to capture temporal patterns for robust multimodal understanding.
-SensorQA and SensorLLM support sensor inputs with varying sequence lengths and channel configurations, allowing broad and realistic HAR evaluation.
   • SensorLLM achieves competitive results across five HAR datasets, matching or surpassing state-of-the-art models.
-Experiments further validate that modality alignment and task-specific prompts significantly enhance the LLM’s ability to interpret and classify sensor data.
-  • We show that SensorLLM maintains strong performance in the Task-Aware Tuning Stage, even when applied to datasets distinct from those used during alignment, highlighting its robustness and generalizability in HAR tasks.
+Experiments further validate that modality alignment and task-specific prompts significantly enhance LLM’s ability to interpret and classify sensor data.
+  • We show that SensorLLM maintains strong performance in the Task-Aware Tuning Stage, even when applied to datasets distinct from those used during alignment, highlighting its robustness and generalizability.
 2 Related Work
-In this section, we discuss recent developments in leveraging LLMs for time-series data, specifically focusing on two categories: (1) LLMs for time series as text and (2) Multimodal Large Language Models (MLLMs) for sensor data.
-A broader overview of other related works, including deep learning approaches to HAR and additional LLM-based forecasting methods, is provided in Appendix A.1.
+In this section, we review recent advances in applying LLMs to time-series data, with a focus on three directions: (1) treating general time series as text for LLM-based modeling, (2) leveraging Multimodal Large Language Models (MLLMs) for sensor data, and (3) employing LLMs for Human Activity Recognition (HAR).
+A broader survey of related work, including deep learning approaches to HAR and additional LLM-based forecasting methods, is provided in Appendix A.1.
 LLMs for Time Series as Text.
-While LLMs excel in processing natural language, applying them directly to time-series data poses unique challenges (Spathis and Kawsar, 2024).
+While LLMs excel in processing natural language, applying them directly to time-series data poses unique challenges (Spathis and Kawsar, 2023).
 Certain methods address this by treating time-series signals as raw text, using the same tokenization as natural language.
-Notable examples include PromptCast (Xue and Salim, 2023), which transforms numeric inputs into textual prompts for zero-shot forecasting, and LLMTime (Gruver et al., 2024), which encodes time-series as numerical strings for GPT-like models.
-However, due to the lack of specialized tokenizers for numeric sequences, LLMs may fail to capture crucial temporal dependencies and repetitive patterns (Spathis and Kawsar, 2024).
-To mitigate these issues, several works employ time-series encoders before mapping the resulting embeddings to language model spaces (Liu et al., 2024a; Zhou et al., 2023c; Xia et al., 2024), thus aligning sensor embeddings with textual embeddings in a contrastive or supervised manner.
+Notable examples include PromptCast (Xue and Salim, 2024), which transforms numeric inputs into textual prompts for zero-shot forecasting, and LLMTime (Gruver et al., 2023), which encodes time-series as numerical strings for GPT-like models.
+However, due to the lack of specialized tokenizers for numeric sequences, LLMs may fail to capture crucial temporal dependencies and repetitive patterns (Spathis and Kawsar, 2023).
+To mitigate these issues, several works employ time-series encoders before mapping the resulting embeddings to language model spaces (Liu et al., 2024a; Zhou et al., 2023b; Xia et al., 2024), thus aligning sensor embeddings with textual embeddings in a contrastive or supervised manner.
 MLLMs for Sensor Data.
@@
 This approach enables wearable AI applications like motion-based media search and LM-based multimodal reasoning, showcasing how sensor data can be integrated into broader multimodal frameworks.
+LLMs for Human Activity Recognition.
+Despite the success of LLMs in NLP, applying them to HAR remains challenging, as inertial signals are difficult to ground in language.
+Xia et al.
+(2023) introduce an unsupervised pipeline that employs a two-stage prompting scheme to infer activities from object sequences without manual descriptions.
+IMUGPT Leng et al.
+(2023) follows a data-synthesis approach: it prompts LLMs to generate diverse textual activity descriptions, converts them into 3D motion sequences, and then renders virtual IMU streams for training.
+HARGPT Ji et al.
+(2024) directly prompts LLMs to classify sensor time-series data, downsampling signals to 10Hz to enable zero-shot HAR from raw IMU inputs, and demonstrates that on simple tasks LLMs can match or even surpass traditional and deep learning models without domain-specific adaptation.
+ZARA Li et al.
+(2025) instead frames HAR as an agentic LLM workflow, supporting plug-and-play zero-shot motion time-series analysis with knowledge-guided reasoning and retrieval-enhanced evidence.
 3 Methods
-In this work, we propose SensorLLM, a two-stage framework that aligns wearable sensor data with descriptive text by a high-precision corpus of question-answering pairs created without any human annotations and tailored for wearable sensor reasoning.
+In this work, we propose SensorLLM, a two-stage framework that aligns wearable sensor data with descriptive text by high-precision question-answering pairs created without any human annotations and tailored for wearable sensor reasoning.
 Our aim is to build a multimodal model capable of interpreting and reasoning over time-series (TS) signals.
 As shown in Figure 2, SensorLLM consists of three core components: (1) a pretrained LLM, (2) a pretrained TS embedder, and (3) a lightweight MLP alignment module.
-In the Sensor–Language Alignment stage, a generative model aligns sensor readings with text, and in the Task–Aware Tuning stage, a lightweight classifier is added on top of the LLM to perform HAR.
-Crucially, only the alignment MLP and this classifier are trainable—both the backbone LLM and the TS embedder remain frozen—resulting in just 5.67% (535.9 M) of parameters being fine-tuned in the first stage and 0.12% (10.5 M) in the second, making training extremely efficient.
+In the Sensor-Language Alignment Stage, a generative model aligns sensor readings with text , and in the Task–Aware Tuning stage, a lightweight classifier is added on top of the LLM to perform HAR.
+Crucially, only the alignment MLP and this classifier are trainable, while the backbone LLM and TS embedder remain frozen.
+This design results in just 5.67% (535.9 M) of parameters being fine-tuned in the first stage and 0.12% (10.5 M) in the second, making training extremely efficient.
 Figure 2: Our proposed SensorLLM framework: (a) Sensor-Language Alignment Stage, where a generative model aligns sensor readings with automatically generated text; (b) Task-Aware Tuning Stage, where a classification model leverages the aligned modalities to perform HAR.
-3.1 SensorQA Dataset
-Aligning TS data with natural language is challenging due to the lack of rich semantic annotations beyond class labels, making manual labeling costly and impractical (Deldari et al., 2024; Haresamudram et al., 2024).
-While prior works often rely on fixed text prototypes (Sun et al., 2024b; Jin et al., 2024a), we introduce SensorQA —a scalable and human-intuitive dataset designed to bridge sensor data and language through structured QA pairs.
-SensorQA is built on the idea that TS signals naturally exhibit semantic patterns—such as trends and statistical behaviors—that can be described in natural language.
-Using predefined templates (Appendix A.2), descriptive QA pairs are automatically generated for each sensor window without any human annotations.
-Each pair includes information such as sensor type, time range, and observed trends.
-Templates are randomly combined to enhance diversity.
+3.1 Sensor-Text Data Generation
+Aligning time-series data with text for LLM-based tasks is challenging due to the lack of rich semantic labels beyond class annotations, making manual annotation impractical (Deldari et al., 2024; Haresamudram et al., 2025).
+While prior works rely on predefined text prototypes (Sun et al., 2024; Jin et al., 2024a), we aim for a more human-intuitive representation of sensor data.
+We argue that time-series data inherently contains semantic patterns that can be expressed through descriptive text, from simple numerical trends to statistical insights.
+To achieve this, we automatically generate descriptive text by analyzing observed trends and fluctuations in the data.
+Using predefined templates, we construct diverse question-answer (QA) pairs that capture trend changes while ensuring accuracy and scalability.
+These templates (see Appendix A.2) are randomly combined to enhance diversity.
 For example:
   • (1)
-    The time-series data represents readings taken from a <S> sensor between <$t_{s}$> and <$t_{e}$> seconds.
+    The time-series data represents readings taken from a <S> sensor between $t_{s}$> and $t_{e}$> seconds.
   • (2)
-    To sum up, the data exhibited a <T> trend for a cumulative period of <$t_{t}$> seconds.
+    To sum up, the data exhibited a <T> trend for a cumulative period of $t_{t}$> seconds.
 where T and S denote specific trends and sensor types, and t corresponds to numerical values.
 3.2 Sensor-Language Alignment
-As shown in Figure 2 (a), the Sensor–Language Alignment stage uses a generative model to create multimodal sentences by combining single-channel sensor readings with text from our SensorQA dataset.
+As shown in Figure 2 (a), the Sensor-Language Alignment stage employs a generative model to create multimodal sentences that combine single-channel sensor readings with textual descriptions.
 The sensor data is represented as a matrix $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels and $T$ is the sequence length.
@@
 Loss is computed only on generated tokens, ensuring SensorLLM effectively integrates sensor and text embeddings to produce coherent, contextually appropriate responses.
+3.3 Task-Aware Tuning
+As shown in Figure 2 (b), the Task-Aware Tuning stage refines the multimodal sensor-text embeddings for HAR.
+This stage integrates multi-channel sensor readings with activity labels, aligning temporal patterns with human activities.
+The input sensor data $\mathbf{X}$ is segmented into overlapping windows of size $L$ with a 50% overlap (Li et al., 2018), forming segments $\mathbf{X}_{S}\in\mathbb{R}^{S\times C\times L}$, where $S$ is the number of segments and $C$ is the number of channels.
+The pretrained alignment module from the first stage maps sensor data to activity labels, preserving inter-channel dependencies while learning activity-related patterns.
+Input Embedding.
+For each sensor channel $c$, we retrieve its aligned sensor embeddings $\hat{o}_{s}^{c}$ from the pretrained alignment module.
+These embeddings are then concatenated across all channels, along with their corresponding statistical features (mean and variance), to form the final input embedding:
+$\hat{z}=\hat{o}_{s}^{1}\oplus\hat{o}_{s}^{2}\oplus\dots\oplus\hat{o}_{s}^{C}\oplus\hat{z}_{\text{stat}},$ (3)
+where $\hat{z}_{\text{stat}}$ represents the statistical information, and $C$ is the number of channels. This ensures the model integrates both temporal and statistical characteristics for HAR.
+Loss Function.
+The input token sequence is processed by the LLM, yielding a latent representation $\mathbf{H}\in\mathbb{R}^{K\times D}$, where $K$ is the number of tokens and $D$ is the embedding dimension.
+Due to causal masking, we extract the final hidden state, $\mathbf{h}=\mathbf{H}_{K}$, which encodes all preceding token information.
+This pooled vector is passed through a fully connected layer to produce a prediction vector of size $M$, where $M$ is the number of activity classes.
+The final class probabilities $\hat{y}_{i}$ are obtained via the softmax function, and the model is optimized using cross-entropy loss:
+$\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i},$ (4)
+where $y_{i}$ is the ground truth label.
 Metric USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
@@
 The column GPT-4o denotes trend descriptions generated by GPT-4o, while the row GPT-4o indicates evaluations conducted by GPT-4o on the model outputs.
-3.3 Task-Aware Tuning
-As shown in Figure 2 (b), the Task-Aware Tuning stage refines the multimodal sensor-text embeddings for HAR.
-This stage integrates multi-channel sensor readings with activity labels, aligning temporal patterns with human activities.
-The input sensor data $\mathbf{X}$ is segmented into overlapping windows of size $L$ with a 50% overlap (Li et al., 2018), forming segments $\mathbf{X}_{S}\in\mathbb{R}^{S\times C\times L}$, where $S$ is the number of segments and $C$ is the number of channels.
-The pretrained alignment module from the first stage maps sensor data to activity labels, preserving inter-channel dependencies while learning activity-related patterns.
-Input Embedding.
-For each sensor channel $c$, we retrieve its aligned embeddings $\hat{o}_{s}^{c}$ from the pretrained alignment module.
-These are concatenated across all channels, along with statistical features (mean and variance) from our SensorQA dataset, to form the final input embedding:
-$\hat{z}=\hat{o}_{s}^{1}\oplus\hat{o}_{s}^{2}\oplus\dots\oplus\hat{o}_{s}^{C}\oplus\hat{z}_{\text{stat}},$ (3)
-where $\hat{z}_{\text{stat}}$ represents the statistical information. This ensures the model integrates both temporal and statistical characteristics for HAR.
-Loss Function.
-The input token sequence is processed by the LLM, yielding a latent representation $\mathbf{H}\in\mathbb{R}^{K\times D}$, where $K$ is the number of tokens and $D$ is the embedding dimension.
-Due to causal masking, we extract the final hidden state, $\mathbf{h}=\mathbf{H}_{K}$, which encodes all preceding token information.
-This pooled vector is passed through a fully connected layer to produce a prediction vector of size $M$, where $M$ is the number of activity classes.
-The final class probabilities $\hat{y}_{i}$ are obtained via the softmax function, and the model is optimized using cross-entropy loss:
-$\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i},$ (4)
-where $y_{i}$ is the ground truth label.
-Method USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
-F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy
-PatchTST 45.2±$\pm 1.48$ 45.6±$\pm 2.19$ 86.8±$\pm 0.84$ 86.0±$\pm 0.71$ 82.0±$\pm 0.71$ 81.2±$\pm 0.84$ 80.0±$\pm 1.58$ 79.4±$\pm 1.34$ 35.6±$\pm 0.89$ 66.2±$\pm 1.10$
-Ns-Transformer 52.6±$\pm 2.30$ 51.8±$\pm 2.86$ 88.0±$\pm 0.71$ 87.4±$\pm 0.55$ 78.8±$\pm 0.84$ 78.8±$\pm 0.84$ 77.2±$\pm 1.48$ 75.8±$\pm 1.48$ 34.8±$\pm 1.10$ 65.4±$\pm 0.55$
-Informer 51.2±$\pm 1.30$ 51.6±$\pm 1.52$ 86.6±$\pm 1.14$ 86.4±$\pm 0.89$ 78.0±$\pm 1.58$ 78.6±$\pm 1.34$ 74.0±$\pm 0.71$ 72.8±$\pm 0.84$ 35.6±$\pm 0.55$ 66.8±$\pm 0.84$
-Transformer 49.6±$\pm 1.67$ 50.6±$\pm 0.55$ 85.4±$\pm 0.89$ 85.2±$\pm 1.10$ 77.0±$\pm 0.71$ 77.6±$\pm 0.89$ 75.2±$\pm 1.30$ 74.6±$\pm 1.34$ 32.8±$\pm 0.84$ 65.4±$\pm 0.89$
-iTransformer 48.4±$\pm 1.82$ 49.6±$\pm 1.67$ 81.8±$\pm 0.84$ 81.8±$\pm 0.84$ 76.6±$\pm 0.55$ 75.8±$\pm 0.45$ 80.4±$\pm 1.14$ 80.0±$\pm 1.22$ 19.8±$\pm 0.84$ 62.4±$\pm 0.89$
-TimesNet 52.2±$\pm 2.39$ 52.6±$\pm 2.07$ 87.4±$\pm 1.14$ 86.6±$\pm 1.14$ 76.2±$\pm 1.92$ 77.4±$\pm 1.14$ 78.4±$\pm 1.52$ 77.2±$\pm 1.48$ 34.8±$\pm 0.84$ 65.8±$\pm 1.79$
-GPT4TS 54.2±$\pm 2.05$ 56.0±$\pm 1.58$ 88.2±$\pm 0.84$ 87.6±$\pm 0.55$ 80.4±$\pm 0.89$ 79.8±$\pm 0.45$ 76.4±$\pm 1.14$ 75.4±$\pm 1.14$ 32.8±$\pm 1.10$ 62.2±$\pm 1.92$
-Chronos+MLP 44.2±$\pm 1.30$ 44.0±$\pm 0.71$ 82.2±$\pm 0.84$ 81.2±$\pm 0.84$ 79.8±$\pm 0.45$ 79.8±$\pm 0.45$ 83.0±$\pm 0.71$ 82.0±$\pm 0.71$ 38.0±$\pm 0.71$ 68.2±$\pm 0.84$
-DeepConvLSTM 48.8±$\pm 2.39$ 50.6±$\pm 2.41$ 89.2±$\pm 0.84$ 89.2±$\pm 0.84$ 78.4±$\pm 1.52$ 78.2±$\pm 1.10$ 75.0±$\pm 1.87$ 76.0±$\pm 1.00$ 40.4±$\pm 0.89$ 69.4±$\pm 1.14$
-DeepConvLSTMAtt 54.0±$\pm 2.12$ 54.4±$\pm 3.21$ 89.6±$\pm 1.14$ 89.4±$\pm 1.14$ 79.2±$\pm 1.30$ 79.6±$\pm 1.14$ 77.4±$\pm 2.19$ 76.8±$\pm 1.48$ 41.4±$\pm 0.55$ 70.4±$\pm 0.55$
-Attend 60.2±$\pm 2.17$ 60.8±$\pm 1.92$ 93.2±$\pm 0.84$ 92.8±$\pm 0.45$ 84.6±$\pm 1.14$ 85.0±$\pm 0.71$ 83.4±$\pm 1.14$ 82.6±$\pm 1.14$ 43.6±$\pm 0.55$ $\pm 0.71$
-SensorLLM 61.2±$\pm 3.56$ 62.6±$\pm 3.36$ 91.2±$\pm 1.48$ 90.8±$\pm 1.30$ 86.2±$\pm 1.48$ 87.2±$\pm 0.84$ 89.4±$\pm 3.85$ 89.0±$\pm 3.54$ 48.6±$\pm 1.14$ 72.0±$\pm 0.71$
-Table 2: F1-macro and accuracy scores (%) for the Human Activity Recognition tasks, presented as the mean and standard deviation over 5 random repetitions.
-Bold for the best and underline for the second-best.
 4 Experiments
@@
 This guarantees that test data in the second stage remains unseen during alignment, ensuring a fair evaluation of generalization.
-We select Chronos as the TS embedder since it has not been pre-trained on HAR-specific data, allowing us to evaluate the robustness of our approach in adapting to raw, domain-agnostic sensor signals.
+We select Chronos as the TS embedder because it has not been pre-trained on motion sensor data, making it an ideal candidate for evaluating the robustness of our approach in adapting to raw, domain-agnostic sensor signals.
 4.1 Datasets
@@
 These datasets vary widely in subject counts, sensor placement, sampling rates, channel configurations, and activity types, covering both controlled laboratory conditions and free-living environments.
-SensorQA (see Appendix A.7) is built on these five benchmarks and supports both alignment and activity classification.
 All datasets are publicly available, containing no personally identifiable information, thus posing minimal ethical or privacy concerns.
@@
 Results.
-Table 1 compares SensorLLM and GPT-4o on the Sensor Data Understanding task.
-SensorLLM consistently outperforms GPT-4o across all metrics, with outputs that align more closely with ground truth and exhibit stronger trend understanding and coherence.
-In contrast, GPT-4o often struggles with complex numerical data and trend detection (Yehudai et al., 2024).
-When serving as an evaluator, GPT-4o also indicates a preference for SensorLLM.
-Human evaluation, conducted on shorter sequences, likewise favors SensorLLM.
-Performance on CAPTURE-24 is comparatively lower, likely due to longer sequences being trained under fixed parameters.
-Overall, these results validate the effectiveness of our alignment method in enabling LLMs to interpret complex TS data.
-Qualitative examples are provided in Appendix A.10.
+Table 1 compares SensorLLM and GPT-4o on the Sensor Data Trend Analysis tasks, showing that our model consistently outperforms GPT-4o across all metrics.
+BLEU-1, ROUGE-1, ROUGE-L, and METEOR primarily focus on surface-level lexical or n-gram overlaps.
+SBERT and SimCSE can capture factual correctness or deeper semantic similarities.
+Across all metrics, SensorLLM generates trend descriptions more closely aligned with the ground truth.
+GPT-4o evaluations further highlight SensorLLM’s superior ability to capture trend details and coherence, whereas GPT-4o struggles with complex numerical data and trend observations (Yehudai et al., 2024).
+Human evaluation also favors SensorLLM, particularly for shorter sequences.
+CAPTURE-24 results are weaker compared to other datasets, likely due to its longer sequences being trained with the same parameters.
+Overall, these findings validate the effectiveness of our Sensor-Language Alignment method in enhancing LLMs’ ability to interpret complex numerical sequence.
+Appendix A.9 provides qualitative examples of outputs from both models.
+Method USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
+F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 2: F1-macro and accuracy scores (%) for the Human Activity Recognition tasks, presented as the mean and standard deviation over 5 random repetitions.
+Bold for the best and underline for the second-best.
 4.3 Human Activity Recognition
@@
 Baselines.
-We benchmark SensorLLM against 11 baselines across two categories: (i) TS models—Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021), NS-Transformer (Liu et al., 2022), PatchTST (Nie et al., 2023), TimesNet (Wu et al., 2023a), and iTransformer (Liu et al., 2024c); (ii) HAR models—DeepConvLSTM (Ordóñez and Roggen, 2016), DeepConvLSTMAttn (Murahari and Plötz, 2018), and Attend (Abedin et al., 2021).
+We benchmark SensorLLM against 11 baselines across two categories: (i) TS models—Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021), NS-Transformer (Liu et al., 2022), PatchTST (Nie et al., 2023), TimesNet (Wu et al., 2023), and iTransformer (Liu et al., 2024c); (ii) HAR models—DeepConvLSTM (Ordóñez and Roggen, 2016), DeepConvLSTMAttn (Murahari and Plötz, 2018), and Attend (Abedin et al., 2021).
 We also include Chronos+MLP and GPT4TS (Zhou et al., 2023a) for a more comprehensive comparison.
@@
 Results.
-Table 2 reports F1-macro and accuracy scores (%) averaged over five runs.
-SensorLLM achieves the best performance on four out of five datasets (USC-HAD, PAMAP2, MHealth, CAPTURE-24), and ranks second on UCI-HAR, slightly behind Attend.
-It shows notable gains on challenging datasets such as CAPTURE-24 and MHealth, demonstrating strong performance in real-world and long-sequence settings.
-Compared to Chronos+MLP, which uses the same TS encoder, SensorLLM significantly improves both F1-macro and accuracy, highlighting the effectiveness of our alignment strategy in enabling LLMs to understand and classify sensor data.
-Strong results on both F1-macro and accuracy indicate that SensorLLM performs well in both overall prediction and per-class balance, showing robust generalization across diverse sensor configurations, activity types, and data collection environments.
+Table 2 reports the F1-macro and accuracy scores (%) averaged over five random runs.
+SensorLLM achieves the best performance on four out of five datasets (USC-HAD, PAMAP2, MHealth, CAPTURE-24), and ranks second on UCI-HAR, indicating strong performance across diverse sensor settings.
+On CAPTURE-24, SensorLLM reaches 48.6% F1-macro, a +5.0% absolute gain over Attend (43.6%).
+It also leads on USC-HAD (61.2%, +1.0%), PAMAP2 (86.2%, +1.6%), and MHealth (89.4%, +6.0%), establishing state-of-the-art performance on these datasets.
+On UCI-HAR, SensorLLM achieves 91.2%, slightly below Attend (93.2%).
+However, Chronos+MLP offers only a marginal improvement over iTransformer on UCI-HAR (82.2% vs.
+81.8%), suggesting that Chronos embeddings alone are of limited utility for this dataset.
+In contrast, using the same time-series encoder (Chronos), SensorLLM substantially boosts both F1-macro and accuracy, with especially large gains on challenging, real-world, long-sequence settings such as CAPTURE-24 and MHealth, highlighting the effectiveness of our alignment strategy in enabling LLMs to understand and classify sensor data.
+Overall, the consistent gains in both F1-macro and accuracy indicate balanced per-class performance alongside strong overall prediction, with robust generalization across sensor configurations, activity types, and data-collection environments.
 5 Ablation Studies
@@
 w/o prompts w/ prompts w/o prompts w/ prompts
-USC-HAD 43.4±$\pm 2.88$ 45.0±$\pm 1.58$ 49.6±$\pm 1.67$ 61.2±$\pm 3.56$
-UCI-HAR 80.0±$\pm 2.12$ 82.0±$\pm 1.58$ 89.2±$\pm 1.10$ 91.2±$\pm 1.48$
-PAMAP2 74.2±$\pm 2.28$ 75.4±$\pm 3.05$ 83.0±$\pm 0.71$ 86.2±$\pm 1.48$
-MHealth 76.6±$\pm 1.34$ 77.4±$\pm 3.13$ 86.6±$\pm 1.14$ 89.4±$\pm 3.85$
-CAPTURE-24 44.8±$\pm 0.84$ 46.0±$\pm 0.71$ 47.2±$\pm 0.84$ 48.6±$\pm 1.14$
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 3: F1-macro scores for models trained with and without text prompts. Task-only refers to conducting Task-Aware Tuning directly bypassing the alignment stage.
@@
 The results highlight the benefits of multimodal inputs, which enrich sensor data representations and improve HAR accuracy.
-More broadly, the ability to jointly process sensor data and textual prompts underscores the potential of LLMs for more generalizable and interpretable sensor-driven applications.
+More broadly, the ability to process both sensor signals and textual prompts not only enhances classification performance but also underscores the potential of LLMs for more generalizable and interpretable sensor-driven applications.
 Figure 3: Effect of the number of alignment module layers.
@@
 Smaller SensorLLM Still Compete.
-To address computational feasibility for deployment in resource-constrained environments, we evaluate SensorLLM-3b—a lighter variant built with Chronos-base and LLaMA3.2-3b.
+To address computational feasibility for deployment in resource-constrained environments, we evaluate SensorLLM-3b, a lighter variant built with Chronos-base and LLaMA3.2-3b.
 Experiments were conducted on USC-HAD, UCI-HAR, and MHealth.
 As shown in Figure 4, SensorLLM-3b achieves slightly lower performance than SensorLLM-8b, reflecting the trade-off between model size and accuracy.
-Nevertheless, it remains competitive—outperforming Attend on USC-HAD and MHealth, and closely trailing it on UCI-HAR.
+Nevertheless, it remains competitive, outperforming Attend on USC-HAD and MHealth, and closely trailing it on UCI-HAR.
 These results suggest that SensorLLM-3b provides a strong balance between efficiency and performance, making it a viable choice for real-world, resource-limited applications.
@@
 w/o ST w/ ST
-MHealth 89.6±$\pm 2.70$ 90.2±$\pm 3.11$ 15
-PAMAP2 84.4±$\pm 1.14$ 85.8±$\pm 0.84$ 27
+[…]
+[…]
 Table 4: Effect of special tokens on HAR based on two-layer alignment MLP. ST refers to special tokens.
@@
 Stage 2 Stage 1 F1-macro
-USC-HAD UCI-HAR 61.6±$\pm 2.07$
-USC-HAD 61.2±$\pm 3.56$
-UCI-HAR UCI-HAR 91.2±$\pm 1.48$
-USC-HAD 91.0±$\pm 1.41$
+[…]
+[…]
+[…]
+[…]
 Table 5: F1-macro scores for cross-dataset experiments.
 6 Conclusions
-We present SensorLLM, a multimodal framework that aligns sensor data with natural language through a QA format at a human-perception level, moving beyond machine-level alignment.
-It effectively captures complex sensor patterns, achieves strong performance on HAR tasks, and generalizes well without requiring dataset-specific alignment.
-Experiments demonstrate its robustness across variable-length sequences, multivariate inputs, and textual metadata.
-To support future research, we release our code and the SensorQA dataset, constructed from five public HAR benchmarks, to advance time-series and language integration, particularly in low-resource domains.
+We present SensorLLM, the first multimodal framework that aligns sensor data with automatically generated text at a human-perception level, moving beyond machine-level alignment.
+SensorLLM effectively captures complex sensor patterns, achieving superior performance in HAR tasks.
+Experiments across diverse datasets demonstrate its robustness in handling variable-length sequences, multi-channel inputs, and textual metadata.
+Cross-dataset results further highlight its strong generalizability without requiring dataset-specific alignment.
+This work establishes a foundation for Sensor-Text MLLMs, with potential applications for sensor data reasoning.
+We release our code and data generation pipeline to facilitate future research on integrating time-series and text, particularly in low-resource domains.
 7 Limitations
 While SensorLLM demonstrates strong performance in aligning sensor data with LLMs, certain limitations remain, offering directions for future exploration.
-Classifier-Based Design.
-To ensure fair comparisons with existing HAR models, we adopt a classifier for downstream tasks rather than fully leveraging the LLM’s generative capabilities.
-While our results demonstrate that the Sensor–Language Alignment Stage can generalize across datasets, relying on a fixed-class classifier may limit adaptability to new activity categories.
-Although zero-shot adaptability is a valuable direction, we did not explore it here due to the lack of comparable baselines.
-To the best of our knowledge, no prior work supports generalization to unseen activity classes under variable-length and variable-channel sensor input, as our framework does, making fair comparison on zero-shot settings infeasible at this stage.
-Future work could explore generative or prompt-based approaches to support broader applications such as activity discovery or open-set recognition.
+Choice of Classifier.
+It is worth noting that the primary contribution of our work lies in proposing a novel TS-text alignment strategy and rigorously validating its effectiveness.
+To ensure fair comparisons with existing HAR baselines, we adopt a classifier for downstream tasks rather than fully exploiting the LLM’s generative abilities.
+While this design allows rigorous validation of our proposed alignment method, it also imposes limitations: relying on a fixed-class classifier may constrain adaptability to new activity categories and does not fully leverage the reasoning potential of LLMs.
+Although our framework is compatible with other downstream heads, such as a language modeling (LM) head for prompt-based generation, we leave this exploration to future work.
+Investigating generative or zero-shot approaches could enable broader applications, including activity discovery and open-set recognition.
 Scope of Sensor-Text Alignment.
 Our alignment focuses on mapping sensor data to trend-descriptive text, demonstrating clear benefits for LLM-based HAR.
-However, human-intuitive descriptions of sensor data extend beyond trend changes—incorporating frequency-domain features, periodicity, and higher-order patterns may further enhance an LLM’s ability to interpret time-series data.
+However, human-intuitive descriptions of sensor data extend beyond trend changes, incorporating frequency-domain features, periodicity, and higher-order patterns may further enhance an LLM’s ability to interpret time-series data.
 Future research could investigate whether aligning text with alternative sensor characteristics improves time-series reasoning.
-This could expand the potential of multimodal NLP applications in sensor-driven tasks beyond activity recognition.
+This could expand the potential of multimodal LLM applications in sensor-driven tasks beyond activity recognition.
 8 Acknowledgements
@@
 Zhou et al.
-(2023b) demonstrates that pre-trained language and image models, such as a Frozen Pretrained Transformer (FPT), can be adapted for diverse time-series tasks like classification, forecasting, and anomaly detection, leveraging self-attention mechanisms to bridge the gap between different data types and achieving state-of-the-art performance across various tasks.
+(2023a) demonstrates that pre-trained language and image models, such as a Frozen Pretrained Transformer (FPT), can be adapted for diverse time-series tasks like classification, forecasting, and anomaly detection, leveraging self-attention mechanisms to bridge the gap between different data types and achieving state-of-the-art performance across various tasks.
 Jin et al.
@@
 Sun et al.
-(2024b) proposes TEST, an innovative embedding technique that integrates time-series data with LLMs through instance-wise, feature-wise, and text-prototype-aligned contrast, yielding improved or comparable results across various applications.
+(2024) proposes TEST, an innovative embedding technique that integrates time-series data with LLMs through instance-wise, feature-wise, and text-prototype-aligned contrast, yielding improved or comparable results across various applications.
 Chang et al.
@@
 (2024b) introduces UniTime, a unified model that leverages language instructions and a Language-TS Transformer to handle multivariate time series across different domains, demonstrating enhanced forecasting performance and zero-shot transferability.
-LLMs for Human Activity Recognition.
-While LLMs like ChatGPT have demonstrated remarkable performance in various NLP tasks, their effectiveness in HAR remains limited due to challenges in interpreting sensor data.
-These models often struggle to distinguish between activities that share similar objects, requiring more advanced prompt engineering to highlight activity-specific details.
-Xia et al.
-(2023) proposed an unsupervised approach to HAR using ChatGPT, leveraging two-stage prompts to infer activities from object sequences without manual descriptions.
-The method demonstrates superior performance on three benchmark datasets, marking a significant advancement in applying language models to activity recognition tasks.
-Similarly, Ji et al.
-(2024) explored LLMs for zero-shot HAR using raw IMU data, showing that GPT-4 can outperform both traditional and deep learning models in simple HAR tasks without domain-specific adaptations, highlighting LLMs’ potential in sensor-based systems.
-A.2 SensorQA Generation
-For SensorQA, we generate text data from sensor readings using predefined sentence templates (Tables 6, 7, 8).
+A.2 Sensor–Text Data Pair Generation
+We generate text data from sensor readings using predefined sentence templates (Tables 6, 8, 9).
 These templates are randomly selected to create diverse question-answer (QA) pairs.
@@
 Table 6: Examples of answer templates used for trend descriptions.
+Dataset Stage 1 Stage 2
+Train Test Train Test
+USC-HAD 300,744 58,704 22,790 4,555
+UCI-HAR 128,292 25,932 7,352 2,947
+PAMAP2 738,666 271,674 14,163 5,210
+MHealth 283,020 60,780 4771 2,039
+CAPTURE-24 72,714 35,688 61,327 30,138
+Table 7: Training and testing sensor-text QA pairs counts for Stage 1 and Stage 2.
 Trend Description Templates
@@
 The AI analyzes a sensor time-series dataset (N points, sampled at {sample_rate}Hz) to answer specific questions, demonstrating its analytical capabilities and the potential for human-AI collaboration in interpreting sensor data.
+Sensor-Language Alignment Stage
+focuses on aligning uni-variate sensor sequence of variable length with descriptive textual responses and includes two types of QA tasks:
+  • Trend Analysis QA, which describes how the signal changes within the window.
+  • Trend Summary QA, which summarizes the overall behavior across a window in a concise natural language phrase.
+Task-Aware Tuning Stage
+focuses on using multi-variate sensor sequences to perform human activity classification, leveraging the aligned representations learned in the alignment stage.
+This stage contains statistical information from each sensor channel as part of the input representation.
+The distribution of training and testing data across both stages is summarized in Table 7.
 A.3 Chronos
@@
 Table 12: Dataset classes and Proportions
-A.7 SensorQA
-We introduce SensorQA, a novel question-answering dataset designed to align time-series sensor data with human-interpretable natural language.
-Each sample pairs a segment of sensor input with natural language questions and answers, capturing trends, patterns, or activity-level semantics.
-Sensor-Language Alignment Stage
-of SensorQA focuses on aligning uni-variate sensor sequence of variable length with descriptive textual responses and includes two types of QA tasks:
-  • Trend Analysis QA, which describes how the signal changes within the window.
-  • Trend Summary QA, which summarizes the overall behavior across a window in a concise natural language phrase.
-Task-Aware Tuning Stage
-focuses on using multi-variate sensor sequences to perform human activity classification, leveraging the aligned representations learned in the alignment stage.
-This stage of SensorQA contains statistical information from each sensor channel as part of the input representation.
-Dataset Stage 1 Stage 2
-Train Test Train Test
-USC-HAD 300,744 58,704 22,790 4,555
-UCI-HAR 128,292 25,932 7,352 2,947
-PAMAP2 738,666 271,674 14,163 5,210
-MHealth 283,020 60,780 4771 2,039
-CAPTURE-24 72,714 35,688 61,327 30,138
-Table 12: Training and testing sample counts for Stage 1 and Stage 2 across datasets of SensorQA.
-The distribution of training and testing data across both stages is summarized in Table 12.
-A.8 Baselines for Task-Aware Tuning Stage
+A.7 Baselines for Task-Aware Tuning Stage
 In Task-Aware Tuning Stage, we compare SensorLLM against several state-of-the-art baseline models for time-series classification and human activity recognition (HAR).
@@
 This design retains local semantics and allows for longer historical context, significantly improving long-term forecasting accuracy.
-TimesNet (Wu et al., 2023a).
+TimesNet (Wu et al., 2023).
 TimesNet is a versatile backbone for time series analysis that transforms 1D time series into 2D tensors to better capture intraperiod and interperiod variations.
@@
 The authors also found that self-attention functions similarly to PCA, providing a theoretical explanation for the versatility of transformers.
-A.9 Evaluation Metrics for Task-Aware Tuning Stage
+A.8 Evaluation Metrics for Task-Aware Tuning Stage
 In our evaluation, we use the F1-macro score to assess the model’s performance across datasets.
@@
 The formula for the F1-macro score is:
-[…]
-[…]
-[…]
+$\text{F1-macro}=\frac{1}{C}\sum_{i=1}^{C}\text{F1}_{i}$ (8)
+where $C$ is the total number of classes, and $\text{F1}_{i}$ is the F1 score for class $i$. The F1 score for each class is calculated as:
+$\text{F1}_{i}=\frac{2\times\text{Precision}_{i}\times\text{Recall}_{i}}{\text{Precision}_{i}+\text{Recall}_{i}}$ (9)
 The precision and recall for each class are defined as:
@@
 This metric ensures that performance is evaluated fairly across all classes, regardless of the frequency of each label, making it a robust measure for imbalanced datasets.
-A.10 Sensor-Language Alignment Stage Output Examples
+A.9 Sensor-Language Alignment Stage Output Examples
 Tables LABEL:tab:stage1_ouput_1 and  LABEL:tab:stage1_ouput_2 present two examples of the trend analysis results generated by SensorLLM and GPT-4o based on the input sensor data.